### v1

# secReferences

Abstract
Typical causal effects are defined based on the marginal distribution of potential outcomes. However, many real-world applications require causal estimands involving the joint distribution of potential outcomes to enable more nuanced treatment evaluation and selection. In this article, we propose a novel framework for identifying and estimating the joint distribution of potential outcomes using multiple experimental datasets. We introduce the assumption of transportability of state transition probabilities for potential outcomes across datasets and establish the identification of the joint distribution under this assumption, along with a regular full-column rank condition. The key identification assumptions are testable in an overidentified setting and are analogous to those in the context of instrumental variables, with the dataset indicator serving as “instrument”. Moreover, we propose an easy-to-use least-squares-based estimator for the joint distribution of potential outcomes in each dataset, proving its consistency and asymptotic normality. We further extend the proposed framework to identify and estimate principal causal effects. We empirically demonstrate the proposed framework by conducting extensive simulations and applying it to evaluate the surrogate endpoint in a real-world application.
Keywords: Causal Inference, Data Fusion, Principal Stratification, Surrogacy Evaluation
Introduction
Estimating the causal effect of a treatment variable on an outcome is a fundamental scientific problem, central to many fields such as social and biomedical sciences (Imbens and Rubin 2015; Pearl and Mackenzie 2018; Hernán and Robins 2020; Rosenbaum 2020). Typical causal effects like average treatment effects are defined in terms of the marginal distribution of potential outcomes. However, in many real-world applications, causal estimands involving the joint distribution of potential outcomes are required to enable more nuanced treatment evaluation and selection. Examples include the probability of causation (Pearl 1999; Dawid and Musio 2022; Lu et al. 2023), the effect of persuasion (Jun and Lee 2023, 2024), treatment benefit rates and treatment harm rates (Shen et al. 2013; Yin et al. 2018b; Kallus 2022a, b; Li et al. 2023b; Wu et al. 2024a), and so on.
Identifying the joint distribution of potential outcomes from a single dataset presents a significant challenge. This difficulty arises from the fact that only one potential outcome is observed for each individual, which is known as the fundamental problem in causal inference (Holland 1986). Consequently, the joint distribution of potential outcomes is generally unidentifiable, even in randomized controlled trials (Pearl et al. 2016). Instead of identifying the joint distribution, many studies focus on deriving bounds for causal estimands involving the joint distribution (e.g., Tian and Pearl 2000; Zhang et al. 2013; Yin et al. 2018a; Li et al. 2022a; Kallus 2022b).
In this article, we propose a novel framework for identifying and estimating the joint distribution of potential outcomes using multiple experimental datasets. Specifically, we make the following three contributions.
First, for binary outcomes, we introduce the assumption of transportability of state transition probability from $Y^{0}$ to $Y^{1}$ — meaning that the distribution of $Y^{1}$ given $Y^{0}$ is invariant across datasets — where $Y^{0}$ and $Y^{1}$ denote the potential outcomes under control and treatment, respectively. Under this assumption, along with a regular full-column rank condition, we establish the identification of the joint distribution of $(Y^{0},Y^{1})$ in each dataset. In clinical studies, we can regard the value of $Y^{0}$ as a measure of the baseline physical status of a patient, so the transportability of state transition probabilities means that patients with the same baseline status would have the same distribution of status if treated. This transportability assumption is weaker than the widely used assumption of transportability of causal effects across datasets, and the full-column rank condition is readily satisfied when the number of datasets is greater than or equal to two (see Section 3.2 for details). Essentially, the proposed identification assumptions are analogous to those in the context of instrumental variables, with dataset indicator serving as “instrument”. We show analogous identification results for general categorical outcomes, requiring the number of datasets to be no smaller than the number of outcome categories. Moreover, we extend the results to identify the joint distribution of potential outcomes in a new dataset that contains only control units.
Second, we propose a least-squares-based method for estimating the transition probabilities from $Y^{0}$ to $Y^{1}$, thereby estimating the joint distribution of $(Y^{0},Y^{1})$ in each dataset. The key idea behind the proposed estimation method is that $\mathbb{P}(Y^{1})$ can be expressed as a linear combination of $\mathbb{P}(Y^{1}\mid Y^{0})$ and $\mathbb{P}(Y^{0})$. We also establish the consistency and asymptotic normality of the proposed estimators of transition probabilities. Additionally, we propose methods for testing the key assumption of transportability of state transition probabilities in an overidentified setting, where the number of datasets exceeds the number of outcome categories.
Third, when an additional binary variable $S$ is observed after the treatment and before the outcome, we extend the proposed framework to identify and estimate the principal stratification average causal effects (PSACEs), i.e., the average causal effects within the principal stratification defined by the joint values of $(S^{0},S^{1})$, where $S^{0}$ and $S^{1}$ represent the potential outcomes for $S$ (Frangakis and Rubin 2002). We consider two sets of identification assumptions in this setting. The first involves the transportability of state transition probabilities from $(S^{0},Y^{0})$ to $(S^{1},Y^{1})$, along with a rank condition requiring at least four datasets. If the monotonicity condition $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$ additionally holds, then only two or more datasets are needed. The second includes the transportability of state transition probabilities from $(S^{0},S^{1})$ to $Y^{0}$ (and $Y^{1}$), the monotonicity condition $S^{1}\geq S^{0}$, and a full-column rank condition requiring at least two datasets. We further propose least-squares-based methods for estimating PSACEs. These complement the identification strategy and the Bayesian estimation methods in Jiang et al. (2016).
The idea of combining two or more datasets to identify causal effects and enhance estimation efficiency has garnered significant attention in the field of causal inference (Colnet et al. 2023; Degtiar and Rose 2023; Wu et al. 2024b; Athey et al. 2019; Yang and Ding 2020; Imbens et al. 2024; Kallus and Mao 2024; Hu et al. 2025; Hünermund and Bareinboim 2025). However, most existing methods consider causal effects involving the marginal distribution of potential outcomes. In contrast, we focus on identifying and estimating causal estimands that involve the joint distribution of potential outcomes, thus extending and complementing this class of methodologies.
The rest of this article is organized as follows. In Section 2, we outline the basic setup. In Section 3, we present the identification assumptions and establish the identification for the joint distribution of potential outcomes. In Section 4, we propose the least-squares-based method for estimating the transition probabilities and present a method for testing the key assumption of transportability of state transition probabilities. Section 5 extends the proposed method to identify and estimate PSACEs. Section 6 evaluates the finite-sample performance of the proposed method through extensive simulations. Section 7 demonstrates the proposed methods in a real-world application, and Section 8 concludes the paper.
Setup
Let $A\in\mathcal{A}=\{0,1\}$ represent a binary treatment, where $A=1$ indicates the treatment condition and $A=0$ indicates the control condition. Let $X\in\mathcal{X}$ be the pre-treatment covariates, and $Y\in\mathcal{Y}$ be the outcome of interest. To define causal estimands, we adopt the potential outcome framework (Neyman 1923; Rubin 1974) and denote $Y^{0}$ and $Y^{1}$ as the potential outcomes under treatment and control, respectively. We make the stable unit treatment value assumption (i.e., no multiple treatment versions and no interference across units), which ensures the well-definedness of potential outcome $Y^{a}$ (Imbens and Rubin 2015). Accordingly, the observed outcome is linked to the potential outcomes through $Y=(1-A)Y^{0}+AY^{1}$.
Suppose we have access to data from individuals in a collection of $m$ trials, indexed by ${\mathcal{G}}=\{1,...,m\}$. For each trial $g\in{\mathcal{G}}$, the observed data consist of realizations of independent random tuples $\{(G_{i}=g,X_{i},A_{i},Y_{i}),i=1,...,n_{g}\}$, where $n_{g}$ denotes the number of individuals in trial $g$, and we let $n=\sum_{g=1}^{m}n_{g}$ represent the total sample size. We adopt a non-nested design (Dahabreh et al. 2021) in which the observed samples from different trials are independent. Without loss of generality, under a superpopulation model $\mathbb{P}$, we assume that the observed data in trial $g$ form an i.i.d. sample from $\mathbb{P}(\cdot\mid G=g)$, which implies that all observed data are obtained by stratified sampling from a population that is stratified by $G$.
In this paper, our primary objective is to identify and estimate the joint distribution of potential outcomes $Y^{0},Y^{1}$ from the data of multiple trials. Given that there might be potential distribution discrepancies across trials, we are particularly interested in the distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for each trial $g\in\mathcal{G}$. Notably, this task diverges from and is more challenging than what most existing data-combination-based causal inference methods can handle. The latter mainly focuses on estimating causal effects defined in terms of the marginal distribution of potential outcomes (e.g., Rosenman et al. 2023; Colnet et al. 2023; Wu et al. 2024b).
We mainly consider the case of binary outcome $Y\in\{0,1\}$ and $m\geq 3$ (i.e., at least three trials). This case is of particular interest because it leads to overidentification and offers an opportunity to test the key identification assumption. Additionally, for identifying the principal causal effects, we require $m\geq 4$ due to the extra complexity introduced by principal stratification, see Section 5 for details.
Nonparametric Identification
In what follows, to illustrate the main ideas, we suppress the dependence on covariates $X$ for simplicity. Alternatively, the statements presented in this section can be interpreted as being implicitly conditional on $X$.
3.1 Basic Assumptions
We begin by presenting the basic assumptions required to identify $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$.
Assumption 1 (Unconfoundedness of trials).
For all $g\in{\mathcal{G}}=\{1,...,m\}$, (i) $A\perp\!\!\!\perp(Y^{0},Y^{1})\mid G=g$, and (ii) $10<\mathbb{P}(A=1\mid G=g)<$ 1.
Assumption 1 means that the treatment $A$ is randomized within each trial and both treatment arms are assigned with a positive probability. This assumption trivially holds for data from randomized trials, which ensures the identification of the average treatment effect within each trial. However, it is insufficient to identify the joint distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$. Therefore, we proceed to invoke Assumption 2 to borrow information from different trials.
Assumption 2 (Transportability of state transition probability).
$Y^{1}\perp\!\!\!\perp G\mid Y^{0}$.
Assumption 2 requires that the probabilities of transitioning from the state $Y^{0}$ to the state $Y^{1}$ are invariant across trials, i.e., $\mathbb{P}(Y^{1}=1\mid Y^{0},G=1)=\mathbb{P}(Y^{1}=1\mid Y^{0},G=2)=\cdots=\mathbb{P}(Y^{1}=1\mid Y^{0},G=m)$. For instance, in Section 7, our experiments focus on Adjuvant Colon Clinical Trials (ACCTs), where the potential outcomes $Y^{1},Y^{0}$ refer to cancer survival with and without the treatment respectively. Hence the value of $Y^{0}$ can be interpreted as a measure of the baseline physical status of a patient. In other words, Assumption 2 means that given the baseline status $Y^{0}$, the survival probability $Y^{1}$ of a patient after receiving the treatment would be the same across different trials. Alternatively, we may understand Assumption 2 through some stylized structural models. For example, the assumption holds if $Y^{1}$ or $G$ depends on only $Y^{0}$, up to some exogenous noises, i.e., $Y^{1}=h(Y^{0},\epsilon^{y})$ or $G=h(Y^{0},\epsilon^{g})$ for a function $h$ with noises $\epsilon^{y}$ and $\epsilon^{g}$ satisfying $\epsilon^{y}\perp\!\!\!\perp G$ and $\epsilon^{g}\perp\!\!\!\perp Y^{1}$ respectively. Moreover, we may adopt the model in Athey and Imbens (2006): $Y^{a}=h(a,U)$ for $a=0,1$ and a latent variable $U$. Although the distribution of $U$ may differ across trials, $Y^{0}$ may provide sufficient information for $U$ (e.g., when $y^{0}\mapsto h(0,u)$ is an invertible mapping), so that we have $G\perp\!\!\!\perp U\mid Y^{0}$, which further implies Assumption 2.
Importantly, Assumption 2 is testable when the number of trials $m$ is strictly larger than the number of outcome categories. For a binary outcome, this means $m\geq 3$. In this setting, the joint distributions of potential outcomes for the $m$ trials are overidentified under Assumption 2. In Section 4.2, we develop a Chi-square test for Assumption 2, which is applied to the ACCTs data to validate this assumption. Moreover, we also use graphs to visually assess this assumption in Section 7.
Assumption 2 is weaker than the assumption of transportability of full potential outcome distribution, $G\perp\!\!\!\perp(Y^{0},Y^{1})$, a widely used assumption in the literature on causal inference via data combination, see e.g., Dahabreh et al. (2019), Dahabreh et al. (2020), Lee et al. (2021), Li et al. (2022b), Li et al. (2023a), Li et al. (2023c), Colnet et al. (2023), Degtiar and Rose (2023), and Wu et al. (2024b). In our analysis of the ACCTs data, as shown in Figure 1 in Section 7, we observe that ATEs vary significantly across trials, clearly refuting this strong assumption. However, Assumption 2 appears plausible according to our hypothesis testing results and visual illustrations in Section 7.
3.2 Identifiability for Joint Distribution of Potential Outcomes
An important implication of Assumptions 12 is that the joint distributions of potential outcomes are identifiable from the data of multiple trials. Assumption 2 imposes invariant potential outcome state transition probabilities across trials:
$\pi_{b|a}:=\mathbb{P}(Y^{1}=b\mid Y^{0}=a)=\mathbb{P}(Y^{1}=b\mid Y^{0}=a,G=g),\quad a,b=0,1;g\in\mathcal{G}.$
By law of total probability, the following holds for all $g\in\mathcal{G}$:
$\mathbb{P}(Y^{1}=1\mid G=g)=\mathbb{P}(Y^{0}=0\mid G=g)\cdot\pi_{1|0}+\mathbb{P}(Y^{0}=1\mid G=g)\cdot\pi_{1|1},$ (1)
where $\mathbb{P}(Y^{1}=1\mid G=g)$ and $\mathbb{P}(Y^{0}=a\mid G=g)$ for $a=0,1$ are all identifiable under Assumption 1. Therefore, the identification of $\pi_{1|0}$ and $\pi_{1|1}$ can be viewed as finding the solution to Eq. (1), a system of $m$ linear equations in terms of $\pi_{1|0}$ and $\pi_{1|1}$. The solution is unique (i.e., point identification) under the following full-rank condition on the coefficient matrix in the linear equation system.
Condition 1 (Full-column rank).
The matrix $(\mathbb{P}(Y^{0}=0\mid G=\cdot),\mathbb{P}(Y^{0}=1\mid G=\cdot)_{m\times 2}$ has a full-column rank (i.e., rank $2$).
Condition 1 requires that $G$ has sufficient variability relative to $Y^{0}$, and it is satisfied only if $m\geq 2$ for a binary outcome. Moreover, this condition only involves identifiable quantities under Assumption 1, so in principle it is testable.
Theorem 1 (Binary outcome).
Under Assumptions 12 and Condition 1, the distributions $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ and $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ for $g\in{\mathcal{G}}$ are identifiable.
Theorem 1 shows the identification of the conditional distribution $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ and joint distribution $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ for each trial $g\in\mathcal{G}$, given the marginal distributions $\mathbb{P}(Y^{1}\mid G=g)$ and $\mathbb{P}(Y^{0}\mid G=g)$ identified from multiple datasets, under the assumptions of trial unconfoundedness (Assumption 1), state transition probability transportability (Assumption 2), and the full-column rank condition (Condition 1). Under these conditions, the conditional distribution $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ can be uniquely identified from the solution to the equation system in Eq. (1), which, together with the identifiable marginal distribution $\mathbb{P}(Y^{0}\mid G=g)$, gives the identification of the joint distribution $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ for each trial $g$.
The identification Assumption 2 and Condition 1 resemble the standard assumptions in the context of instrumental variable (IV) analysis (Angrist et al. 1996; Imbens 2004). We may view the dataset indicator $G$ as an “IV”, the potential outcome $Y^{0}$ as the “treatment”, and the potential outcome $Y^{1}$ as the “outcome”. Then Assumption 2 corresponds to the IV exclusion restriction assumption, and Condition 1 corresponds to the full rank assumption of IV on treatment (or IV relevance), which requires a strong enough association between $G$ and $Y^{0}$.
Following Theorem 1, we can identify various causal estimands involving the joint distribution of potential outcomes. Below are some examples.
Example 1 (Probability of causation).
Causal inference involves not only evaluating the effects of causes but also deducing the causes of given effects (Dawid and Musio 2022), where the latter is also referred to as attribution analysis (Pearl 2009; Pearl et al. 2016; Pearl and Mackenzie 2018). The probability of sufficient causation (PS) and the probability of necessary causation (PN) are two standard quantities for attribution analysis. They are defined by $\textup{PS}(A\Rightarrow Y)=\mathbb{P}(Y^{1}=1\mid A=0,Y=0)$ and $\textup{PN}(A\Rightarrow Y)=\mathbb{P}(Y^{0}=0\mid A=1,Y=1)$ respectively. For example, the quantity $\textup{PN}(A\Rightarrow Y)$ can be written as $\mathbb{P}(Y^{0}=0,Y^{1}=1\mid A=1)/\mathbb{P}(Y^{1}=1\mid A=1)$, where the denominator is an identifiable quantity and the numerator equals to $\mathbb{P}(Y^{0}=0,Y^{1}=1)$ for randomized treatment assignment.
Example 2 (Effect of persuasion).
Let $A\in\{0,1\}$ be a binary indicator for an individual’s exposure to certain persuasive information, and let $Y$ be a binary indicator representing the individual’s behavior. Jun and Lee (2023) defined the persuasion rate as $\mathbb{P}(Y^{1}=1\mid Y^{0}=0)$, which quantifies the proportion of individuals in the population who changed their behavior as a result of their exposure to persuasive information.
Example 3 (Treatment benefit and harm rates).
Denote $Y=1$ as a favorable outcome (e.g., survival) and $Y=0$ as an unfavorable outcome (e.g., death). The average treatment harm rate (Shen et al. 2013; Wu et al. 2024a) is defined as $\textup{THR}=\mathbb{P}(Y^{0}=1,Y^{1}=0),$ which quantifies the percentage of individuals experiencing worse outcomes under treatment than under control. Similarly, the average treatment benefit rate is defined as $\textup{TBR}=\mathbb{P}(Y^{0}=0,Y^{1}=1).$
3.3 Generalization to a Control-Only Target Population
In some scenarios, the target data may contain only control units and not any treated units (Li et al. 2023c). For example, this could arise if the target population has only received an old drug ($A=0$) and researchers want to assess the efficacy of a new drug ($A=1$) unavailable to this population. This subsection aims to generalize identification results to this setting. To this end, we consider a simple random sample from the target population with only control units (denoted by $G=0$), in addition to the $m$ observed experimental datasets.
Assumption 3 (Control-only target population).
We have $A=0$ for $G=0$, so $Y=Y^{0}$ in the target population.
Under Assumption 3, the treatment $A$ has no variation, so the conditional independence $A\perp\!\!\!\perp(Y^{0},Y^{1})\mid G=0$ holds naturally. Thus, Assumption 3 is a special case of Assumption 1(i) for the target population. On the dataset $G=0$, we can straightforwardly identify $\mathbb{P}(Y^{0}=a\mid G=0)$. Moreover, we can apply Theorem 1 to identify the conditional potential outcome distribution (i.e., $\pi_{b\mid a}$ for $a,b=0,1$), from the $m$ experimental datasets. These together lead to the identification of the joint distribution on the target population $\mathbb{P}(Y^{1}=b,Y^{0}=a\mid G=0)=\pi_{b|a}\cdot\mathbb{P}(Y^{0}=a\mid G=0)$. This is summarized as follows.
Corollary 1.
Under Assumptions 13 and Condition 1, the joint distribution $\mathbb{P}(Y^{1},Y^{0}\mid G=0)$ is identifiable.
3.4 Extension to Categorical Outcomes
The identifiability results for binary outcomes can be extended to the cases of categorical outcomes. Specifically, for categorical outcome with cardinality $k$, let $\mathbb{P}(Y^{0}\mid G)=(\mathbb{P}(Y^{1}=i\mid G=g))_{m\times k}$ be a $m\times k$ matrix whose $(g,i)$-th element is $\mathbb{P}(Y^{1}=i\mid G=g)$.
Condition 2.
The matrix $\mathbb{P}(Y^{0}|G)$ is full-column rank.
Similar to Condition 1, Condition 2 holds only when $m\geq k$. Under Assumptions 12 and Condition 2, we have the following identifiability results.
Theorem 2 (Categorical outcome).
Let $Y$ be a categorical outcome with $k$ possible values. Under Assumptions 12 and Condition 2, the joint distributions $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ for $g\in{\mathcal{G}}$ are identifiable.
Theorem 2 extends the result of Theorem 1 for identifying the joint distribution of potential outcomes. After establishing identifiability, we then consider the estimation of the joint distributions and the testing of the key Assumption 2.
Estimation and Inference
In this section, we first present an estimator of $\pi_{b|a}$ for $a,b=0,1$, which are key invariant parameters for identifying the joint distribution $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ by noting that $\mathbb{P}(Y^{1}=b,Y^{0}=a\mid G=g)=\pi_{b|a}\cdot\mathbb{P}(Y^{0}=a\mid G=g)$ for any $g\in{\mathcal{G}}$. Then, we provide a method to test Assumption 2 when $m=|\mathcal{G}|\geq 3$.
4.1 Estimation
We now present a simple least-squares-based estimation method for $\pi_{b|a}$, $a,b\in\{0,1\}$. Note that under Assumptions 12, equation (1) gives the following for $g\in\mathcal{G}$:
$\mathbb{P}(Y=1\mid G=g,A=1)=\pi_{1|0}\mathbb{P}(Y=0\mid G=g,A=0)+\pi_{1|1}\mathbb{P}(Y=1\mid G=g,A=0)$ (2)
When $m=2$, the quantities $\pi_{b|a}$ for $a,b\in\{0,1\}$ are just-identified as the number of parameters to be identified equals the number of equations. For example, when $m=2$, solving the two equations in equation (2) identifies $\pi_{1|0}$ and $\pi_{1|1}$. Then $\pi_{0|0}$ and $\pi_{0|1}$ are identified through $\pi_{0|0}=1-\pi_{1|0}$ and $\pi_{0|1}=1-\pi_{1|1}$, respectively. In contrast, if $m>2$, equation (2) includes more equations than parameters. In such a case, the parameters $\pi_{b|a}$ for $a,b\in\{0,1\}$ are overidentified.
In both just-identified and overidentified cases, we can estimate $\theta:=(\pi_{1|0},\pi_{1|1})$ via a linear least-squares estimator. Formally, we let $\tilde{Y}_{g}=\mathbb{P}(Y=1\mid G=g,A=1),$ $\tilde{X}_{1g}=\mathbb{P}(Y=0\mid G=g,A=0),$ and $\tilde{X}_{2g}=\mathbb{P}(Y=1\mid G=g,A=0)$, and denote $\tilde{X}_{g}=(\tilde{X}_{1g},\tilde{X}_{2g})^{\intercal}$. By equation (2), we have $\tilde{Y}_{g}=\tilde{X}_{g}^{\top}\theta$ for $g\in\mathcal{G}$. Under the full rank condition (Condition 1), we can further write $\theta$ as
$\theta=\left(\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{Y}_{g}.$
Although the conditional probabilities $\tilde{X}_{g}$ and $\tilde{Y}_{g}$ are unknown, they can be easily estimated by the corresponding sample frequencies, e.g., $\hat{Y}_{g}=\sum_{i=1}^{n}\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)\big{/}\sum_{i=1}^{n}\mathbb{I}(G_{i}=g,A_{i}=1)$. Given estimators $\hat{X}_{g}$ and $\hat{Y}_{g}$, the resulting estimator for $\hat{\theta}$ is
$\hat{\theta}=\left(\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g}.$
This amounts to the ordinary least squares (OLS) coefficient estimator with $\hat{Y}_{g}$ as the response and $\hat{X}_{g}$ as the covariates. It is noteworthy that while $\hat{\theta}$ resembles the OLS estimator in a linear model, it is different in two aspects: the “effective sample size” $m$ is fixed, and the true value $\theta$ also has a least-squares formulation. A potential strength of this estimator is that it does not need the individual-level data. Instead, it only involves some sample summary statistics $\hat{X}_{g}$ and $\hat{Y}_{g}$. This may be helpful for privacy protection (Han et al. 2023).
Next, we show that estimator $\hat{\theta}$ is consistent and asymptotically normal.
Theorem 3.
[…]
$V=$ $\frac{1}{m^{2}}\sum_{g=1}^{m}\mathrm{Var}\Biggl{\{}\tilde{Y}_{g}\begin{pmatrix}\dfrac{\mathbb{I}(Y=0,G=g,A=0)-\mathbb{P}(Y=0,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\\ \dfrac{\mathbb{I}(Y=1,G=g,A=0)-\mathbb{P}(Y=1,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\end{pmatrix}$
$+\dfrac{\mathbb{I}(Y=1,G=g,A=1)-\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(G=g,A=1)}\tilde{X}_{g}\Biggr{\}}.$
4.2 Testing Assumption 2 under Overidentifiability
If the parameter $\theta$ is overidentified, we can further test for Assumption 2—the key identification assumption, provided that Assumption 1 already holds. Specifically, given that Assumption 1 holds, Assumption 2 leads to the null hypothesis
$H_{0}:\tilde{Y}_{g}=\tilde{X}_{g}^{\intercal}\theta,\text{ for all }g=1,...,m,$
that is, there is a linear relationship between $\tilde{Y}_{g}$ and $\tilde{X}_{g}$ for all $g\in{\mathcal{G}}$. We use the following statistic to test Assumption 2,
$J=\sum_{g=1}^{m}\frac{(\hat{Y}_{g}-\hat{X}_{g}^{\intercal}\hat{\theta})^{2}}{\hat{\sigma}_{g}^{2}}.$
[…]
The test on Assumption 2 is analogous to the test of exclusion restriction assumption in the IV setting, such as Windmeijer (2019); Kiviet (2020); Carrasco and Doukali (2022), and overidentified testing method in the GMM framework (Newey 1985; Newey and McFadden 1994).
Extension to Principal Causal Effect
In this section, we extend the proposed method developed in Sections 34 to identify and estimate the principal causal effects, complementing and extending the framework of Jiang et al. (2016). Again, we focus on the binary outcome setting.
In addition to the observed variables $(G,A,Y)$, we have access to a binary post-treatment variable $S$ (e.g., surrogate endpoint) observed after the treatment $A$ and before the outcome $Y$. Let $S^{0}$ and $S^{1}$ be the potential outcomes of $S$ under treatment arms 00 and $1$, respectively. Our goal is to identify the joint distributions $\mathbb{P}(Y^{a},S^{0},S^{1}\mid G=g)$ for $g\in{\mathcal{G}}$ and $a=0,1$. If these joint distributions are identifiable, we can identify the principal stratification average causal effects (PSACEs) for each trial $g$ defined as follows:
$\text{PSACE}_{ab|g}=\mathbb{E}[Y^{1}-Y^{0}\mid S^{0}=a,S^{1}=b,G=g],\quad a=0,1;b=0,1.$
This quantity has many applications, such as noncompliance (Imbens and Angrist 1994), truncation by death (Rubin 2006), and surrogate endpoint evaluation (Jiang et al. 2016). See Frangakis and Rubin (2002) for more examples.
5.1 Extension
Following the similar spirit of the proposed method in Sections 34, and parallel to Assumptions 12 and Condition 1, we introduce the following identifiability Assumptions 45 and Condition 3.
Assumption 4.
$A\perp\!\!\!\perp(S^{0},S^{1},Y^{0},Y^{1})\mid G=g$ and $10<\mathbb{P}\left(A=1\mid G=g\right)<$ 1 for all $g\in\mathcal{G}$.
Assumption 5.
$G\perp\!\!\!\perp(S^{1},Y^{1})\mid S^{0},Y^{0}$.
Condition 3.
The number of trials $m\geq 4$, and the matrix $(\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=g),\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=g))_{m\times 4}$ is full-column rank.
Assumption 4 is similar to Assumption 1 and holds when all trials are randomized experiments. Assumption 5 corresponds to Assumption 2. In this setting, we view the values of $(S^{0},Y^{0})$ as the baseline physical status of a patient, in which case Assumption 5 again implies the transportability of state transition probabilities across different trials.
Corollary 2.
Under Assumptions 4, 5 and Condition 3, the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ for all $g\in\mathcal{G}$ are identifiable, and therefore $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ and $g\in\mathcal{G}$ are also identifiable.
Corollary 2 presents the identifiability of $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$. Essentially, Corollary 2 can be seen as an instantiation of Theorem 2 where $(Y^{1},S^{1})$ and $(Y^{0},S^{0})$ are virtually two categorical variables with four different values. This is why Condition 3 needs the number of trials $m$ to be no smaller than $4$. In fact, this condition can be weakened to the following condition, requiring only $m\geq 2$ rather than $m\geq 4$, under an additional monotonicity assumption (i.e., $Y^{1}\geq Y^{0}$ and $S^{1}\geq S^{0}$).
Condition 4.
(i) the matrix $(\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=g))_{m\times 2}$ is full-column rank; (ii) the matrix $(\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=g))_{m\times 2}$ is full-column rank.
The identification under this relaxed condition is formally stated as follows.
Theorem 4.
If $Y^{1}\geq Y^{0}$ and $S^{1}\geq S^{0}$, then under Assumptions 4, 5 and Condition 4, the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ for $g\in\mathcal{G}$ are identifiable, and therefore $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ and $g\in\mathcal{G}$ are also identifiable.
With the additional monotonicity assumption $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$, Theorem 4 extends Corollary 2, by relaxing the full rank requirement from Condition 3 to Condition 4. In addition, we remark that under the identification assumptions in Corollary 2 and Theorem 4, we can also use least-squares-based method for estimation. Please refer to Supplementary Material S2 for details.
If only partial monotonicity holds (i.e., $S^{1}\geq S^{0}$ or $Y^{1}\geq Y^{0}$, but not both), then under Assumptions 4, 5 and Condition 4, only half of the quantities in $\{\mathbb{P}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g),a,b,c,d=0,1;g\in\mathcal{G}\}$ are identifiable (see Supplementary Material S3). This means that partial monotonicity alone cannot completely weaken Condition 3 to Condition 4 for the sake of identifying the whole joint distribution. But partial monotonicity can indeed simplify the identification of Corollary 2 provided that Condition 3 still holds. The corresponding simplified estimation methods are given in Supplementary Material S4. Moreover, if we relax Assumption 5 to partial but not joint conditional independence $G\perp\!\!\!\perp S^{1}\mid(S^{0},Y^{0})$ and $G\perp\!\!\!\perp Y^{1}\mid(S^{0},Y^{0})$ in Corollary 2, then we can identify $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$ but not the full joint distribution needed to identify the principal causal effects. See Supplementary Material S5 for a detailed discussion.
5.2 Connection and Comparison
Our paper is closely related to Jiang et al. (2016), who aim to identify and estimate the principal effects by leveraging multiple trials. They adopt Assumption 4, along with the key Assumptions 67 and Condition 5 presented below.
Assumption 6.
$G\perp\!\!\!\perp Y^{a}\mid S^{0},S^{1}$ for $a=0,1$.
Assumption 7.
$S^{1}\geq S^{0}$.
Condition 5.
(i) The matrix $(\mathbb{P}(S^{0}=0,S^{1}=1\mid G=g),\mathbb{P}(S^{0}=1,S^{1}=1\mid G=g))_{m\times 2}$ is full-column rank; (ii) The matrix $(\mathbb{P}(S^{0}=0,S^{1}=1\mid G=g),\mathbb{P}(S^{0}=0,S^{1}=0\mid G=g))_{m\times 2}$ is full-column rank.
Assumption 6 means that there is no dependence between $G$ and $Y^{0}$ or $Y^{1}$ within the principal stratum defined by $(S^{0},S^{1})$. This implies that $\text{PSACE}_{ab|g}$ for $g\in{\mathcal{G}}$ are invariant across trials. In contrast, the assumptions in Section 5.1 allow different principal effects across trials.
The monotonicity condition in Assumption 7, together with the unconfoundedness Assumption 4, imply the identifiability of principal scores $\delta_{ab|g}=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$ for $a,b\in\{0,1\}$. Specifically, Assumption 7 implies $\delta_{10\mid g}=0$, thus the joint distribution $\mathbb{P}(S^{0},S^{1}\mid G=g)$ involves only three free parameters ($\delta_{00\mid g},\delta_{01\mid g},\delta_{11\mid g}$). These parameters satisfy three equations $\delta_{10\mid g}+\delta_{11\mid g}=\mathbb{P}(S^{0}=1\mid G=g),~{}\delta_{01\mid g}+\delta_{11\mid g}=\mathbb{P}(S^{1}=1\mid G=g),~{}\sum_{a=0}^{1}\sum_{b=0}^{1}\delta_{ab\mid g}=1.$ Solving these equations easily gives the identification of the principal scores.
Condition 5 is also a full rank condition, which, like Condition 4, also requires $m\geq 2$. Under Assumption 4 and the three above assumptions, Jiang et al. (2016) show the identifiability of $\text{PSACE}_{ab|g}$ in their Theorem 1. We reproduce their results in Theorem 5 below and provide an alternative proof (see Supplementary Material S1.7). This alternative proof will motivate our new estimator in Section 5.3.
Theorem 5.
Under Assumptions 4, 6, and 7, for for $g=1,...,m$, we have that
(a) $\mathbb{P}(Y^{1}\mid S^{0},S^{1},G=g)$ is identifiable if Condition 5(i) holds.
(b) $\mathbb{P}(Y^{0}\mid S^{0},S^{1},G=g)$ is identifiable if Condition 5(ii) holds.
(c) $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ are identifiable if Condition 5 holds.
In Theorem 5, the monotonicity condition $S^{1}\geq S^{0}$ in Assumption 6 is used to identify the principal scores as discussed above. One may wonder whether this condition can be replaced by an alternative condition, $G\perp\!\!\!\perp S^{1}\mid S_{0}$, that is also relevant for identifying the principal scores [^1] According to our theory in Section 3, the condition $G\perp\!\!\!\perp S^{1}\mid S_{0}$ is relevant for the identification of the joint distribution of $(S^{1},S^{0})$ and thus the principal scores.. Our Supplementary Material S6 gives a negative answer to this question, showing the importance of the monotonicity condition. Moreover, Theorem 4 also considers another monotonicity condition $Y^{1}\geq Y^{0}$. This alternative monotonicity condition is not relevant here because the distributions $\mathbb{P}(Y^{1}\mid S^{0},S^{1},G=g)$ and $\mathbb{P}(Y^{0}\mid S^{0},S^{1},G=g)$ and the principal effects $\text{PSACE}_{ab|g}$ do not involve the joint distribution of $Y^{1}$ and $Y^{0}$.
Without the monotonicity condition $S^{1}\geq S^{0}$, Proposition 2 of Jiang et al. (2016) further shows that a necessary condition for local identifiability of the principal effects is $m\geq 3$. In contrast, our previous Corollary 2, which also considers identification without any monotonicity condition, establishes sufficient conditions for global identifiability when $m\geq 4$. In this sense, our new identification results provide some alternative conditions that lead to a stronger identification, which complements the findings of Jiang et al. (2016).
5.3 Least-squares Estimation Based on Theorem 5
Under the identification in Theorem 5, Jiang et al. (2016) propose a Bayesian estimation method. In this part, we follow our new proof for Theorem 5 and extend the least-square estimation method in Section 4 to the setting of Theorem 5.
We first define the following parameters for $a,b\in\{0,1\}$:
$\pi_{1|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b),~{}\tilde{\pi}_{1|ab}=\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b).$
These are key invariant parameters across trials. Clearly, under the monotonicity condition $S^{1}\geq S^{0}$, $\pi_{1|10}$ and $\tilde{\pi}_{1|10}$ are undefined and do not need estimation. Now we describe how to estimate $\beta:=(\pi_{1|00},\pi_{1|01},\pi_{1|11})$ and $\gamma:=(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01},\tilde{\pi}_{1|11})$.
Step 1: estimate the principal scores $\delta_{ab|g}:=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$, and the probabilities $\mathbb{P}(Y^{1}=1\mid G=g)$, and $\mathbb{P}(Y^{0}=1\mid G=g)$. Specifically, for any given $g$, $\delta_{10|g}=0$ by Assumption 7, and
$\begin{cases}\delta_{11|g}={}&\mathbb{P}(S^{0}=1,S^{1}=1|G=g)=\mathbb{P}(S^{0}=1|G=g)=\mathbb{P}(S=1|A=0,G=g),\\ \delta_{01|g}={}&\mathbb{P}(S^{0}=0,S^{1}=1|G=g)=\mathbb{P}(S^{1}=1|G=g)-\delta_{11|g}\\ \quad\quad={}&\mathbb{P}(S=1|A=1,G=g)-\delta_{11|g},\\ \delta_{00|g}={}&1-\delta_{11|g}-\delta_{01|g}.\end{cases}$
In addition, $\mathbb{P}(Y^{1}=1\mid G=g)=\mathbb{P}(Y=1\mid G=g,A=1)$ and $\mathbb{P}(Y^{0}=1\mid G=g)=\mathbb{P}(Y=1\mid G=g,A=0)$. We can construct their estimators $\hat{\delta}_{11|g},\hat{\delta}_{01|g},\hat{\delta}_{00|g}$, ^$\hat{\mathbb{P}}(Y^{1}=1\mid G=g)$, and ^$\hat{\mathbb{P}}(Y^{0}=1\mid G=g)$ via replacing the population probabilities by the empirical frequencies and solving the resulting linear equations.
Step 2: estimate $\pi_{1|00}$ and $\tilde{\pi}_{1|11}$. Clearly, due to $S^{1}\geq S^{0}$, we have $\pi_{1|00}=\mathbb{P}(Y^{1}=1\mid S^{0}=0,S^{1}=0)=\mathbb{P}(Y^{1}=1\mid S^{1}=0)=\mathbb{P}(Y=1\mid S=0,A=1),$ and $\tilde{\pi}_{1|11}=\mathbb{P}(Y^{0}=1\mid S^{0}=1,S^{1}=1)=\mathbb{P}(Y^{0}=1\mid S^{0}=1)=\mathbb{P}(Y=1\mid S=1,A=0).$ These two can again be estimated by the empirical frequencies. We denote the estimators by $\hat{\pi}_{1|00}$ and $\hat{\tilde{\pi}}_{1|11}$ respectively.
Step 3: estimate $(\pi_{1|01},\pi_{1|11})$ and $(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01})$. Note that
$\mathbb{P}(Y^{1}=1|G=g)-\pi_{1|00}\delta_{00|g}={}$ $\pi_{1|01}\delta_{01|g}+\pi_{1|11}\delta_{11|g},\quad g=1,...,m,$
$\mathbb{P}(Y^{0}=1|G=g)-\tilde{\pi}_{1|11}\delta_{11|g}={}$ $\tilde{\pi}_{1|00}\delta_{00|g}+\tilde{\pi}_{1|01}\delta_{01|g},\quad g=1,...,m.$
Then we can estimate $(\pi_{1|01},\pi_{1|11})$ by running linear regression of ^$\hat{\mathbb{P}}(Y^{1}=1\mid G=g)-\hat{\pi}_{1|00}\hat{\delta}_{00|g}$ on $(\hat{\delta}_{01|g},\hat{\delta}_{11|g})$, and estimate $(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01})$ by running linear regression of ^$\hat{\mathbb{P}}(Y^{0}=1\mid G=g)-\hat{\tilde{\pi}}_{1|11}\hat{\delta}_{11|g}$ on $(\hat{\delta}_{00|g},\hat{\delta}_{01|g})$. Let $(\hat{\pi}_{1|01},\hat{\pi}_{1|11})$ and $(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01})$ be the corresponding estimators, and denote $\hat{\beta}=(\hat{\pi}_{1|00},\hat{\pi}_{1|01},\hat{\pi}_{1|11})$ and $\hat{\gamma}=(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01},\hat{\tilde{\pi}}_{1|11})$.
Step 4: estimate principal effects $\text{PSACE}_{ab|g}$. The estimators are given by $\hat{\pi}_{1|ab}-\hat{\tilde{\pi}}_{1|ab}$ for $ab=00,01$, and 11. Due to the monotonicity condition $S^{1}\geq S^{0}$, $\text{PSACE}_{ab|g}$ for $ab=10$ is undefined and needs no estimation.
Simulation
We perform extensive simulation studies to evaluate the finite-sample performance of the proposed method. We consider both scenarios, with and without the post-treatment variable $S$. The R codes to reproduce the results for both simulation and application are available at https://github.com/pengwu1224/The-Promises-of-Multiple-Experiments. Throughout this simulation, the number of trials is set to 10, and for each trial $g$, the binary treatment $A$ is randomly assigned with probability $\mathbb{P}(A=1\mid G=g)=0.5$, the sample size is set to 100, 200, and 500, respectively.
Study I (without the post-treatment variable $S$). We first examine the methods developed in Sections 34 by considering two data-generating processes for $(Y^{1},Y^{0})$.
  • (C1)
    $\mathbb{P}(Y^{1}=1\mid Y^{0},G=g)=\text{expit}(Y^{0}-0.5)$ for $g=1,...,10$, where $\text{expit}(x)=\exp(x)/\{1+\exp(x)\}$ is the standard logistic function. For each trial $g=1,2,...,10$, the potential outcome $Y^{0}$ follows from a Bernoulli distribution with $\mathbb{P}(Y^{0}=1\mid G=g)=0.5+(g-1)/30$, i.e., taking evenly spaced values at equal intervals from 0.5 to 0.8.
  • (C2)
    $\mathbb{P}(Y^{1}=1\mid Y^{0},G=g)=\text{expit}(Y^{0}+0.5)$ for $g=1,...,10$ and $Y^{0}$ is generated according to the process described in (C1).
Assumptions 12 and Condition 1 hold for both cases (C1) and (C2) in Study I. The true values of $\theta=(\pi_{1|0},\pi_{1|1})=(\mathbb{P}(Y^{1}=1\mid Y^{0}=0),\mathbb{P}(Y^{1}=1\mid Y^{0}=1))$ are $(0.378,0.622)$ and $(0.622,0.818)$ for cases (C1) and (C2), respectively. We replicate each simulation case 1,000 times and calculate the Bias, SD, ESE, and CP95 as evaluation metrics, where Bias and SD represent the Monte Carlo bias and standard deviation of the point estimates over the 1,000 replicates, ESE denotes the square root of the average of the estimated asymptotic variances, and CP95 refers to the coverage proportion of the 95% confidence intervals. Both ESE and CP95 are calculated using the estimated asymptotic variance obtained from 100 bootstraps for each replicate.
Table 1 summarizes the numerical results for the proposed estimator of $\theta$ for cases (C1)–(C2). From the table, we observe that the Bias is small and decreases as the sample size increases, demonstrating the consistency of the proposed estimator. As the sample size grows, ESE approaches SD, and CP95 converges to its nominal value of 0.95, indicating the asymptotic normality of the proposed estimator and validating the bootstrap method for estimating asymptotic variance. Additionally, the numerical results for both cases (C1) and (C2) follow similar patterns, suggesting the stability of the proposed estimator across different true values of $\theta$.
Table 1: Simulation results for cases (C1)–(C2).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C1) $\pi_{1|0}$ 0.041 0.136 0.141 0.938 0.017 0.104 0.107 0.946 0.009 0.068 0.070 0.952
$\pi_{1|1}$0.024 0.074 0.077 0.943 -0.009 0.057 0.058 0.945 -0.006 0.037 0.038 0.956
(C2) $\pi_{1|0}$ 0.037 0.123 0.122 0.937 0.024 0.092 0.092 0.944 0.007 0.060 0.060 0.951
$\pi_{1|1}$0.021 0.067 0.067 0.935 -0.013 0.049 0.050 0.939 -0.003 0.032 0.033 0.947
Note: Bias and SD are the Monte Carlo bias and standard deviation over the 1000 simulations of the point estimates, ESE and CP95 are the estimated asymptotic variances and coverage proportions of the 95% confidence intervals based on 100 bootstraps, respectively.
Study II (with the post-treatment variable $S$). We then explore the proposed method in the presence of a post-treatment variable $S$. Two additional data-generating processes for the potential outcomes $(S^{0},S^{1},Y^{0},Y^{1})$ are considered.
  • (C3)
    $\mathbb{P}(Y^{1}=1\mid Y^{0},S^{0},G=g)=\text{expit}\{(S^{0}+Y^{0}+1)/2\}$ and $\mathbb{P}(S^{1}=1\mid Y^{0},S^{0},G=g)=\text{expit}\{(S^{0}+Y^{0}-1)/2\}$, and $S^{0}$ and $Y^{0}$ are independent, both drawn from the Bernoulli distribution with success probability $300.5+(g-1)/$ 30.
  • (C4)
    $\mathbb{P}(Y^{1}=1\mid S^{0},S^{1},G=g)=\text{expit}\{(S^{0}+S^{1}+1)/2\}$ and $\mathbb{P}(Y^{0}=1\mid S^{0},S^{1},G=g)=\text{expit}\{(S^{0}+S^{1}-1)/2\}$, and $S^{0}$ and $S^{1}$ are independent, both drawn from the Bernoulli distribution, with the success probabilities $300.3+(g-1)/$ 30 and $300.5+(g-1)/$ 30, respectively. After generating $(S^{0},S^{1})$, we further adjust the value of $S^{0}$, setting $S^{0}$ to 0 if $S^{1}=0$ to ensure monotonicity.
Assumptions 45 and Condition 3 hold for case (C3), while Assumptions 4, 67 and Condition 5 hold for case (C4). In case (C3), we denote $\pi_{s|ab}=\mathbb{P}(S^{1}=1\mid S^{0}=a,Y^{0}=b)$ and $\pi_{y|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,Y^{0}=b)$ for $a,b=0,1$. The parameters of interest are $\theta=(\pi_{s|00},\pi_{s|01},\pi_{s|10},\pi_{s|11},\pi_{y|00},\pi_{y|01},\pi_{y|10},\pi_{y|11}),$ which are the key invariant parameters for estimating the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},S^{1},Y^{1}\mid G=g)$ for $g=1,...,10$. In case (C4), we define $\pi_{1|ab}=\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b)$ and $\tilde{\pi}_{1|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b)$ for $a,b=0,1$. By the monotonicity condition $S^{1}\geq S^{0}$, $\pi_{1|10}$ and $\tilde{\pi}_{1|10}$ are undefined. Thus, the key invariance parameters are $\theta=(\pi_{1|00},\pi_{1|01},\pi_{1|11},\tilde{\pi}_{1|00},\tilde{\pi}_{1|01},\tilde{\pi}_{1|11}).$ Moreover, it is noteworthy that cases (C3) and (C4) satisfy different sets of identifiability conditions. Therefore, we use the method proposed in Section 5.1 to estimate $\theta$ in Case (C3) the method outlined in Sections 5.25.3 to estimate $\theta$ in Case (C4).
Table 2: Simulation results for case (C3).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C3) $\pi_{s|00}$ 0.032 0.431 0.420 0.951 0.012 0.382 0.388 0.964 0.017 0.357 0.354 0.952
$\pi_{s|01}$ 0.005 0.427 0.375 0.933 -0.008 0.393 0.378 0.942 -0.010 0.381 0.365 0.945
$\pi_{s|10}$0.005 0.396 0.379 0.944 0.020 0.416 0.374 0.939 -0.003 0.375 0.362 0.948
$\pi_{s|11}$0.011 0.140 0.141 0.955 -0.011 0.118 0.121 0.952 0.000 0.101 0.102 0.963
$\pi_{y|00}$ 0.026 0.396 0.363 0.940 0.013 0.337 0.334 0.956 0.014 0.310 0.308 0.961
$\pi_{y|01}$ 0.007 0.363 0.324 0.937 -0.013 0.346 0.323 0.945 -0.021 0.349 0.320 0.943
$\pi_{y|10}$0.002 0.342 0.327 0.944 0.006 0.340 0.322 0.948 0.007 0.339 0.314 0.940
$\pi_{y|11}$0.009 0.127 0.120 0.937 -0.001 0.102 0.104 0.962 0.003 0.089 0.088 0.957
Table 3: Simulation results for case (C4).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C4) $\pi_{1|00}$ 0.040 0.155 0.142 0.931 0.033 0.130 0.128 0.938 0.020 0.094 0.101 0.954
$\pi_{1|01}$0.043 0.160 0.144 0.923 -0.037 0.131 0.131 0.943 -0.021 0.097 0.103 0.946
$\pi_{1|11}$0.000 0.039 0.039 0.943 0.001 0.028 0.028 0.949 -0.001 0.018 0.018 0.940
$\tilde{\pi}_{1|00}$ 0.001 0.037 0.037 0.942 0.000 0.026 0.026 0.952 -0.000 0.017 0.016 0.948
$\tilde{\pi}_{1|01}$ 0.019 0.109 0.113 0.953 0.012 0.092 0.094 0.962 0.004 0.064 0.065 0.947
$\tilde{\pi}_{1|11}$0.022 0.125 0.127 0.954 -0.013 0.106 0.106 0.956 -0.004 0.073 0.074 0.940
Tables 2 and 3 present the numerical results for cases (C3) and (C4), respectively. The simulation results are similar to those in cases (C1)–(C2), indicating that the extension of the proposed method presented in Section 5 performs well.
Application to the Adjuvant Colon Clinical Trials
We demonstrate the proposed methodology using the data from phase III adjuvant colon clinical trials (ACCTs). The initial ACCTs data consist of 20,898 patients from 18 randomized phase III clinical trials, with the enrollment period spanning from 1977 to 1999 (Sargent et al. 2005). Among these 18 trials, the data from 10 trials are publicly available in Baker et al. (2012), including a total of 9,102 patients. In each trial, we have a contingency table of observed frequencies for three binary variables: treatment $A$, surrogate $S$, and outcome $Y$. Here $S=0$ indicates cancer recurrence within 3 years and $S=1$ otherwise; $Y=0$ denotes mortality within 5 years and $Y=1$ indicates survival beyond 5 years; $A=1$ and $A=0$ denote receiving treatment and not, respectively.
The goal of the ACCTs is to determine whether cancer recurrence within 3 years ($S$) can serve as an appropriate surrogate for overall survival with a 5-year follow-up ($Y$). Using the ACCTs data, Sargent et al. (2005) identified a strong correlation between $S$ and $Y$. From the perspective of principal stratification in causal inference, Jiang et al. (2016) investigated this question by first estimating PSACEs and then evaluating the surrogate using the causal necessity and causal sufficiency criteria (Gilbert and Hudgens 2008): for all $g\in{\mathcal{G}}$, causal necessity requires $\text{PSACE}_{11|g}=0$ and $\text{PSACE}_{00|g}=0$, and causal sufficiency requires $\text{PSACE}_{10|g}\neq 0$ and $\text{PSACE}_{01|g}\neq 0$.
In this article, we further explore the problem by estimating both the joint distributions of $\mathbb{P}(S^{0},S^{1}\mid G=g)$ and $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$, as well as the PSACEs under different sets of identification assumptions. Notably, for the ACCTs data, Assumptions 1 and 4 naturally hold due to randomization. In addition, by calculation, all full rank regularity conditions (Conditions 1, 3, 4, and 5) are satisfied.
7.1 State Transition Probabilities and Joint Distribution of Potential Surrogates and Potential Outcomes
To get an intuitive understanding of the heterogeneity across different trials, we estimate the average treatment effects (ATEs) of $A$ on $S$ (or $Y$) for each trial $g\in{\mathcal{G}}$ by computing the average contrasts of $S$ (or $Y$) between the groups $(A=1,G=g)$ and $(A=0,G=g)$. As illustrated in the barplots of Figure 1, the ATEs of $A$ on $S$ and $Y$ vary significantly across trials, indicating substantial heterogeneity. This means that the transportability of causal effects across datasets imposed in many existing literature is implausible. In addition, the ATEs of $A$ on $S$ and $Y$ show a similar pattern, indicating that $S$ may have the potential to serve as an appropriate surrogate for $Y$.
Figure 1: (a) Estimated ATEs of $A$ on $S$, and scatter plots of ^$\hat{\mathbb{P}}(S^{1}=1\mid G=g)$ on ^$\hat{\mathbb{P}}(S^{0}=0\mid G=g)$ and ^$\hat{\mathbb{P}}(S^{0}=1\mid G=g)$, respectively; (b) Estimated ATEs of $A$ on $Y$, and scatter plots of ^$\hat{\mathbb{P}}(Y^{1}=1\mid G=g)$ on ^$\hat{\mathbb{P}}(Y^{0}=0\mid G=g)$ and ^$\hat{\mathbb{P}}(Y^{0}=1\mid G=g)$, respectively;
By employing the proposed method outlined in Sections 34, and under the key Assumption 2 (i.e., $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$), we can estimate the state transition probabilities from $Y^{0}$ to $Y^{1}$, i.e., $\mathbb{P}(Y^{1}\mid Y^{0})$. This allows us to subsequently estimate the joint distributions $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for $g\in{\mathcal{G}}$. Similarly, if we assume $S^{1}\perp\!\!\!\perp G\mid S^{0}$ (referred to as Assumption 2 hereafter), then we can apply the same methods to estimate the transition probabilities $\mathbb{P}(S^{1}\mid S^{0})$ and the joint distributions $\mathbb{P}(S^{0},S^{1}\mid G=g)$ for $g\in{\mathcal{G}}$.
The key identification assumptions $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$ and $S^{1}\perp\!\!\!\perp G\mid S^{0}$ imply a linear relationship between $\mathbb{P}(Y^{1}=1\mid G=g)$ on $\mathbb{P}(Y^{0}=1\mid G=g)$ and $\mathbb{P}(Y^{0}=0\mid G=g))$, as well as between $\mathbb{P}(S^{1}=1\mid G=g)$ on $\mathbb{P}(S^{0}=1\mid G=g)$ and $\mathbb{P}(S^{0}=0\mid G=g))$. We illustrate these possible linear relationships through scatter plots, where probabilities are replaced with observed frequencies, as shown in Figure 1. The scatter plots suggest that the linear relationships likely hold.
Table 4: State transition probabilities and tests for Assumptions 2 and 2.
Parameters Estimate (ESE) 95% CI $J$-Statistic $p$-value null hypothesis $H_{0}$
$\mathbb{P}(S^{1}=1\mid S^{0}=0)$ 0.379 (0.107) (0.170, 0.590) 4.190 0.840 $S^{1}\perp\!\!\!\perp G\mid S^{0}$
$\mathbb{P}(S^{1}=1\mid S^{0}=1)$ 0.896 (0.049) (0.800, 0.992)
$\mathbb{P}(S^{1}=0\mid S^{0}=0)$ 0.621 (0.107) (0.411, 0.830)
$\mathbb{P}(S^{1}=0\mid S^{0}=1)$ 0.104 (0.488) (0.008, 0.200)
$\mathbb{P}(Y^{1}=1\mid Y^{0}=0)$ 0.275 (0.101) (0.062, 0.488) 8.793 0.360 $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$
$\mathbb{P}(Y^{1}=1\mid Y^{0}=1)$ 0.946 (0.051) (0.846, 1.044)
$\mathbb{P}(Y^{1}=0\mid Y^{0}=0)$ 0.725 (0.101) (0.512, 0.938)
$\mathbb{P}(Y^{1}=0\mid Y^{0}=1)$ 0.054 (0.051) (-0.045, 0.154)
Note: ESE is the estimated asymptotic standard error based on 500 bootstraps, 95% CI represents 95% confidence interval, and the $J$-Statistic and $p$-value correspond to the null hypothesis.
We then estimate the key invariant parameters $(\mathbb{P}(S^{1}=1\mid S^{0}=0),\mathbb{P}(S^{1}=1\mid S^{0}=1))$ and $(\mathbb{P}(Y^{1}=1\mid Y^{0}=0),\mathbb{P}(Y^{1}=1\mid Y^{0}=1))$ among trials, and the associated numerical results are presented in Table 4. Additionally, we run the overidentification tests for the key Assumptions 2 ($Y^{1}\perp\!\!\!\perp G\mid Y^{0}$) and 2 ($S^{1}\perp\!\!\!\perp G\mid S^{0}$). The corresponding $p$-value are 0.860 and 0.360, respectively, both of which are well above 0.100. Thus, we cannot reject these two null hypotheses.
We further estimate the joint distributions $\mathbb{P}(S^{0}=1,S^{1}=0\mid G=g)$ and $\mathbb{P}(Y^{0}=1,Y^{1}=0\mid G=g)$ for $g\in{\mathcal{G}}$. Figure 2 displays the point estimates along with the corresponding pointwise 95% confidence intervals (CIs). The lower bounds of the 95% CIs for ^$\hat{\mathbb{P}}(S^{0}=1,S^{1}=0\mid G=g)$ for all trials are strictly positive, suggesting that the monotonicity condition (Assumption 7, $S^{1}\geq S^{0}$) may not hold. In contrast, the 95% CIs for ^$\hat{\mathbb{P}}(Y^{0}=1,Y^{1}=0\mid G=g)$ cover 0 in all trials. The other estimated values of $\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$ and $\mathbb{P}(Y^{0}=a,Y^{1}=b\mid G=g)$ for $a,b\in\{0,1\}$ are provided in Supplementary Material S7.
Figure 2: Estimated joint distributions of $\mathbb{P}(S^{0}=1,S^{1}=0\mid G=g)$ and $\mathbb{P}(Y^{0}=1,Y^{1}=0\mid G=g)$ for all $g\in{\mathcal{G}}$.
7.2 Evaluation of Principal Surrogate
To evaluate the surrogate, we estimate the principal stratification average causal effects $\text{PSACE}_{ab|g}$ for $g\in{\mathcal{G}}$. Based on the methods proposed in Section 5, we adopt the following four ways to estimate $\text{PSACE}_{ab|g}$ under different sets of assumptions.
  • Method 1: Based on Theorem 5, relying on Assumptions 4, 6, and 7 ($S^{1}\geq S^{0}$) like Jiang et al. (2016). See Section 5.3 for estimation details.
  • Method 2: Based on Theorem 4, relying on Assumptions 4, 5, $Y^{1}\geq Y^{0}$, and $S^{1}\geq S^{0}$. See Supplementary Material S2 for details.
  • Method 3: Based on Corollary 2, relying on Assumptions 4, 5, and additional partial monotonicity $Y^{1}\geq Y^{0}$ (see Supplementary Material S4). We include this method for comparison, primarily because $S^{1}\geq S^{0}$ may not hold from Figure 2.
  • Method 4: Based on Corollary 2, relying on Assumptions 4 and 5 but no monotonicity. See Supplementary Material S2 for details.
Figure 3: (a) Estimated joint distribution $\mathbb{P}(S^{0},S^{1}\mid G=g)$ for all $g\in{\mathcal{G}}$; (b) Estimated joint distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for all $g\in{\mathcal{G}}$. Methods 1 and 2 assume monotonicity $S^{1}\geq S^{0}$ so $\text{PSACE}_{10\mid g}$ is undefined.
Figure 3 presents the point estimates and the associated 95% CIs of $\text{PSACE}_{ab|g}$ across trials, with the results displayed in four columns corresponding to Methods 1, 2, 3, and 4, respectively. Note that when $S^{1}\geq S^{0}$ is assumed (Methods 1 and 2), $\text{PSACE}_{10|g}$ is undefined so we set it as null. Additionally, under Assumption 6 (Method 1), for fixed $a,b\in\{0,1\}$, $\text{PSACE}_{ab|g}$ remains invariant across trials.
From Figure 3, we have the following observations: (1) From the third row, one can see that the potential surrogate may not satisfy the full causal sufficiency, as all 95% CIs of $\text{PSACE}_{10|g}$ cover zero; (2) From the second row, the point estimates of $\text{PSACE}_{01|g}$ are significantly greater than zero, suggesting that causal sufficiency partially holds; (3) From the first and fourth rows, the potential surrogate may satisfy the causal necessity, as all 95% CIs of $\text{PSACE}_{00|g}$ and $\text{PSACE}_{11|g}$ include zero; (4) Considering that $S^{1}\geq S^{0}$ may not hold based on Figure 2, the results from Methods 3 and 4 appear more plausible.
Conclusion
In this work, we propose a novel framework that leverages multiple experimental datasets to identify and estimate the joint distribution of potential outcomes. First, we established the identifiability of joint distributions for binary and categorical outcomes under the assumption of transportability of state transition probabilities (Assumption 2) and a rank condition (Conditions 1 and 2). Second, we introduce a simple least-squares-based method for estimating the joint distribution of potential outcomes. Finally, we extended our framework to identify and estimate principal stratification causal effects. These extend and complement existing methods in causal inference that integrate information from multiple datasets.