### v2

# The Promises of Multiple Experiments: Identifying Joint Distribution of Potential Outcomes

Abstract
Typical causal effects are defined based on the marginal distribution of potential outcomes. However, many real-world applications require causal estimands involving the joint distribution of potential outcomes to enable more nuanced treatment evaluation and selection. In this article, we propose a novel framework for identifying and estimating the joint distribution of potential outcomes using multiple experimental datasets. We introduce the assumption of transportability of state transition probabilities for potential outcomes across datasets and establish the identification of the joint distribution under this assumption, along with a regular full-column rank condition. The key identification assumptions are testable in an overidentified setting and are analogous to those in the context of instrumental variables, with the dataset indicator serving as “instrument”. Moreover, we propose an easy-to-use least-squares-based estimator for the joint distribution of potential outcomes in each dataset, proving its consistency and asymptotic normality. We further extend the proposed framework to identify and estimate principal causal effects. We empirically demonstrate the proposed framework by conducting extensive simulations and applying it to evaluate the surrogate endpoint in a real-world application.
Keywords: Causal Inference, Data Fusion, Principal Stratification, Surrogacy Evaluation
Introduction
Estimating the causal effect of a treatment variable on an outcome is a fundamental scientific problem, central to many fields such as social and biomedical sciences (Imbens and Rubin 2015; Pearl and Mackenzie 2018; Hernán and Robins 2020; Rosenbaum 2020). Typical causal effects like average treatment effects are defined in terms of the marginal distribution of potential outcomes. However, in many real-world applications, causal estimands involving the joint distribution of potential outcomes are required to enable more nuanced treatment evaluation and selection. Examples include the probability of causation (Pearl 1999; Dawid and Musio 2022; Lu et al. 2023), the effect of persuasion (Jun and Lee 2023, 2024), treatment benefit rates and treatment harm rates (Shen et al. 2013; Yin et al. 2018b; Kallus 2022a, b; Li et al. 2023a; Wu et al. 2024a), and so on.
Identifying the joint distribution of potential outcomes from a single dataset presents a significant challenge. This difficulty arises from the fact that only one potential outcome is observed for each individual, which is known as the fundamental problem in causal inference (Holland 1986). Consequently, the joint distribution of potential outcomes is generally unidentifiable, even in randomized controlled trials (Pearl et al. 2016). Instead of identifying the joint distribution, many studies focus on deriving bounds for causal estimands involving the joint distribution (e.g., Tian and Pearl 2000; Zhang et al. 2013; Yin et al. 2018a; Li et al. 2022; Kallus 2022b).
In this article, we propose a novel framework for identifying and estimating the joint distribution of potential outcomes using multiple experimental datasets. Specifically, we make the following three contributions. First, our primary contribution is to study identification of the joint distribution of potential outcomes, which in turn defines many important causal quantities. For binary outcomes, we introduce the assumption of transportability of state transition probabilities, namely that the conditional law of the treated potential outcome $Y^{1}$ given the untreated potential outcome $Y^{0}$ is invariant across trials, together with a full-column rank condition requiring sufficient cross-trial variation in the untreated potential outcome distribution. Under these conditions, we establish identification of the joint distribution of $(Y^{1},Y^{0})$ for each trial. We further extend this result to general categorical outcomes. In an overidentified setting, the same identifying restrictions also generate observable implications that can be assessed empirically. The resulting identification strategy is analogous to instrumental-variable reasoning, with the dataset indicator serving as a source of identifying variation.
Second, we propose a simple least-squares-based framework for estimation and inference. The key idea is that the observed treated-arm outcome distribution in each trial can be written as a linear function of the untreated potential outcome distribution and the common transition parameters. Building on this representation, we derive least-squares estimators for the invariant transition probabilities and the induced joint distributions, establish their consistency and asymptotic normality, and develop an overidentification test for the observable restrictions implied by the transportability assumption. The resulting procedure is easy to implement and only relies on arm-specific summary statistics, which may be useful when individual-level data are not directly available.
Third, when an additional binary post-treatment variable is observed before the final outcome, we extend the framework to principal stratification. We first develop new transportability-based identification routes for the joint distribution of the potential surrogates and outcomes, with and without additional monotonicity restrictions. These routes can accommodate trial-specific principal stratification average causal effects (PSACEs). We then show how these results complement the monotonicity-based route of Jiang et al. (2016). In this sense, PSACEs are treated here as an important application of a broader joint-distribution framework.
Throughout the main text, we focus on a covariate-free, nonparametric multi-trial identification framework. We make this choice deliberately rather than incidentally, and it aligns with our empirical application, in which no individual-level baseline covariates are available. In Supplementary Material S2, we discuss how to extend the proposed framework to accommodate baseline covariates.
The idea of combining two or more datasets to identify causal effects and enhance estimation efficiency has garnered significant attention in the field of causal inference (Colnet et al. 2023; Degtiar and Rose 2023; Wu et al. 2024b; Athey et al. 2019; Yang and Ding 2020; Imbens et al. 2024; Kallus and Mao 2024; Hu et al. 2025; Hünermund and Bareinboim 2025). However, most existing methods consider causal effects involving the marginal distribution of potential outcomes. In contrast, we focus on identifying and estimating causal estimands that involve the joint distribution of potential outcomes, thus extending and complementing this class of methodologies.
The rest of this article is organized as follows. In Section 2, we outline the basic setup. In Section 3, we present the identification assumptions and establish the identification for the joint distribution of potential outcomes. In Section 4, we propose the least-squares-based method for estimating the transition probabilities and present a method for testing the key assumption of transportability of state transition probabilities. Section 5 extends the proposed method to identify and estimate PSACEs. Section 6 evaluates the finite-sample performance of the proposed method through extensive simulations. Section 7 demonstrates the proposed methods in a real-world application, and Section 8 concludes the paper.
Setup
Let $A\in\mathcal{A}=\{0,1\}$ represent a binary treatment, where $A=1$ indicates the treatment condition and $A=0$ indicates the control condition. Let $Y\in\mathcal{Y}$ be the outcome of interest. To define causal estimands, we adopt the potential outcome framework (Neyman 1923; Rubin 1974) and denote $Y^{1}$ and $Y^{0}$ as the potential outcomes under treatment and control, respectively. We make the stable unit treatment value assumption (i.e., no multiple treatment versions and no interference across units), which ensures the well-definedness of potential outcome $Y^{a}$ (Imbens and Rubin 2015). Accordingly, the observed outcome is linked to the potential outcomes through $Y=(1-A)Y^{0}+AY^{1}$.
Suppose we have access to data from individuals in a collection of $m$ trials, indexed by ${\mathcal{G}}=\{1,...,m\}$. For each trial $g\in{\mathcal{G}}$, the observed data consist of realizations of independent random tuples $\{(G_{i}=g,X_{i},A_{i},Y_{i}),i=1,...,n_{g}\}$, where $n_{g}$ denotes the number of individuals in trial $g$, and we let $n=\sum_{g=1}^{m}n_{g}$ represent the total sample size. We adopt a non-nested design (Dahabreh et al. 2021) in which the observed samples from different trials are independent. Without loss of generality, under a superpopulation model $\mathbb{P}$, we assume that the observed data in trial $g$ form an i.i.d. sample from $\mathbb{P}(\cdot\mid G=g)$, which implies that all observed data are obtained by stratified sampling from a population that is stratified by $G$.
In this paper, our primary objective is to identify and estimate the joint distribution of potential outcomes $Y^{0},Y^{1}$ from the data of multiple trials. Given that there might be potential distribution discrepancies across trials, we are particularly interested in the distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for each trial $g\in\mathcal{G}$. Notably, this task diverges from and is more challenging than what most existing data-combination-based causal inference methods can handle. The latter mainly focuses on estimating causal effects defined in terms of the marginal distribution of potential outcomes (e.g., Rosenman et al. 2023; Colnet et al. 2023; Wu et al. 2024b).
We mainly consider the case of a binary outcome $Y\in\{0,1\}$ and number of trials $m\geq 3$ (i.e., at least three trials). This case is of particular interest because it leads to overidentification and offers an opportunity to test the key identification assumption. Additionally, for identifying the principal causal effects, we require $m\geq 4$ due to the extra complexity introduced by principal stratification and relax this to $m\geq 2$ under additional monotonicity conditions, see Section 5 for details.
Nonparametric Identification
In this section, we establish the nonparametric identification of the joint distribution of potential outcomes. We do not consider any covariates in the main text and defer the discussion of covariates to Supplementary Material S2.
3.1 Basic Assumptions
We begin by presenting the basic assumptions required to identify $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$.
Assumption 1 (Within-trial Randomization).
For all $g\in{\mathcal{G}}=\{1,...,m\}$, (i) $A\perp\!\!\!\perp(Y^{0},Y^{1})\mid G=g$, and (ii) $0<\mathbb{P}(A=1\mid G=g)<1$.
Assumption 1 means that the treatment $A$ is randomized within each trial and both treatment arms are assigned with a positive probability. This assumption trivially holds for data from randomized trials, and it identifies the marginal potential outcome distributions within each trial. However, it is insufficient to identify the joint distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$, because the observed data only provide information about single-world marginals rather than the cross-world dependence between $Y^{0}$ and $Y^{1}$. Therefore, if our goal is to identify the joint distribution of potential outcomes, some cross-world structure is unavoidable. Moreover, because we aim to combine information across multiple trials, some cross-trial structure is also needed; otherwise there is no basis for borrowing information across trials. We therefore introduce the following invariance assumption.
Assumption 2 (Transportability of state transition probabilities).
$Y^{1}\perp\!\!\!\perp G\mid Y^{0}$.
Assumption 2 requires that the cross-world conditional distribution of $Y^{1}$ given $Y^{0}$ is invariant across trials, i.e., $\mathbb{P}(Y^{1}=1\mid Y^{0},G=1)=\mathbb{P}(Y^{1}=1\mid Y^{0},G=2)=\cdots=\mathbb{P}(Y^{1}=1\mid Y^{0},G=m)$. This means that the transition probabilities from the untreated state $Y^{0}$ to the treated state $Y^{1}$ are the same across trials.
A useful way to interpret Assumption 2 is through untreated prognosis. In many applications, $Y^{0}$ can be viewed as a latent summary of an individual’s untreated prognosis, or natural outcome tendency under control, whereas $Y^{1}$ represents the outcome after treatment acts on this prognosis. Under this interpretation, Assumption 2 states that, once untreated prognosis is fixed, the treatment-response mechanism is stable across trials, even though the marginal distribution of untreated prognosis may differ across trials. In Section 7, our application focuses on the Adjuvant Colon Clinical Trials (ACCTs), where $Y^{1}$ and $Y^{0}$ refer to cancer survival with and without treatment, respectively. In this setting, Assumption 2 means that, given the untreated prognosis summarized by $Y^{0}$, the survival response under fluorouracil-based chemotherapy is approximately the same across trials. The same logic may also arise in other contexts, where populations may differ in baseline natural outcome tendency across sites while the incremental response mechanism conditional on the natural tendency is comparatively stable.
Assumption 2 can also be understood from the perspective of effect modifiers. In particular, it implies that any effect modification by $G$ operates through the untreated potential outcome $Y^{0}$. That is, once we condition on $Y^{0}$, the trial indicator $G$ does not further modify the treatment response. Thus, Assumption 2 still allows substantial heterogeneity across trials in the marginal distributions of potential outcomes and in average treatment effects, but it requires the conditional transition law from $Y^{0}$ to $Y^{1}$ to remain stable. This is weaker than the full transportability assumption $G\perp\!\!\!\perp(Y^{0},Y^{1})$ that requires identical effects across trials. This stronger assumption is ruled out in our analysis of the ACCT data, as shown in Figure 1 in Section 7, where the average treatment effects vary across trials.
Assumption 2 has both strengths and limitations. On the one hand, it has a clear scientific interpretation as a stable transition mechanism from untreated prognosis to treated outcome, while allowing substantial cross-trial heterogeneity through differences in $\mathbb{P}(Y^{0}\mid G=g)$. On the other hand, it is a substantive cross-world assumption and is not guaranteed by randomized design alone. It may fail when treatment versions, adherence, supportive care, endpoint definitions, or latent effect modifiers differ across trials in ways that still affect treatment response after conditioning on $Y^{0}$.
Importantly, Assumption 2 has observable implications in an overidentified setting, namely when the number of trials is strictly larger than the number of outcome categories. For a binary outcome, this means $m\geq 3$. In this case, the same transition probabilities must simultaneously explain the observed arm-specific outcome distributions across all trials. In Section 4.2, we develop a chi-squared overidentification test for the restrictions implied by Assumption 2. Such a test assesses whether the observable implications of Assumption 2 are compatible with the data. A rejection indicates that these implied restrictions are violated by the observed data, whereas a failure to reject should only be interpreted as a lack of evidence against Assumption 2. We also provide visual diagnostics in Section 7.
3.2 Identifiability for Joint Distribution of Potential Outcomes
An important implication of Assumptions 12 is that the joint distributions of potential outcomes are identifiable from the data of multiple trials. Assumption 2 imposes invariant potential outcome state transition probabilities across trials:
$\pi_{b|a}:=\mathbb{P}(Y^{1}=b\mid Y^{0}=a)=\mathbb{P}(Y^{1}=b\mid Y^{0}=a,G=g),\quad a,b=0,1;g\in\mathcal{G}.$
By law of total probability, the following holds for all $g\in\mathcal{G}$:
$\mathbb{P}(Y^{1}=1\mid G=g)=\mathbb{P}(Y^{0}=0\mid G=g)\cdot\pi_{1|0}+\mathbb{P}(Y^{0}=1\mid G=g)\cdot\pi_{1|1},$ (1)
where $\mathbb{P}(Y^{1}=1\mid G=g)$ and $\mathbb{P}(Y^{0}=a\mid G=g)$ for $a=0,1$ are all identifiable under Assumption 1. Therefore, identifying $\pi_{1|0}$ and $\pi_{1|1}$ amounts to solving the system of $m$ linear equations in (1). The solution is unique, and hence point identification holds, under the following full-rank condition on the coefficient matrix in (1).
Condition 1 (Full-column rank).
The matrix $(\mathbb{P}(Y^{0}=0\mid G=g),\mathbb{P}(Y^{0}=1\mid G=g)_{m\times 2}$ has a full-column rank (i.e., rank $2$).
For a binary outcome, Condition 1 is equivalent to requiring that the untreated outcome distribution differs across at least two trials, namely that the untreated base rate $\mathbb{P}(Y^{0}=1\mid G=g)$ is not the same for all $g$. Thus, for point identification, the condition does not require the untreated base rates to be very different; any nonzero cross-trial variation is sufficient in principle. At the same time, larger variation generally improves numerical stability and finite-sample precision, whereas near collinearity in the design matrix can make estimation more variable.
At first glance, Assumption 2 and Condition 1 may appear to pull in different directions. Assumption 2 requires the conditional law $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ to be invariant across trials, whereas Condition 1 requires the marginal untreated outcome distribution $\mathbb{P}(Y^{0}\mid G=g)$ to vary across trials. We do not view these two requirements as being in any necessary tension. Rather, they concern different aspects of the data-generating structure and play complementary roles in identification. Condition 1 provides the identifying leverage through cross-trial variation in untreated prognosis, whereas Assumption 2 imposes a common conditional transition law that links these heterogeneous trial populations. Different trials may therefore have different untreated base rates, and even different treated marginal risks, while still sharing the same conditional response mechanism once $Y^{0}$ is fixed. The key substantive issue is whether the factors driving cross-trial variation in $Y^{0}$ are adequately summarized by $Y^{0}$ itself. If they are, then marginal heterogeneity in $Y^{0}$ and conditional invariance of $Y^{1}\mid Y^{0}$ are compatible; if they are not, then Assumption 2 becomes less plausible.
Condition 1 is satisfied only if $m\geq 2$ for a binary outcome. If we only have access to data from a single trial, one might imagine splitting the data into two independent subsamples to create two pseudo-trials. However, this would not generate new identifying variation: although the two subsamples would be independent, their untreated outcome distributions would be the same at the population level, so Condition 1 would fail.
Theorem 1 (Binary outcome).
Under Assumptions 12 and Condition 1, the distributions $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ and $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ for $g\in{\mathcal{G}}$ are identifiable.
Theorem 1 shows that the conditional distribution $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ and the joint distribution $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ are identified for each trial $g\in{\mathcal{G}}$. Intuitively, Assumption 2 implies that the same transition probabilities $(\pi_{1|0},\pi_{1|1})$ must fit all trials simultaneously, and Condition 1 ensures that the resulting linear system has a unique solution. Once $\mathbb{P}(Y^{1}\mid Y^{0},G=g)$ is identified, the joint distribution follows immediately from $\mathbb{P}(Y^{1}=b,Y^{0}=a\mid G=g)=\mathbb{P}(Y^{1}=b\mid Y^{0}=a,G=g)\mathbb{P}(Y^{0}=a\mid G=g)$ for $a,b\in\{0,1\}$.
The identification Assumption 2 and Condition 1 resemble the standard assumptions in the context of instrumental variable (IV) analysis (Angrist et al. 1996; Imbens 2004). We may view the dataset indicator $G$ as an “IV”, the potential outcome $Y^{0}$ as the “treatment”, and the potential outcome $Y^{1}$ as the “outcome”. Then Assumption 2 corresponds to the IV exclusion restriction assumption, and Condition 1 corresponds to the full rank assumption of IV on treatment (or IV relevance), which requires a strong enough association between $G$ and $Y^{0}$.
We note that one may instead impose the symmetric assumption $Y^{0}\perp\!\!\!\perp G\mid Y^{1}$. Then, by the same argument as above, $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ remains identifiable provided that the marginal distribution of $Y^{1}$ varies sufficiently across trials, namely that the matrix $\bigl(\mathbb{P}(Y^{1}=0\mid G=g),,\mathbb{P}(Y^{1}=1\mid G=g)\bigr)_{m\times 2}$ has full column rank. We focus on Assumption 2 because it is more natural scientifically: in many applications, $Y^{0}$ represents untreated prognosis, and $\mathbb{P}(Y^{1}\mid Y^{0},G)$ then describes how treatment acts on that prognosis. By contrast, $\mathbb{P}(Y^{0}\mid Y^{1},G)$ has backward-looking interpretation. Thus, although the two formulations are mathematically symmetric, Assumption 2 provides the more natural primary formulation.
Following Theorem 1, we can identify various causal estimands involving the joint distribution of potential outcomes. Below are some examples.
Example 1 (Probability of causation).
Causal inference involves not only evaluating the effects of causes but also deducing the causes of given effects (Dawid and Musio 2022), where the latter is also referred to as attribution analysis (Pearl 2009; Pearl et al. 2016; Pearl and Mackenzie 2018). The probability of sufficient causation (PS) and the probability of necessary causation (PN) are two standard quantities for attribution analysis. They are defined by $\textup{PS}(A\Rightarrow Y)=\mathbb{P}(Y^{1}=1\mid A=0,Y=0)$ and $\textup{PN}(A\Rightarrow Y)=\mathbb{P}(Y^{0}=0\mid A=1,Y=1)$ respectively. For example, the quantity $\textup{PN}(A\Rightarrow Y)$ can be written as $\mathbb{P}(Y^{0}=0,Y^{1}=1\mid A=1)/\mathbb{P}(Y^{1}=1\mid A=1)$, where the denominator is an identifiable quantity and the numerator equals to $\mathbb{P}(Y^{0}=0,Y^{1}=1)$ for randomized treatment assignment.
Example 2 (Effect of persuasion).
Let $A\in\{0,1\}$ be a binary indicator for an individual’s exposure to certain persuasive information, and let $Y$ be a binary indicator representing the individual’s behavior, where $Y=0$ denotes a negative response to the information and $Y=1$ denotes a positive response. Jun and Lee (2023) defined the persuasion rate as $\mathbb{P}(Y^{1}=1\mid Y^{0}=0)$, which quantifies the proportion of individuals in the subpopulation with $Y^{0}=0$ who would change from negative to positive behavior as a result of the exposure to persuasive information.
Example 3 (Treatment benefit and harm rates).
Denote $Y=1$ as a favorable outcome (e.g., survival) and $Y=0$ as an unfavorable outcome (e.g., death). The average treatment harm rate (Shen et al. 2013; Wu et al. 2024a) is defined as $\textup{THR}=\mathbb{P}(Y^{0}=1,Y^{1}=0),$ which quantifies the percentage of individuals experiencing worse outcomes under treatment than under control. Similarly, the average treatment benefit rate is defined as $\textup{TBR}=\mathbb{P}(Y^{0}=0,Y^{1}=1).$
In some scenarios, the target data may contain only control units (Li et al. 2023b). For example, this could arise if the target population has only received an old drug ($A=0$) and researchers want to assess the efficacy of a new drug ($A=1$) unavailable in this population. Our identification result can be readily generalized to such a control-only target population.
Example 4 (Generalization to a control-only target population).
Consider a simple random sample from the target population with only control units (denoted by $G=0$). In the control-only dataset, we can straightforwardly identify $\mathbb{P}(Y^{0}=a\mid G=0)$. Moreover,applying Theorem 1 allows us to identify $\pi_{b\mid a}$ for $a,b=0,1$ from the $m$ trials. Taken together, these results yield identification of the joint distribution in the control-only target population: $\mathbb{P}(Y^{0}=a,Y^{1}=b\mid G=0)=\pi_{b|a}\cdot\mathbb{P}(Y^{0}=a\mid G=0)$.
3.3 Extension to Categorical Outcomes
The identifiability results can be also extended to general categorical outcomes. Specifically, for a categorical outcome with cardinality $k$, let $\mathbb{P}(Y^{0}\mid G)=(\mathbb{P}({\color[rgb]{0,0,0}Y^{0}}=i\mid G=g))_{m\times k}$ be a $m\times k$ matrix whose $(g,i)$-th element is $\mathbb{P}({\color[rgb]{0,0,0}Y^{0}}=i\mid G=g)$.
Condition 2.
The matrix $\mathbb{P}(Y^{0}\mid G)$ has full-column rank.
Similar to Condition 1, Condition 2 holds only when $m\geq k$.
Theorem 2 (Categorical outcome).
Let $Y$ be a categorical outcome with $k$ possible values. Under Assumptions 12 and Condition 2, the joint distributions $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$ for $g\in{\mathcal{G}}$ are identifiable.
Theorem 2 extends the result of Theorem 1. We also discuss extensions to more general treatment settings; see Supplementary Material S8 for details. After establishing identifiability, we then consider the estimation of the joint distributions and the testing of the key Assumption 2.
Estimation and Inference
In this section, we first present an estimator of $\pi_{b|a}$ for $a,b=0,1$, which are key invariant parameters for identifying the joint distribution $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$. Then, we provide a method to test Assumption 2 when $m=|\mathcal{G}|\geq 3$.
4.1 Estimation
We now present a simple least-squares-based estimation method for $\pi_{b|a}$, $a,b\in\{0,1\}$. Under Assumptions 12, (1) gives the following for $g\in\mathcal{G}$:
$\mathbb{P}(Y=1\mid G=g,A=1)=\pi_{1|0}\mathbb{P}(Y=0\mid G=g,A=0)+\pi_{1|1}\mathbb{P}(Y=1\mid G=g,A=0)$ (2)
When $m=2$, the quantities $\pi_{b|a}$ for $a,b\in\{0,1\}$ are just-identified as the number of parameters to be identified equals the number of equations. For example, when $m=2$, solving the two equations in (2) identifies $\pi_{1|0}$ and $\pi_{1|1}$. Then $\pi_{0|0}$ and $\pi_{0|1}$ are identified through $\pi_{0|0}=1-\pi_{1|0}$ and $\pi_{0|1}=1-\pi_{1|1}$, respectively. In contrast, if $m>2$, (2) includes more equations than parameters. In such a case, the parameters $\pi_{b|a}$ for $a,b\in\{0,1\}$ are overidentified.
In both just-identified and overidentified cases, we can estimate $\theta:=(\pi_{1|0},\pi_{1|1})$ via a linear least-squares estimator. Formally, we let $\tilde{Y}_{g}=\mathbb{P}(Y=1\mid G=g,A=1),$ $\tilde{X}_{1g}=\mathbb{P}(Y=0\mid G=g,A=0),$ and $\tilde{X}_{2g}=\mathbb{P}(Y=1\mid G=g,A=0)$, and denote $\tilde{X}_{g}=(\tilde{X}_{1g},\tilde{X}_{2g})^{\intercal}$. By (2), we have $\tilde{Y}_{g}=\tilde{X}_{g}^{\top}\theta$ for $g\in\mathcal{G}$. Under the full rank condition (Condition 1), we can further write $\theta$ as
$\theta=\left(\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{Y}_{g}.$
Although the conditional probabilities $\tilde{X}_{g}$ and $\tilde{Y}_{g}$ are unknown, they can be easily estimated by the corresponding sample frequencies, e.g., $\hat{Y}_{g}=\sum_{i=1}^{n}\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)\big/\sum_{i=1}^{n}\mathbb{I}(G_{i}=g,A_{i}=1)$. Given estimators $\hat{X}_{g}$ and $\hat{Y}_{g}$, the resulting estimator for $\hat{\theta}$ is
$\hat{\theta}=\left(\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g}.$
This amounts to the ordinary least squares (OLS) coefficient estimator with $\hat{Y}_{g}$ as the response and $\hat{X}_{g}$ as the covariates. It is noteworthy that while $\hat{\theta}$ resembles the OLS estimator in a linear model, it differs in that the “effective sample size” $m$ is fixed. A potential strength of this estimator is that it does not need the individual-level data. Instead, it only involves some sample summary statistics $\hat{X}_{g}$ and $\hat{Y}_{g}$. This may be helpful for privacy protection (Han et al. 2023).
Next, we show that estimator $\hat{\theta}$ is consistent and asymptotically normal. In our theoretical analysis, for each trial $g$, the sample size $n_{g}\to\infty$, and the proportion $n_{g}/n$ converges to a strictly positive constant in $(0,1)$.
Theorem 3.
The estimator $\hat{\theta}$ is consistent and asymptotically normal, satisfying $\sqrt{n}(\hat{\theta}-\theta)\xrightarrow{d}N(0,\Sigma),$ where $\Sigma=C^{-1}VC^{-1}$, $C=m^{-1}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}$, and
$V=$ $\frac{1}{m^{2}}\sum_{g=1}^{m}\mathrm{Var}\Biggl\{\tilde{Y}_{g}\begin{pmatrix}\dfrac{\mathbb{I}(Y=0,G=g,A=0)-\mathbb{P}(Y=0,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\\ \dfrac{\mathbb{I}(Y=1,G=g,A=0)-\mathbb{P}(Y=1,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\end{pmatrix}$
$+\dfrac{\mathbb{I}(Y=1,G=g,A=1)-\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(G=g,A=1)}\tilde{X}_{g}\Biggr\}.$
From Theorem 3, we could use the bootstrap method to estimate the asymptotic variance of the proposed estimator, as detailed in Supplementary Material S4.3.
4.2 Testing Assumption 2 under Overidentification
If the parameter $\theta$ is overidentified, we can further test for the observable implications of Assumption 2—the key identification assumption. Specifically, given that Assumption 1 holds, Assumption 2 leads to the null hypothesis $H_{0}:\tilde{Y}_{g}=\tilde{X}_{g}^{\intercal}\theta,\text{ for all }g=1,...,m,$ that is, there is a linear relationship between $\tilde{Y}_{g}$ and $\tilde{X}_{g}$ for all $g\in{\mathcal{G}}$.
[…]
$J=n\;\hat{\boldsymbol{\epsilon}}^{\intercal}\,\boldsymbol{\Sigma}_{*}^{+}\,\hat{\boldsymbol{\epsilon}}$,
where $\boldsymbol{\Sigma}_{*}^{+}$ denotes the Moore-Penrose pseudo-inverse of $\boldsymbol{\Sigma}_{*}$, and $\boldsymbol{\Sigma}_{*}$ is the asymptotic covariance matrix of $\sqrt{n}\boldsymbol{\hat{\epsilon}}$, which can be estimated via the bootstrap. In Supplementary Material S1.8, we show that the test statistic is asymptotically chi-squared, i.e., $J\xrightarrow{d}\chi^{2}_{m-2}$, where $m$ is the number of trials and $2$ is the number of outcome categories. In addition, we provide the procedure for implementing the hypothesis test in Supplementary Material S3. Intuitively, if Assumption 2 is significantly violated, the $J$-statistic will be far from 0, leading to a rejection of the null hypothesis $H_{0}$. However, it should be noted that the failure to reject $H_{0}$ does not necessarily imply that Assumption 2 holds, because this may simply result from the test’s lack of power or insufficient evidence to reject Assumption 2. Nevertheless, the test could be useful for detecting the violation of Assumption 2.
We emphasize that the overidentification test is used here as a diagnostic tool for the observable implications of Assumption 2, not as a pretest that determines whether the estimation and inference result is reported. The asymptotic results are therefore derived under the maintained identifying assumptions and do not condition on the test outcome. If the test were instead used as a pretest for model selection, additional analysis would be needed to address the resulting post-selection inference, which is beyond the scope of this paper.
The test on Assumption 2 is analogous to the test of exclusion restriction assumption in the IV setting, such as Windmeijer (2019); Kiviet (2020); Carrasco and Doukali (2022), and overidentified testing method in the GMM framework (Newey 1985; Newey and McFadden 1994).
Extension to Principal Causal Effects
In this section, we extend the framework developed in Sections 34 to the setting with an additional binary post-treatment variable $S$, and study the identification of principal stratification average causal effects (PSACEs). As before, we focus on the binary outcome setting. In addition to variables $(G,A,Y)$, we now observe a binary post-treatment variable $S$ (e.g., a surrogate endpoint) measured before the outcome $Y$. Let $S^{0}$ and $S^{1}$ denote the potential outcomes of $S$ under treatment arms 0 and $1$, respectively. Our goal is to identify $\mathbb{P}(Y^{a},S^{0},S^{1}\mid G=g)$ for $g\in{\mathcal{G}}$ and $a=0,1$. Once these joint distributions are identified, we can identify the PSACEs:
$\text{PSACE}_{ab|g}=\mathbb{E}[Y^{1}-Y^{0}\mid S^{0}=a,S^{1}=b,G=g],\quad a=0,1;b=0,1.$
These quantities have many applications, such as noncompliance (Imbens and Angrist 1994), truncation by death (Rubin 2006), and surrogate endpoint evaluation (Jiang et al. 2016). See Frangakis and Rubin (2002) for more examples.
We consider two conceptually distinct identification routes in this setting. The first route, developed in Section 5.1, extends our transportability of state transition probabilities from Sections 34 to the pair of states $(S^{0},Y^{0})$ and $(S^{1},Y^{1})$. This route yields two versions: Corollary 1 without monotonicity, and Theorem 4 with additional monotonicity, where monotonicity weakens the rank requirement. The second route, discussed in Section 5.2, is the monotonicity-based route of Jiang et al. (2016). We focus on the first route and the new transportability-based results. Then we compare these two routes in detail.
5.1 Our Transportability-Based Route
Following the spirit of Sections 34, and in parallel to Assumptions 12 and Condition 1, we introduce the following assumptions and rank condition.
Assumption 3.
$A\perp\!\!\!\perp(S^{0},S^{1},Y^{0},Y^{1})\mid G=g$ and $0<\mathbb{P}\left(A=1\mid G=g\right)<1$ for all $g\in\mathcal{G}$.
Assumption 4.
$G\perp\!\!\!\perp(S^{1},Y^{1})\mid S^{0},Y^{0}$.
Condition 3.
The number of trials $m\geq 4$, and the matrix $(\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=g),\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=g))_{m\times 4}$ has full-column rank.
Assumption 3 is analogous to Assumption 1 and holds when all trials are randomized experiments. Assumption 4 is the principal-stratification analogue of Assumption 2: it imposes transportability of the state transition from the untreated state $(S^{0},Y^{0})$ to the treated state $(S^{1},Y^{1})$ across trials. In this setting, we may again regard $(S^{0},Y^{0})$ as a latent summary of untreated prognosis, so that Assumption 4 describes a stable transition mechanism across trials.
Corollary 1.
Under Assumptions 3, 4 and Condition 3, the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ for all $g\in\mathcal{G}$ are identifiable, and therefore $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ and $g\in\mathcal{G}$ are also identifiable.
Corollary 1 is the baseline transportability-based route in this section. It does not impose monotonicity, but requires the stronger four-column rank condition (Condition 3), and hence at least four trials. Essentially, we can view Corollary 1 as an instantiation of Theorem 2, treating $(Y^{1},S^{1})$ and $(Y^{0},S^{0})$ as two categorical variables, each with four possible values.
In addition, we can relax Condition 3 by additionally imposing monotonicity.
Condition 4.
(i) The matrix $(\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=g))_{m\times 2}$ has full column rank; (ii) the matrix $(\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g),\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=g))_{m\times 2}$ has full column rank.
Theorem 4.
If $Y^{1}\geq Y^{0}$ and $S^{1}\geq S^{0}$, then under Assumptions 3, 4 and Condition 4, the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ for $g\in\mathcal{G}$ are identifiable, and therefore $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ and $g\in\mathcal{G}$ are also identifiable.
Theorem 4 should be read as a monotonicity-strengthened version of Corollary 1. By imposing the additional monotonicity conditions $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$, it relaxes the identifying-variation requirement from Condition 3 to Condition 4, and hence from $m\geq 4$ to $m\geq 2$. In this sense, Corollary 1 and Theorem 4 are best viewed as two versions of the same transportability-based route: Corollary 1 is the no-monotonicity version with a stronger rank condition, whereas Theorem 4 is the monotonicity-strengthened version with a weaker rank condition.
In addition, under the identification assumptions in Corollary 1 and Theorem 4, we can also construct least-squares-based estimators for the invariant transition parameters and the induced PSACEs. To keep the main text concise, we defer these estimation details to Supplementary Material S4.
If only partial monotonicity holds (i.e., $S^{1}\geq S^{0}$ or $Y^{1}\geq Y^{0}$, but not both), then under Assumptions 3, 4 and Condition 4, only half of the quantities in $\{\mathbb{P}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g):a,b,c,d=0,1;\ g\in\mathcal{G}\}$ are identifiable (see Supplementary Material S5). Thus, partial monotonicity alone cannot replace Condition 3 by Condition 4 for the purpose of identifying the whole joint distribution. Nevertheless, partial monotonicity can still simplify estimation when Condition 3 holds; see Supplementary Material S5.3. Moreover, if we relax Assumption 4 to partial but not joint conditional independence $G\perp\!\!\!\perp S^{1}\mid(S^{0},Y^{0})$ and $G\perp\!\!\!\perp Y^{1}\mid(S^{0},Y^{0})$ in Corollary 1, then we can identify $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$ but not the full joint distribution needed to identify the PSACEs; see Supplementary Material S6 for details.
5.2 Comparison with Monotonicity-based Route in Jiang et al. (2016)
Our paper is closely related to Jiang et al. (2016), as both works leverage multiple trials to identify causal quantities that are not identifiable from a single trial. The two papers, however, have different targets. Jiang et al. (2016) focus on the PSACEs for surrogate endpoint evaluation. By contrast, our primary target is the joint distribution of potential outcomes itself, with PSACEs treated as one important extension. This broader target covers causal estimands beyond PSACEs (see the examples in Section 3.2).
A second key difference is in the cross-trial invariance assumption. Jiang et al. (2016) impose outcome homogeneity within principal strata, namely $Y^{a}\perp\!\!\!\perp G\mid(S^{0},S^{1})$ for $a=0,1$, which implies that the PSACEs are invariant across trials. In contrast, our framework is based on transportability of state transition probabilities. This allows the marginal distributions of $(S^{0},Y^{0})$ and, in general, the PSACEs themselves to vary across trials. Conceptually, Jiang et al. (2016) assume invariance of outcomes within principal strata, whereas we assume invariance of the transition mechanism from untreated states to treated states.
In addition to the cross-trial invariance assumption, Jiang et al. (2016) crucially relies on the monotonicity condition $S^{1}\geq S^{0}$ for the identification of PSACEs. For brevity, we do not restate their theorem and detailed assumptions here. Instead, Supplementary Material S7 provides a formal restatement of the monotonicity-based route of Jiang et al. (2016). There we also provide an alternative proof for their identification theorem, which motivates a least-squares estimator. These complement the identification proof and Bayesian estimator in Jiang et al. (2016). We remark that this monotonicity-based route is a benchmark for comparison, whereas Corollary 1 and Theorem 4 are our new identification results.
Without the monotonicity condition $S^{1}\geq S^{0}$, Jiang et al. (2016) also shows that a necessary condition for local identifiability of the principal effects is $m\geq 3$. In contrast, our Corollary 1, which does not impose monotonicity, establishes sufficient conditions for global identifiability when $m\geq 4$. In this sense, our new identification results provide alternative conditions leading to stronger global identification, which complement the findings of Jiang et al. (2016).
In summary, our paper complements Jiang et al. (2016) by developing new transportability-based identification routes from the broader perspective of joint potential-outcome distributions, while their monotonicity-based route serves as a useful benchmark. In Sections 6 and 7, we compare these two routes in both simulation and real-data analyses.
Simulation
We perform simulation studies to evaluate the finite-sample performance of the proposed method. We consider both scenarios, with and without the post-treatment surrogate $S$. The replication code for both simulation and application is available at https://github.com/pengwu1224/The-Promises-of-Multiple-Experiments. Throughout this simulation, the number of trials is set to 10, and for each trial $g$, the binary treatment $A$ is randomly assigned with probability $\mathbb{P}(A=1\mid G=g)=0.5$, the sample size is set to 100, 200, and 500, respectively.
Study I (without the post-treatment variable $S$). We first examine the methods developed in Sections 34 by considering two data-generating processes for $(Y^{1},Y^{0})$.
  • (C1)
    $\mathbb{P}(Y^{1}=1\mid Y^{0},G=g)=\text{expit}(Y^{0}-0.5)$ for g=1,,10g=1,...,10, where $\text{expit}(x)=\exp(x)/\{1+\exp(x)\}$ is the standard logistic function. For each trial g=1,2,,10g=1,2,...,10, the potential outcome $Y^{0}$ follows from a Bernoulli distribution with $\mathbb{P}(Y^{0}=1\mid G=g)=0.5+(g-1)/30$, i.e., taking evenly spaced values at equal intervals from 0.5 to 0.8.
  • (C2)
    $\mathbb{P}(Y^{1}=1\mid Y^{0},G=g)=\text{expit}(Y^{0}+0.5)$ for g=1,,10g=1,...,10 and $Y^{0}$ is generated according to the process described in (C1).
Assumptions 1-2 and Condition 1 hold for both cases (C1) and (C2). The true values of $\theta=(\pi_{1|0},\pi_{1|1})=(\mathbb{P}(Y^{1}=1\mid Y^{0}=0),\mathbb{P}(Y^{1}=1\mid Y^{0}=1))$ are $(0.378,0.622)$ and $(0.622,0.818)$ for cases (C1) and (C2), respectively. We replicate each simulation case 1,000 times and calculate the Bias, SD, ESE, and CP95 as evaluation metrics, where Bias and SD represent the Monte Carlo bias and standard deviation of the point estimates over the 1,000 replicates, ESE denotes the square root of the average of the estimated asymptotic variances, and CP95 refers to the coverage proportion of the 95% confidence intervals. Both ESE and CP95 are calculated using the estimated asymptotic variance based on 100 bootstraps.
Table 1 summarizes the numerical results for the proposed estimator of $\theta$ for cases (C1)–(C2). From the table, the Bias is small and decreases as the sample size increases, demonstrating the consistency of the proposed estimator. As the sample size grows, ESE approaches SD, and CP95 converges to its nominal value of 0.95, indicating the asymptotic normality of the proposed estimator and validating the bootstrap method for estimating asymptotic variance.
Table 1: Simulation results for cases (C1)–(C2).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C1) $\pi_{1|0}$ 0.041 0.136 0.141 0.938 0.017 0.104 0.107 0.946 0.009 0.068 0.070 0.952
$\pi_{1|1}$ -0.024 0.074 0.077 0.943 -0.009 0.057 0.058 0.945 -0.006 0.037 0.038 0.956
(C2) $\pi_{1|0}$ 0.037 0.123 0.122 0.937 0.024 0.092 0.092 0.944 0.007 0.060 0.060 0.951
$\pi_{1|1}$ -0.021 0.067 0.067 0.935 -0.013 0.049 0.050 0.939 -0.003 0.032 0.033 0.947
Note: Bias and SD are the Monte Carlo bias and standard deviation over the 1000 simulations of the point estimates, ESE and CP95 are the estimated asymptotic variances and coverage proportions of the 95% confidence intervals based on 100 bootstraps, respectively.
Study II (with the post-treatment variable $S$). We then explore the proposed method in the presence of a post-treatment variable $S$. Two additional data-generating processes for the potential outcomes $(S^{0},S^{1},Y^{0},Y^{1})$ are considered.
  • (C3)
    $\mathbb{P}(Y^{1}=1\mid Y^{0},S^{0},G=g)=\text{expit}\{(S^{0}+Y^{0}+1)/2\}$ and $\mathbb{P}(S^{1}=1\mid Y^{0},S^{0},G=g)=\text{expit}\{(S^{0}+Y^{0}-1)/2\}$, and $S^{0}$ and $Y^{0}$ are independent, both drawn from the Bernoulli distribution with success probability 0.5+(g1)/300.5+(g-1)/30.
  • (C4)
    $\mathbb{P}(Y^{1}=1\mid S^{0},S^{1},G=g)=\text{expit}\{(S^{0}+S^{1}+1)/2\}$ and $\mathbb{P}(Y^{0}=1\mid S^{0},S^{1},G=g)=\text{expit}\{(S^{0}+S^{1}-1)/2\}$, and $S^{0}$ and $S^{1}$ are independent, both drawn from the Bernoulli distribution, with the success probabilities 0.3+(g1)/300.3+(g-1)/30 and 0.5+(g1)/300.5+(g-1)/30, respectively. After generating $(S^{0},S^{1})$, we further adjust the value of $S^{0}$, setting $S^{0}$ to 0 if $S^{1}=0$ to ensure monotonicity.
Assumptions 34 and Condition 3 hold for case (C3), while Assumptions 3, S4S5 and Condition S3 in Supplementary Material S7 hold for case (C4). In case (C3), we denote $\pi_{s|ab}=\mathbb{P}(S^{1}=1\mid S^{0}=a,Y^{0}=b)$ and $\pi_{y|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,Y^{0}=b)$ for $a,b=0,1$. The parameters of interest are $\theta=(\pi_{s|00},\pi_{s|01},\pi_{s|10},\pi_{s|11},\pi_{y|00},\pi_{y|01},\pi_{y|10},\pi_{y|11}),$ which are the key invariant parameters for estimating the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},S^{1},Y^{1}\mid G=g)$ for g=1,,10g=1,...,10. In case (C4), we define $\pi_{1|ab}=\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b)$ and $\tilde{\pi}_{1|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b)$ for $a,b=0,1$. By the monotonicity condition $S^{1}\geq S^{0}$, $\pi_{1|10}$ and $\tilde{\pi}_{1|10}$ are undefined. Thus, the key invariance parameters are $\theta=(\pi_{1|00},\pi_{1|01},\pi_{1|11},\tilde{\pi}_{1|00},\tilde{\pi}_{1|01},\tilde{\pi}_{1|11}).$ It is noteworthy that cases (C3) and (C4) correspond to two different identification routes in Section 5. Case (C3) illustrates our new transportability-based identification route in Corollary 1. By contrast, case (C4) corresponds to the monotonicity-based route of Jiang et al. (2016), which is restated in Supplementary Material S7.
Table 2: Simulation results for case (C3).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C3) $\pi_{s|00}$ 0.032 0.431 0.420 0.951 0.012 0.382 0.388 0.964 0.017 0.357 0.354 0.952
$\pi_{s|01}$ 0.005 0.427 0.375 0.933 -0.008 0.393 0.378 0.942 -0.010 0.381 0.365 0.945
$\pi_{s|10}$ -0.005 0.396 0.379 0.944 0.020 0.416 0.374 0.939 -0.003 0.375 0.362 0.948
$\pi_{s|11}$ -0.011 0.140 0.141 0.955 -0.011 0.118 0.121 0.952 0.000 0.101 0.102 0.963
$\pi_{y|00}$ 0.026 0.396 0.363 0.940 0.013 0.337 0.334 0.956 0.014 0.310 0.308 0.961
$\pi_{y|01}$ 0.007 0.363 0.324 0.937 -0.013 0.346 0.323 0.945 -0.021 0.349 0.320 0.943
$\pi_{y|10}$ -0.002 0.342 0.327 0.944 0.006 0.340 0.322 0.948 0.007 0.339 0.314 0.940
$\pi_{y|11}$ -0.009 0.127 0.120 0.937 -0.001 0.102 0.104 0.962 0.003 0.089 0.088 0.957
Table 3: Simulation results for case (C4).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C4) $\pi_{1|00}$ 0.040 0.155 0.142 0.931 0.033 0.130 0.128 0.938 0.020 0.094 0.101 0.954
$\pi_{1|01}$ -0.043 0.160 0.144 0.923 -0.037 0.131 0.131 0.943 -0.021 0.097 0.103 0.946
$\pi_{1|11}$ -0.000 0.039 0.039 0.943 0.001 0.028 0.028 0.949 -0.001 0.018 0.018 0.940
$\tilde{\pi}_{1|00}$ 0.001 0.037 0.037 0.942 0.000 0.026 0.026 0.952 -0.000 0.017 0.016 0.948
$\tilde{\pi}_{1|01}$ 0.019 0.109 0.113 0.953 0.012 0.092 0.094 0.962 0.004 0.064 0.065 0.947
$\tilde{\pi}_{1|11}$ -0.022 0.125 0.127 0.954 -0.013 0.106 0.106 0.956 -0.004 0.073 0.074 0.940
Tables 2 and 3 present the numerical results for cases (C3) and (C4), respectively. The simulation results are similar to those in cases (C1)–(C2), indicating that the extension of the proposed method presented in Section 5 performs well. Taken together, these two cases show that the proposed least-squares framework performs well both under our new transportability-based route and under the monotonicity-based route corresponding to Jiang et al. (2016). To further evaluate the robustness of the proposed method, we present additional simulations for the just-identified case and for the violation of Assumption 2 in Supplementary Material S9.
Application to the Adjuvant Colon Clinical Trials
We demonstrate the proposed methodology using the data from phase III adjuvant colon clinical trials (ACCTs). The initial ACCTs data consist of 20,898 patients from 18 randomized phase III clinical trials, with the enrollment period spanning from 1977 to 1999 (Sargent et al. 2005). Among these 18 trials, the data from 10 trials are publicly available in Baker et al. (2012), including a total of 9,102 patients. We base our analysis on these 10 trials. In each trial, we have a contingency table of observed frequencies for three binary variables: treatment $A$, surrogate $S$, and outcome $Y$. Here $S=0$ indicates cancer recurrence within 3 years and $S=1$ otherwise; $Y=0$ denotes mortality within 5 years and $Y=1$ indicates survival beyond 5 years; $A=1$ and $A=0$ denote receiving treatment ( fluorouracil-based chemotherapy) and not, respectively. The public data do not contain individual-level baseline covariates. Accordingly, as in Baker et al. (2012) and Jiang et al. (2016), our empirical analysis here does not adjust for covariates.
The goal of the ACCTs is to determine whether cancer recurrence within 3 years ($S$) can serve as an appropriate surrogate for overall survival with a 5-year follow-up ($Y$). Using the ACCTs data, Sargent et al. (2005) identified a strong correlation between $S$ and $Y$. From the perspective of principal stratification in causal inference, Jiang et al. (2016) investigated this question by first estimating PSACEs and then evaluating the surrogate using the causal necessity and causal sufficiency criteria (Gilbert and Hudgens 2008): for all $g\in{\mathcal{G}}$, causal necessity requires $\text{PSACE}_{11|g}=0$ and $\text{PSACE}_{00|g}=0$, and causal sufficiency requires $\text{PSACE}_{10|g}\neq 0$ and $\text{PSACE}_{01|g}\neq 0$. Causal necessity means that if $A$ has no effect on $S$ (i.e., $S^{1}=S^{0}$), then it should also have no effect on $Y$, whereas causal sufficiency means that if $A$ affects $S$ (i.e., $S^{1}\neq S^{0}$), then it should also affect $Y$.
In this article, we further explore the problem by estimating both the joint distributions of $\mathbb{P}(S^{0},S^{1}\mid G=g)$ and $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$, as well as the PSACEs under different sets of identification assumptions. Notably, for the ACCTs data, Assumptions 1 and 3 naturally hold due to randomization.
7.1 State Transition Probabilities and Joint Distribution of Potential Surrogates and Potential Outcomes
To get an intuitive understanding of the heterogeneity across different trials, we estimate the average treatment effects (ATEs) of $A$ on $S$ (or $Y$) for each trial $g\in{\mathcal{G}}$ by computing the average contrasts of $S$ (or $Y$) between the groups $(A=1,G=g)$ and $(A=0,G=g)$. As illustrated in the barplots of Figure 1, the ATEs of $A$ on $S$ and $Y$ vary substantially across trials, indicating strong cross-trial heterogeneity and ruling out stronger homogeneity assumptions that require common ATEs across trials. In addition, the ATEs of $A$ on $S$ and $Y$ show a similar pattern, indicating that $S$ may have the potential to serve as an appropriate surrogate for $Y$.
Figure 1: (a) Estimated ATEs of $A$ on $S$, and scatter plots of $\hat{\mathbb{P}}(S^{1}=1\mid G=g)$ against $\hat{\mathbb{P}}(S^{0}=0\mid G=g)$ and $\hat{\mathbb{P}}(S^{0}=1\mid G=g)$, respectively. The scatter plots show strong linear correlations, providing informal visual evidence for the linear restrictions implied by $S^{1}\perp\!\!\!\perp G\mid S^{0}$. A linear regression yields an $R^{2}$ of 0.9996; the formal over-identification test assessment is reported in Table 4. (b) Estimated ATEs of $A$ on $Y$, and scatter plots of $\hat{\mathbb{P}}(Y^{1}=1\mid G=g)$ against $\hat{\mathbb{P}}(Y^{0}=0\mid G=g)$ and $\hat{\mathbb{P}}(Y^{0}=1\mid G=g)$, respectively. The scatter plots likewise provide informal visual evidence for the linear restrictions implied by $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$. A linear regression yields an $R^{2}$ of $0.9991$; the formal assessment is again reported in Table 4.
By employing the proposed method outlined in Sections 34, and under Assumption 2 (i.e., $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$), we can estimate the state transition probabilities from $Y^{0}$ to $Y^{1}$, i.e., $\mathbb{P}(Y^{1}\mid Y^{0})$. This allows us to subsequently estimate the joint distributions $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for $g\in{\mathcal{G}}$. Similarly, if we assume $S^{1}\perp\!\!\!\perp G\mid S^{0}$, then we can apply the same methods to estimate the transition probabilities $\mathbb{P}(S^{1}\mid S^{0})$ and the joint distributions $\mathbb{P}(S^{0},S^{1}\mid G=g)$.
The key assumptions $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$ and $S^{1}\perp\!\!\!\perp G\mid S^{0}$ imply a linear relationship between $\mathbb{P}(Y^{1}=1\mid G=g)$ on $\mathbb{P}(Y^{0}=1\mid G=g)$ and $\mathbb{P}(Y^{0}=0\mid G=g))$, as well as between $\mathbb{P}(S^{1}=1\mid G=g)$ on $\mathbb{P}(S^{0}=1\mid G=g)$ and $\mathbb{P}(S^{0}=0\mid G=g))$. We illustrate these possible linear relationships through the scatter plots in Figure 1, where probabilities are replaced by observed frequencies. These plots provide informal visual evidence for the implied restrictions. Moreover, a linear regression of $\hat{\mathbb{P}}(S^{1}=1\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0\mid G=g),\hat{\mathbb{P}}(S^{0}=1\mid G=g))$ yields an $R^{2}$ of 0.9996. Similarly, A linear regression of $\hat{\mathbb{P}}(Y^{1}=1\mid G=g)$ on $(\hat{\mathbb{P}}(Y^{0}=0\mid G=g),\hat{\mathbb{P}}(Y^{0}=1\mid G=g))$ yields an $R^{2}$ of $0.9991$. These results indicate a strong linear relationship. We then estimate the invariant parameters $(\mathbb{P}(S^{1}=1\mid S^{0}=0),\mathbb{P}(S^{1}=1\mid S^{0}=1))$ and $(\mathbb{P}(Y^{1}=1\mid Y^{0}=0),\mathbb{P}(Y^{1}=1\mid Y^{0}=1))$, with the results reported in Table 4. Table 4 also reports the overidentification tests for $S^{1}\perp\!\!\!\perp G\mid S^{0}$ and $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$. The corresponding $p$-values are 0.846 and 0.325, respectively, both of which are very large. So we do not reject the observable restrictions implied by these two assumptions.
Table 4: State transition probabilities and tests for $S^{1}\perp\!\!\!\perp G\mid S^{0}$ and $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$.
Parameters Estimate (ESE) 95% CI $J$-Statistic $p$-value null hypothesis $H_{0}$
$\mathbb{P}(S^{1}=1\mid S^{0}=0)$ 0.379 (0.107) (0.170, 0.590) 4.119 0.846 $S^{1}\perp\!\!\!\perp G\mid S^{0}$
$\mathbb{P}(S^{1}=1\mid S^{0}=1)$ 0.896 (0.049) (0.800, 0.992)
$\mathbb{P}(S^{1}=0\mid S^{0}=0)$ 0.621 (0.107) (0.411, 0.830)
$\mathbb{P}(S^{1}=0\mid S^{0}=1)$ 0.104 (0.488) (0.008, 0.200)
$\mathbb{P}(Y^{1}=1\mid Y^{0}=0)$ 0.275 (0.101) (0.062, 0.488) 9.208 0.325 $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$
$\mathbb{P}(Y^{1}=1\mid Y^{0}=1)$ 0.946 (0.051) (0.846, 1.044)
$\mathbb{P}(Y^{1}=0\mid Y^{0}=0)$ 0.725 (0.101) (0.512, 0.938)
$\mathbb{P}(Y^{1}=0\mid Y^{0}=1)$ 0.054 (0.051) (-0.045, 0.154)
Note: ESE is the estimated asymptotic standard error based on 500 bootstraps, 95% CI represents 95% confidence interval, and the $J$-Statistic and $p$-value correspond to the null hypothesis.
Figure 2: Estimated joint probabilities $\hat{\mathbb{P}}(S^{0}=1,S^{1}=0\mid G=g)$ and $\hat{\mathbb{P}}(Y^{0}=1,Y^{1}=0\mid G=g)$ for all $g\in\mathcal{G}$. Since these quantities would be zero under the monotonicity conditions $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$, respectively, the figure provides direct evidence for assessing the plausibility of the two monotonicity assumptions.
Another useful by-product of our framework is that it allows us to assess whether the monotonicity assumptions are plausible in the ACCT data. Jiang et al. (2016) also examined monotonicity by comparing models with and without monotonicity. By contrast, once our framework identifies the joint distributions $\mathbb{P}(S_{0},S_{1}\mid G=g)$ and $\mathbb{P}(Y_{0},Y_{1}\mid G=g)$, we can directly examine quantities such as $\mathbb{P}(S_{0}=1,S_{1}=0\mid G=g)$ and $\mathbb{P}(Y_{0}=1,Y_{1}=0\mid G=g)$, which would be zero under the corresponding monotonicity assumptions. Figure 2 displays the point estimates along with the corresponding 95% confidence intervals (CIs). The lower bounds of the 95% CIs for $\hat{\mathbb{P}}(S^{0}=1,S^{1}=0\mid G=g)$ for all trials are strictly positive, suggesting that the monotonicity condition (Assumption S5, $S^{1}\geq S^{0}$) may not hold. In contrast, the 95% CIs for $\hat{\mathbb{P}}(Y^{0}=1,Y^{1}=0\mid G=g)$ cover 0 in all trials. The other estimated values of $\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$ and $\mathbb{P}(Y^{0}=a,Y^{1}=b\mid G=g)$ for $a,b\in\{0,1\}$ are provided in Supplementary Material S9.
7.2 Evaluation of Principal Surrogate
To evaluate the surrogate, we estimate the principal stratification average causal effects $\text{PSACE}_{ab|g}$ for $g\in{\mathcal{G}}$. Based on the methods proposed in Section 5, we consider the following four approaches under different sets of assumptions. The four methods serve both as a sensitivity analysis and as a way to distinguish the monotonicity-based route most closely associated with Jiang et al. (2016) from our transportability-based routes.
  • Method 1: Based on the monotonicity-based identification route of Jiang et al. (2016), relying on Assumptions 3, S4, and Condition S3, with monotonicity $S^{1}\geq S^{0}$. See Supplementary Material S7 for details on identification and estimation.
  • Method 2: Based on Theorem 4, relying on Assumptions 3, 4, and Condition 4, with monotonicity $Y^{1}\geq Y^{0}$ and $S^{1}\geq S^{0}$. See Supplementary Material S4 for estimation details.
  • Method 3: Based on Corollary 1, relying on Assumptions 3, 4, and Condition 3. We additionally impose the partial monotonicity $Y^{1}\geq Y^{0}$ to simplify the identification and estimation of a subset of parameters. See Supplementary Material S5 for estimation details. We include this method for comparison primarily because Figure 2 suggests that $S^{1}\geq S^{0}$ may fail.
  • Method 4: Based on Corollary 1, relying on Assumptions 3, 4, and Condition 3, without any monotonicity. See Supplementary Material S4 for estimation details.
Figure 3 presents the point estimates and the associated 95% confidence intervals of $\text{PSACE}_{ab\mid g}$ across trials, with the four columns corresponding to Methods 1–4, respectively. When $S^{1}\geq S^{0}$ is assumed (Methods 1 and 2), $\text{PSACE}_{10\mid g}$ is undefined and is therefore left blank. In addition, under Assumption S4 ($G\perp\!\!\!\perp Y^{a}\mid(S^{0},S^{1})$, Method 1), $\text{PSACE}_{ab\mid g}$ is invariant across trials for fixed $a,b\in\{0,1\}$.
Method 1
Method 2
Method 3
Method 4
Figure 3: Point estimates and 95% confidence intervals of $\textup{PSACE}_{ab\mid g}$ across trials under Methods 1–4. From top to bottom, the four rows correspond to $(a,b)=(0,0),(0,1),(1,0),(1,1)$. Methods 1 and 2 assume monotonicity $S^{1}\geq S^{0}$, so $\textup{PSACE}_{10\mid g}$ is undefined. The vertical scale in Method 4 differs from that in the other columns because its confidence intervals are substantially wider.
Figure 3 suggests that causal necessity may hold, as the 95% confidence intervals of $\text{PSACE}_{00\mid g}$ and $\text{PSACE}_{11\mid g}$ all include zero, indicating no clear evidence that treatment affects survival when it does not materially affect cancer recurrence. By contrast, the evidence for causal sufficiency is mixed. The point estimates of $\text{PSACE}_{01\mid g}$ are uniformly positive, suggesting that patients whose cancer recurrence would be prevented by treatment are also likely to benefit in survival. However, the 95% confidence intervals of $\text{PSACE}_{10\mid g}$ generally include zero, indicating ambiguous survival responses among patients whose cancer recurrence would be worsened by treatment. So these results provide only partial support for causal sufficiency.
Method 1 corresponds most closely to the monotonicity-based identification route in Jiang et al. (2016), although our estimation procedure here is least-squares-based rather than Bayesian. In their preferred empirical analysis without monotonicity, Jiang et al. (2016) conclude that both causal necessity and causal sufficiency are supported. Our Methods 3 and 4, which do not rely on the monotonicity condition $S^{1}\geq S^{0}$ and are therefore more plausible in light of Figure 2, likewise support causal necessity and also show positive evidence for the causal sufficiency within the latent stratum with $S^{0}=0$ and $S^{1}=1$. However, because the confidence intervals for $\text{PSACE}_{10\mid g}$ generally cover zero, our analysis yields a more cautious conclusion about full causal sufficiency. Thus, the two analyses agree that monotonicity is questionable for the ACCT data and that causal necessity is supported, but our approach yields a more cautious conclusion about full causal sufficiency.
Conclusion
In this work, we propose a novel framework that leverages multiple experimental datasets to identify and estimate the joint distribution of potential outcomes. First, we established the identifiability of joint distributions for binary and categorical outcomes under the assumption of transportability of state transition probabilities (Assumption 2) and a rank condition (Conditions 1 and 2). Second, we introduce a simple least-squares-based method for estimating the joint distribution of potential outcomes. Finally, we extended our framework to identify and estimate principal stratification causal effects. These extend and complement existing methods in causal inference that integrate information from multiple datasets.
Supplementary Material
Technical Proofs
Throughout the proofs, for events $B_{1}$ and $B_{2}$, we set $\mathbb{P}(B_{1}\mid B_{2})=0$ whenever $\mathbb{P}(B_{2})=0$ to ensure well-definedness.
S1.1 Proof of Theorem 1
Proof of Theorem 1.
Recall that $\pi_{b|a}=\mathbb{P}(Y^{1}=b\mid Y^{0}=a)$ for $a,b=0,1$ are invariant across trials under Assumption 2. Note that
$\begin{cases}\mathbb{P}(Y^{1}=1\mid G=1)=\pi_{1|0}\cdot\mathbb{P}(Y^{0}=0\mid G=1)+\pi_{1|1}\cdot\mathbb{P}(Y^{0}=1\mid G=1)\\ \mathbb{P}(Y^{1}=1\mid G=2)=\pi_{1|0}\cdot\mathbb{P}(Y^{0}=0\mid G=2)+\pi_{1|1}\cdot\mathbb{P}(Y^{0}=1\mid G=2)\\ \vdots\\ \mathbb{P}(Y^{1}=1\mid G=m)=\pi_{1|0}\cdot\mathbb{P}(Y^{0}=0\mid{\color[rgb]{0,0,0}G=m})+\pi_{1|1}\cdot\mathbb{P}(Y^{0}=1\mid G=m)\end{cases}$
and $\mathbb{P}(Y^{1}=b\mid G=g)$ and $\mathbb{P}(Y^{0}=a\mid G=g)$ are identifiable under Assumption 1. Then, the above system of equations contains a total of $m$ equations and $2$ unknown free parameters ($\pi_{1|0}$ and $\pi_{1|1}$). Thus, under Condition 1 $\pi_{1|0}$ and $\pi_{1|1}$ are identifiable by solving the system of equations. In addition, $\pi_{0|0}=1-\pi_{1|0}$ and $\pi_{0|1}=1-\pi_{1|1}$ are identifiable.
The identifiability of $\mathbb{P}(Y^{1}=b,Y^{0}=a\mid G=g)$ follows immediately from the fact that
$\mathbb{P}(Y^{1}=b,Y^{0}=a\mid G=g)={}$ $\mathbb{P}(Y^{1}=b\mid Y^{0}=a)\cdot\mathbb{P}(Y^{0}=a\mid G=g)$
$={}$ $\pi_{b|a}\cdot\mathbb{P}(Y=a\mid A=0,G=g)~\text{for }g={\color[rgb]{0,0,0}1,2,...,m};a,b=0,1.$
S1.2 Proof of Theorem 2
Proof of Theorem 2.
Under Assumption 2, $\pi_{i|j}:=\mathbb{P}(Y^{1}=i\mid Y^{0}=j)$ for i,j=1,2,,ki,j=1,2,...,k are invariant across trials. Thus, we have that
$\mathbb{P}(Y^{1}=i\mid G=g)=\sum_{j=1}^{k}\pi_{i|j}\mathbb{P}(Y^{0}=j\mid G=g),\quad i=1,...,k-1;g=1,...,m,$
where $\mathbb{P}(Y^{1}=i\mid G=g)$ and $\mathbb{P}(Y^{0}=j\mid G=g)$ are identifiable under Assumption 1. Under Condition 2, the above system of equations include a total of $m\times(k-1)$ equations, however, the quantities $\{\pi_{i|j},i=1,...,k;j=1,...,k\}$ contains $k\times(k-1)$ free parameters due to the truth that $\sum_{i=1}^{k}\pi_{i|j}=1$ for each $j$. Thus, when $m\geq k$, the quantities $\{\pi_{i|j},i=1,...,k;j=1,...,k\}$ are identifiable, and the conclusion of Theorem 2 holds immediately from the fact that
$\mathbb{P}(Y^{1}=i,Y^{0}=j\mid G=g)=\pi_{i|j}\cdot\mathbb{P}(Y^{0}=j\mid G=g)~\text{for }i,j=1,...,k.$
This finishes the proof.
S1.3 Proof of Theorem 3
Proof of Theorem 3.
Recall that $m$ is the number of trials, which is a fixed number; and $n_{g}$ is the sample size of observed data in the $g$-th trial, which converges to infinity when we discussing the large sample properties, and $n=\sum_{g=1}^{m}n_{g}$ is the total sample size of the observed data.
For ease of presentation, we let $\tilde{Y}_{g}=\mathbb{P}(Y=1\mid G=g,A=1),$ $\tilde{X}_{1g}=\mathbb{P}(Y=0\mid G=g,A=0),$ and $\tilde{X}_{2g}=\mathbb{P}(Y=1\mid G=g,A=0)$. Denote $\tilde{X}_{g}=(\tilde{X}_{1g},\tilde{X}_{2g})^{\intercal}$. Then by equation (2), the true value of $\theta=(\pi_{1|0},\pi_{1|1})$ can be written as
$\theta=\left(\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{Y}_{g}.$
However, if we use linear regression for estimating $\theta$, its estimator is
$\hat{\theta}=\left(\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g},$
where $\hat{X}_{g}$ and $\hat{Y}_{g}$ are the estimators of $\tilde{X}_{g}$ and $\tilde{Y}_{g}$, respectively, by replacing probabilities with frequencies. For example,
$\hat{Y}_{g}=\frac{n^{-1}\sum_{i=1}^{n}\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)}{n^{-1}\sum_{i=1}^{n}\mathbb{I}(G_{i}=g,A_{i}=1)}.$
For $\hat{Y}_{g}$, we have that
$\sqrt{n}(\hat{Y}_{g}-\tilde{Y}_{g})={}$ $\frac{n^{-1/2}\sum_{i=1}^{n}\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)}{\mathbb{P}(G=g,A=1)}-\frac{n^{-1/2}\sum_{i=1}^{n}\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(G=g,A=1)}+o_{\mathbb{P}}(1)$
$={}$ $\frac{1}{\mathbb{P}(G=g,A=1)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)-\mathbb{P}(Y=1,G=g,A=1)]+o_{\mathbb{P}}(1).$
Similarly,
$\sqrt{n}(\hat{X}_{g}-\tilde{X}_{g})={}$ $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{pmatrix}\dfrac{\mathbb{I}(Y_{i}=0,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=0,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\\ \dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=1,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\end{pmatrix}+o_{\mathbb{P}}(1).$
By the strong law of large numbers, $\hat{X}_{g}$ converges to $\tilde{X}_{g}$ almost surely, which implies that $m^{-1}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}$ converges to $m^{-1}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}$ almost surely. Thus, $\hat{\theta}$ has the same asymptotical distribution as
$\bar{\theta}=\left(\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g}.$
Next, we focus on analyzing $\tilde{\theta}$. Let $C:=m^{-1}\sum_{g=1}^{m}\tilde{X}_{g}\tilde{X}_{g}^{\intercal}$, then
$\sqrt{n}(\bar{\theta}-\theta)={}$ $C^{-1}\cdot\frac{\sqrt{n}}{m}\sum_{g=1}^{m}(\hat{X}_{g}\hat{Y}_{g}-\tilde{X}_{g}\tilde{Y}_{g}).$
$={}$ $C^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\{\sqrt{n}(\hat{X}_{g}-\tilde{X}_{g})\hat{Y}_{g}+\sqrt{n}(\hat{Y}_{g}-\tilde{Y}_{g})\tilde{X}_{g}\}$
$={}$ $C^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\{\sqrt{n}(\hat{X}_{g}-\tilde{X}_{g})\tilde{Y}_{g}+\sqrt{n}(\hat{Y}_{g}-\tilde{Y}_{g})\tilde{X}_{g}\}+o_{\mathbb{P}}(1)$
$={}$ $C^{-1}\cdot\left\{\frac{1}{m}\sum_{g=1}^{m}\tilde{Y}_{g}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{pmatrix}\dfrac{\mathbb{I}(Y_{i}=0,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=0,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\\ \dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=1,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\end{pmatrix}\right\}$
$+{}$ $~C^{-1}\cdot\left\{\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)-\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(G=g,A=1)}\right\}+o_{\mathbb{P}}(1)$
$={}$ $C^{-1}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{\frac{1}{m}\sum_{g=1}^{m}\tilde{Y}_{g}\begin{pmatrix}\dfrac{\mathbb{I}(Y_{i}=0,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=0,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\\ \dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=1,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\end{pmatrix}\right\}$
$+{}$ $~C^{-1}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{1}{m}\sum_{g=1}^{m}\tilde{X}_{g}\cdot\dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)-\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(G=g,A=1)}\right)+o_{\mathbb{P}}(1).$
Therefore, we have that
$\sqrt{n}(\bar{\theta}-\theta)\xrightarrow{d}N(0,\sigma^{2}),$
where $\sigma^{2}=C^{-1}VC^{-1}$, $V$ equals to the variance of
$\frac{1}{m}\sum_{g=1}^{m}\Biggl\{\tilde{Y}_{g}\begin{pmatrix}\dfrac{\mathbb{I}(Y_{i}=0,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=0,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\\ \dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=0)-\mathbb{P}(Y=1,G=g,A=0)}{\mathbb{P}(G=g,A=0)}\end{pmatrix}$
$+\tilde{X}_{g}\cdot\dfrac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)-\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(G=g,A=1)}\Biggr\}.$
This finishes the proof.
S1.4 Proof of Corollary 1
Proof of Corollary 1.
Define $\pi_{cd|ab}=\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b)$ for $a,b,c,d=0,1$, which are invariant parameters across trials by Assumption 4. Taking $(S^{1},Y^{1})$ and $(Y^{0},S^{0})$ as two category variables taking four different values, then following the proof of Theorem 2, under Assumption 4, we have the following system of equations
$\begin{cases}\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)=\sum_{a,b}\pi_{cd|ab}\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g),~\text{for }c,d=0,1;g=1,...,m\\ \sum_{c,d}\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)=1,\qquad g=1,...,m\\ \sum_{a,b}\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)=1,\qquad g=1,...,m\end{cases}$
where $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ are identifiable quantities under Assumption 3.
The above system of equations contains a total of $m\times 3$ linearly independent equations, and $3\times 4=12$ unknown free parameters (i.e., $\{(\pi_{11|ab},\pi_{10|ab},\pi_{01|ab}):a,b=0,1\}$) due to $\sum_{c,d}\pi_{cd|ab}=1$ for $c=0,1;d=0,1$. Then under Condition 3, $\pi_{cd|ab}$ for $c=0,1;d=0,1$ are identifiable. This also implies the identifiability of $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ by noting that
$\mathbb{P}(S^{0}=a,S^{1}=c,Y^{0}=b,Y^{1}=d\mid G=g)=\pi_{cd|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g).$
Therefore, $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ and $g\in{\mathcal{G}}$ are also identifiable.
S1.5 Proof of Theorem 4
Proof of Theorem 4.
Under Assumption 4,
$\pi_{cd|ab}=\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b),\quad a,b,c,d=0,1,$
are invariant parameters across trials. When the monotonicity condition $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$ holds, it has 9 possible values of $(a,b,c,d)$ with 5 free parameters in $\{\pi_{cd|ab},a,b,c,d=0,1\}$, as shown in Table S1.
Table S1: Parameters under Assumption 4 and the monotonicity condition $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$
$S^{0}=a$ $Y^{0}=b$ $S^{1}=c$ $Y^{1}=d$ Restriction Number of Free Parameters
0 0 0 0 $\sum_{c=0}^{1}\sum_{d=0}^{1}\pi_{cd|00}=1$ 3
0 0 0 1
0 0 1 0
0 0 1 1
0 1 0 1 $\sum_{c=0}^{1}\pi_{c1|01}=1$ 1
0 1 1 1
1 0 1 0 $\sum_{d=0}^{1}\pi_{1d|10}=1$ 1
1 0 1 1
1 1 1 1 $\pi_{11|11}=1$ 0
On the other hand, under Assumption 4, we have the following system of equations
$\begin{cases}\mathbb{P}(S^{1}=0,Y^{1}=0|G=g)={}&\pi_{00|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g),\\ \mathbb{P}(S^{1}=0,Y^{1}=1|G=g)={}&\pi_{01|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{01|01}\cdot\mathbb{P}(S^{0}=0,Y^{0}=1|G=g),\\ \mathbb{P}(S^{1}=1,Y^{1}=0|G=g)={}&\pi_{10|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{10|10}\cdot\mathbb{P}(S^{0}=1,Y^{0}=0|G=g),\\ \mathbb{P}(S^{1}=1,Y^{1}=1|G=g)={}&\pi_{11|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{11|01}\cdot\mathbb{P}(S^{0}=0,Y^{0}=1|G=g)\\ \hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad+{}&\pi_{11|10}\cdot\mathbb{P}(S^{0}=1,Y^{0}=0|G=g)+\pi_{11|11}\cdot\mathbb{P}(S^{0}=1,Y^{0}=1|G=g),\\ \qquad\text{for }g=1,...,m.\end{cases}$
where $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ are identifiable quantities under Assumption 3. Clearly, $\pi_{00|00}$ is identifiable from the first row of the above system of equations. When Condition 4 holds, $\{\pi_{01|00},\pi_{01|01},\pi_{10|00},\pi_{10|10}\}$ are identifiable from the second and third rows of the above system of equations. Then by the restriction (the second column) in Table S1, all the parameters in $\{\pi_{cd|ab},a,b,c,d=0,1\}$ are identifiable. Then, $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ is identifiable due to
$\mathbb{P}(S^{0}=a,S^{1}=c,Y^{0}=b,Y^{1}=d\mid G=g)=\pi_{cd|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g).$
Therefore, $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ and $g\in{\mathcal{G}}$ are also identifiable.
S1.6 Proof of Theorem S2
Proof of Theorem S2.
As discussed below Assumption S5 in the main text, under Assumptions 3 and S5, the principal scores $\delta_{ab|g}=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$ for $a=0,1$ and $b=0,1$ are identifiable.
We first prove Theorem S2(a). Define
$\pi_{1|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b),\quad a=0,1;b=0,1,$
which are invariant parameters across trials by Assumption S4. Note that by Assumptions 3 and S5, $\pi_{1|10}=0$ and
$\pi_{1|00}=\mathbb{P}(Y^{1}=1\mid S^{0}=0,S^{1}=0)=\mathbb{P}(Y^{1}=1\mid S^{1}=0)=\mathbb{P}(Y=1\mid S=0,A=1)$
is identifiable. Thus, the set $\{\pi_{1|ab},a=0,1;b=0,1\}$ contains only two free parameters $\pi_{1|01}$ and $\pi_{1|11}$. On the other hand, we have the following decomposition
$\mathbb{P}(Y^{1}=1\mid G=g)={}$ $\mathbb{P}(Y^{1}=1\mid S^{0}=0,S^{1}=0,G=g)\mathbb{P}(S^{0}=0,S^{1}=0\mid G=g)$
$+\mathbb{P}(Y^{1}=1\mid S^{0}=0,S^{1}=1,G=g)\mathbb{P}(S^{0}=0,S^{1}=1\mid G=g)$
$+\mathbb{P}(Y^{1}=1\mid S^{0}=1,S^{1}=1,G=g)\mathbb{P}(S^{0}=1,S^{1}=1\mid G=g)$
$={}$ $\pi_{1|00}\delta_{00|g}+\pi_{1|01}\delta_{01|g}+\pi_{1|11}\delta_{11|g},\qquad g=1,...,m,$
which has $m$ equations. Clearly, when Condition S3(i) holds, $\pi_{1|01}$ and $\pi_{1|11}$ are identifiable by solving the above equations. Therefore, $\mathbb{P}(Y^{1}\mid S^{0},S^{1},G=g)$ is identifiable.
In addition, the joint distribution $\mathbb{P}(Y^{1},S^{0},S^{1}|G=g)$ for $g\in{\mathcal{G}}$ are also identifiable by noting that
$\mathbb{P}(Y^{1}=d,S^{0}=a,S^{1}=b\mid G=g)=\pi_{d|ab}\cdot\delta_{ab|g}.$
and both $\pi_{d|ab}$ and $\delta_{ab|g}$ are identifiable.
Next, we prove Theorem S2(b). Define
$\tilde{\pi}_{1|ab}=\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b),\quad a=0,1;b=0,1,$
which are invariant across trials by Assumption S4. By Assumptions 3 and S5, $\tilde{\pi}_{1|10}=0$ and
$\tilde{\pi}_{1|11}=\mathbb{P}(Y^{0}=1|S^{0}=1,S^{1}=1)=\mathbb{P}(Y^{0}=1|S^{0}=1)=\mathbb{P}(Y=1|S=1,A=0)$
is identifiable. Thus, the set $\{\tilde{\pi}_{1|ab},a=0,1;b=0,1\}$ contains only two free parameters $\pi_{1|00}$ and $\pi_{1|01}$. In addition, by Assumption S4, we have the following decomposition
$\mathbb{P}(Y^{0}=1|G=g)={}$ $\tilde{\pi}_{1|00}\delta_{00|g}+\tilde{\pi}_{1|01}\delta_{01|g}+\tilde{\pi}_{1|11}\delta_{11|g},\qquad g=1,...,m,$
which has $m$ equations. When Condition S3(b) holds, $\tilde{\pi}_{1|00}$ and $\tilde{\pi}_{1|01}$ are identifiable by solving the above equations. Therefore, $\mathbb{P}(Y^{0}\mid S^{0},S^{1},G=g)$ is identifiable. Also, the joint distribution of $(Y^{0},S^{0},S^{1}|G=g)$ for $gin{\mathcal{G}}$ are also identifiable by noting that
$\mathbb{P}(Y^{0}=c,S^{0}=a,S^{1}=b|G=g)=\tilde{\pi}_{c|ab}\cdot\delta_{ab|g}.$
and both $\tilde{\pi}_{c|ab}$ and $\delta_{ab|g}$ are identifiable.
Theorem S2(c) is a direct corollary of Theorems S2(a) and S2(b). This finishes the proof.
S1.7 Proof of Theorem S3
Before presenting the detailed proof of Theorem S3, we first give a general result (Lemma S1) that will be used in the proof of Theorem S3.
Lemma S1.
Consider a general estimator of $\theta\in\mathbb{R}^{p}$. Suppose that the true parameter $\theta$ and its estimator $\hat{\theta}$ take the forms
$\theta=\left(\frac{1}{m}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}X_{g}Y_{g},$
$\hat{\theta}=\left(\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g},$
where $m$ is a fixed constant, $X_{g}$ and $Y_{g}$ are constants indexed by $g$, $\hat{X}_{g}$ and $\hat{Y}_{g}$ are estimators of $X_{g}\in\mathbb{R}^{p}$ and $Y_{g}\in\mathbb{R}$, respectively. Assume that they admit asymptotic linear representations:
$\sqrt{n}(\hat{X}_{g}-X_{g})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{g}(X_{i},A_{i},S_{i},Y_{i})+o_{\mathbb{P}}(1),$
$\sqrt{n}(\hat{Y}_{g}-Y_{g})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})+o_{\mathbb{P}}(1),$
where $\phi_{g}(X_{i},A_{i},S_{i},Y_{i})\in\mathbb{R}^{p}$ and $\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})\in\mathbb{R}$ have mean zero and finite variance, $o_{\mathbb{P}}(1)$ converges to zero in probability uniformly as $n\to\infty$. Then we have $\sqrt{n}(\hat{\theta}-\theta)\xrightarrow{d}N(0,\sigma^{2}),$ where $\sigma^{2}=C^{-1}VC^{-1}$, $C=m^{-1}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}$, and
$V=$ $\mathrm{Var}\Biggl\{\frac{1}{m}\sum_{g=1}^{m}\Big(\phi_{g}(X_{i},A_{i},S_{i},Y_{i})Y_{g}+\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})X_{g}\Big)\Biggr\}.$
Proof of Lemma S1.
By the strong law of large numbers, $\hat{\theta}$ has the same asymptotical distribution as
$\bar{\theta}=\left(\frac{1}{m}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g}=C^{-1}\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g}.$
Thus, it is sufficient to analyzing $\tilde{\theta}$. Note that
$\sqrt{n}(\bar{\theta}-\theta)={}$ $C^{-1}\cdot\frac{\sqrt{n}}{m}\sum_{g=1}^{m}(\hat{X}_{g}\hat{Y}_{g}-X_{g}Y_{g}).$
$={}$ $C^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\{\sqrt{n}(\hat{X}_{g}-X_{g})\hat{Y}_{g}+\sqrt{n}(\hat{Y}_{g}-Y_{g})X_{g}\}$
$={}$ $C^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\{\sqrt{n}(\hat{X}_{g}-X_{g})Y_{g}+\sqrt{n}(\hat{Y}_{g}-Y_{g})X_{g}\}+o_{\mathbb{P}}(1)$
$={}$ $C^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\left\{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big(\phi_{g}(X_{i},A_{i},S_{i},Y_{i})Y_{g}+\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})X_{g}\Big)\right\}+o_{\mathbb{P}}(1)$
$={}$ $C^{-1}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{\frac{1}{m}\sum_{g=1}^{m}\Big(\phi_{g}(X_{i},A_{i},S_{i},Y_{i})Y_{g}+\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})X_{g}\Big)\right\}+o_{\mathbb{P}}(1).$
This finishes the proof.
Next, we prove Theorem S3.
Proof of Theorem S3.
Recall that
$\pi_{1|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b),~\tilde{\pi}_{1|ab}=\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b).$
for $a,b\in\{0,1\}$. In addition, under the monotonicity condition $S^{1}\geq S^{0}$, $\pi_{1|10}$ and $\tilde{\pi}_{1|10}$ are undefined and do not need estimation. Let $\beta:=(\pi_{1|00},\pi_{1|01},\pi_{1|11})$ and $\gamma:=(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01},\tilde{\pi}_{1|11})$. The estimation of $\beta$ and $\gamma$ relies on estimating $\delta_{ab\mid g}:=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$ for $a,b\in\{0,1\}$. We therefore first establish the asymptotic linear representation of these estimators.
First, we focus on $\delta_{ab\mid g}$ for $a,b\in\{0,1\}$. For any given $g$, $\hat{\delta}_{10|g}=0$ by Assumption S5, and the estimator of $\delta_{11|g},\delta_{01|g}$, and $\delta_{00|g}$ are given as
$\hat{\delta}_{11|g}={}$ $\frac{n^{-1}\sum_{i=1}^{n}\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)}{n^{-1}\sum_{i=1}^{n}\mathbb{I}(A_{i}=0,G_{i}=g)}$
$\hat{\delta}_{01|g}={}$ $\frac{n^{-1}\sum_{i=1}^{n}\mathbb{I}(S_{i}=1,A_{i}=1,G_{i}=g)}{n^{-1}\sum_{i=1}^{n}\mathbb{I}(A_{i}=1,G_{i}=g)}-\hat{\delta}_{11|g}$
$\hat{\delta}_{00|g}={}$ $\frac{n^{-1}\sum_{i=1}^{n}\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)}{n^{-1}\sum_{i=1}^{n}\mathbb{I}(A_{i}=1,G_{i}=g)},$
where the estimate of $\delta_{00|g}$ is based on the observation that $\delta_{00|g}=\mathbb{P}(S^{0}=0,S^{1}=0\mid G=g)=\mathbb{P}(S^{1}=0\mid G=g)=\mathbb{P}(S=0\mid A=1,G=g)$. It is easy to show that
$\sqrt{n}(\hat{\delta}_{11|g}-\delta_{11|g})={}$ $\frac{1}{\mathbb{P}(A=0,G=g)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)]+o_{\mathbb{P}}(1),$
$\sqrt{n}(\hat{\delta}_{01|g}-\delta_{01|g})={}$ $\frac{1}{\mathbb{P}(A=1,G=g)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(S_{i}=1,A_{i}=1,G_{i}=g)-\mathbb{P}(S=1,A=1,G=g)]$
-{} $\frac{1}{\mathbb{P}(A=0,G=g)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)]+o_{\mathbb{P}}(1),$
$\sqrt{n}(\hat{\delta}_{00|g}-\delta_{00|g})={}$ $\frac{1}{\mathbb{P}(A=1,G=g)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)-\mathbb{P}(S=0,A=1,G=g)]+o_{\mathbb{P}}(1).$
Second, we focus on $\hat{\pi}_{1|00}$ and $\hat{\tilde{\pi}}_{1|11}$.
$\hat{\pi}_{1|00}={}$ $\frac{n^{-1}\sum_{i=1}^{n}\mathbb{I}(Y_{i}=1,S_{i}=0,A_{i}=1)}{n^{-1}\sum_{i=1}^{n}\mathbb{I}(S_{i}=0,A_{i}=1)}$
$\hat{\tilde{\pi}}_{1|11}={}$ $\frac{n^{-1}\sum_{i=1}^{n}\mathbb{I}(Y_{i}=1,S_{i}=1,A_{i}=0)}{n^{-1}\sum_{i=1}^{n}\mathbb{I}(S_{i}=1,A_{i}=0)}.$
We could show that
$\sqrt{n}(\hat{\pi}_{1|00}-\pi_{1|00})$
$={}$ $\frac{1}{\mathbb{P}(S=0,A=1)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(Y_{i}=1,S_{i}=0,A_{i}=1)-\mathbb{P}(Y=1,S=0,A=1)]+o_{\mathbb{P}}(1),$
$\sqrt{n}(\hat{\tilde{\pi}}_{1|11}-\tilde{\pi}_{1|11})$
$={}$ $\frac{1}{\mathbb{P}(S=1,A=0)}\cdot\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[\mathbb{I}(Y_{i}=1,S_{i}=1,A_{i}=0)-\mathbb{P}(Y=1,S=1,A=0)]+o_{\mathbb{P}}(1).$
Third, we focus on $(\hat{\pi}_{1|01},\hat{\pi}_{1|11})$ and $(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01})$. For $(\hat{\pi}_{1\mid 01},\hat{\pi}_{1\mid 11})$, for ease of presentation, we let
$Y_{g}:={}$ $\mathbb{P}(Y=1\mid G=g,A=1)-\pi_{1|00}\delta_{00|g},$
$X_{g}:={}$ $(X_{1g},X_{2g})^{\intercal}=(\delta_{01|g},\delta_{11|g})^{\intercal}.$
Then the true value of $(\pi_{1|01},\pi_{1|11})^{\intercal}$ can be written as
$(\pi_{1|01},\pi_{1|11})^{\intercal}=\left(\frac{1}{m}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}X_{g}Y_{g},$
and the estimator is
$(\hat{\pi}_{1\mid 01},\hat{\pi}_{1\mid 11})^{\intercal}=\left(\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g},$
where $\hat{X}_{g}$ and $\hat{Y}_{g}$ are estimates of $X_{g}$ and $Y_{g}$, respectively. Note that
$\sqrt{n}(\hat{\pi}_{1|00}\hat{\delta}_{00|g}-\pi_{1|00}\delta_{00|g})$
$={}$ $\sqrt{n}(\hat{\pi}_{1|00}\hat{\delta}_{00|g}-\hat{\pi}_{1|00}\delta_{00|g})+\sqrt{n}(\hat{\pi}_{1|00}\delta_{00|g}-\pi_{1|00}\delta_{00|g})$
$={}$ $\pi_{1|00}\sqrt{n}(\hat{\delta}_{00|g}-\delta_{00|g})+\delta_{00|g}\sqrt{n}(\hat{\pi}_{1|00}-\pi_{1|00})+o_{\mathbb{P}}(1)$
$={}$ $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\pi_{1|00}\frac{\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)-\mathbb{P}(S=0,A=1,G=g)}{\mathbb{P}(A=1,G=g)}$
$+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\delta_{00|g}\frac{\mathbb{I}(Y_{i}=1,S_{i}=0,A_{i}=1)-\mathbb{P}(Y=1,S=0,A=1)}{\mathbb{P}(S=0,A=1)}+o_{\mathbb{P}}(1),$
we have
$\sqrt{n}(\hat{Y}_{g}-Y_{g})={}$ $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{\mathbb{I}(Y_{i}=1,G_{i}=g,A_{i}=1)-\mathbb{P}(Y=1,G=g,A=1)}{\mathbb{P}(A=1,G=g)}$
-{} $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\pi_{1|00}\frac{\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)-\mathbb{P}(S=0,A=1,G=g)}{\mathbb{P}(A=1,G=g)}$
-{} $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\delta_{00|g}\frac{\mathbb{I}(Y_{i}=1,S_{i}=0,A_{i}=1)-\mathbb{P}(Y=1,S=0,A=1)}{\mathbb{P}(S=0,A=1)}+o_{\mathbb{P}}(1).$
In addition,
$\sqrt{n}(\hat{X}_{g}-X_{g})=o_{\mathbb{P}}(1)+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}$
$\begin{pmatrix}\dfrac{\mathbb{I}(S_{i}=1,A_{i}=1,G_{i}=g)-\mathbb{P}(S=1,A=1,G=g)}{\mathbb{P}(A=1,G=g)}-\dfrac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}\\ \dfrac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}\end{pmatrix}$.
Then, by Lemma S1, Theorem S3(a) holds.
Similarly, we could show Theorem S3(b) by setting
$Y_{g}:={}$ $\mathbb{P}(Y=1\mid A=0,G=g)-\tilde{\pi}_{1|11}\delta_{11|g},$
$X_{g}:={}$ $(X_{1g},X_{2g})^{\intercal}=(\delta_{00|g},\delta_{01|g})^{\intercal},$
observing that the true value of $({\tilde{\pi}}_{1|00},{\tilde{\pi}}_{1|01})$ can be written as
$(\pi_{1|01},\pi_{1|11})^{\intercal}=\left(\frac{1}{m}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}X_{g}Y_{g},$
and the estimator is
$(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01})^{\intercal}=\left(\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{X}_{g}^{\intercal}\right)^{-1}\cdot\frac{1}{m}\sum_{g=1}^{m}\hat{X}_{g}\hat{Y}_{g},$
and
$\sqrt{n}(\hat{X}_{g}-X_{g})=o_{\mathbb{P}}(1)+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}$
$\begin{pmatrix}\dfrac{\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)-\mathbb{P}(S=0,A=1,G=g)}{\mathbb{P}(A=1,G=g)}\\ \dfrac{\mathbb{I}(S_{i}=1,A_{i}=1,G_{i}=g)-\mathbb{P}(S=1,A=1,G=g)}{\mathbb{P}(A=1,G=g)}-\dfrac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}\end{pmatrix}$
$\sqrt{n}(\hat{Y}_{g}-Y_{g})={}$ $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{\mathbb{I}(Y_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(Y=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}$
-{} $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\tilde{\pi}_{1|11}\frac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}$
-{} $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\delta_{11|g}\frac{\mathbb{I}(Y_{i}=1,S_{i}=1,A_{i}=0)-\mathbb{P}(Y=1,S=1,A=0)}{\mathbb{P}(S=1,A=0)}+o_{\mathbb{P}}(1).$
S1.8 Proof of the Asymptotic Distribution of the Test Statistic
Recall that for each trial $g=1,\dots,m$, $\tilde{X}_{g}=(\tilde{X}_{1g},\tilde{X}_{2g})^{\intercal}$, $\hat{X}_{g}=(\hat{X}_{1g},\hat{X}_{2g})^{\intercal}$, $\theta=(\pi_{1|0},\pi_{1|1})^{\intercal}$ is the true value of the parameter. Under the null hypothesis,
$\tilde{Y}_{g}=\tilde{X}_{g}^{\intercal}\theta\qquad\text{for all }g.$
Define the residual evaluated at $\theta$ as $\hat{\epsilon}_{\theta,g}=\hat{Y}_{g}-\hat{X}_{g}^{\intercal}\theta$, and let
$\hat{\boldsymbol{\epsilon}}_{\theta}=(\hat{\epsilon}_{\theta,1},\dots,\hat{\epsilon}_{\theta,m})^{\intercal}.$
We note that $\hat{Y}_{g}$ and $\hat{X}_{g}$ are asymptotically normal with convergence rate of order $1/\sqrt{n_{g}}$. It then follows that $\hat{\boldsymbol{\epsilon}}_{\theta}$ is asymptotically normal. When the proportion $n_{g}/n$ converges to a strictly positive constant in $(0,1)$, we can use CLT to easily prove that
$\sqrt{n}\,\hat{\boldsymbol{\epsilon}}_{\theta}\xrightarrow{d}N(\mathbf{0},\boldsymbol{\Omega}),$
where $\boldsymbol{\Omega}$ is a diagonal matrix of full rank $m$. This follows from the independence of samples across trials, which implies that $\operatorname{cov}(\hat{\epsilon}_{\theta,g},\hat{\epsilon}_{\theta,g^{\prime}})=0$ for $g\neq g^{\prime}$.
Let $\hat{\epsilon}_{g}=\hat{Y}_{g}-\hat{X}_{g}^{\intercal}\hat{\theta}$ be the residual evaluated at $\hat{\theta}$, and $\hat{\boldsymbol{\epsilon}}=(\hat{\epsilon}_{1},\dots,\hat{\epsilon}_{m})^{\intercal}$, then
$J=(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})^{\intercal}\,\boldsymbol{\Sigma}_{*}^{+}\,(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})=n\;\hat{\boldsymbol{\epsilon}}^{\intercal}\,\boldsymbol{\Sigma}_{*}^{+}\,\hat{\boldsymbol{\epsilon}}$,
where $\boldsymbol{\Sigma}_{*}^{+}$ denotes the Moore–Penrose pseudo‑inverse of $\boldsymbol{\Sigma}_{*}$,
[…]
$\tilde{\mathbf{P}}=\tilde{\mathbf{X}}(\tilde{\mathbf{X}}^{\intercal}\tilde{\mathbf{X}})^{-1}\tilde{\mathbf{X}}^{\intercal}$ is the orthogonal projection matrix onto the column space of $\tilde{\mathbf{X}}$ (rank $2$), and $\tilde{\mathbf{X}}$ be the $m\times 2$ matrix whose $g$-th row is $\tilde{X}_{g}^{\intercal}=(\tilde{X}_{1g},\tilde{X}_{2g})$.
Next, we focus on showing that $J\xrightarrow{d}\chi^{2}_{m-2}$.
OLS Estimation and Residuals. Set $\hat{\mathbf{Y}}=(\hat{Y}_{1},\dots,\hat{Y}_{m})^{\intercal}$ and let $\hat{\mathbf{X}}$ be the $m\times 2$ matrix whose $g$-th row is $\hat{X}_{g}^{\intercal}$. The OLS estimator is
$\hat{\theta}=(\hat{\mathbf{X}}^{\intercal}\hat{\mathbf{X}})^{-1}\hat{\mathbf{X}}^{\intercal}\hat{\mathbf{Y}}$.
By the definition of $\hat{\boldsymbol{\epsilon}}_{\theta}$, we could write $\hat{\mathbf{Y}}=\hat{\mathbf{X}}\theta+\hat{\boldsymbol{\epsilon}}_{\theta}$. Then
$\hat{\theta}-\theta=(\hat{\mathbf{X}}^{\intercal}\hat{\mathbf{X}})^{-1}\hat{\mathbf{X}}^{\intercal}\hat{\boldsymbol{\epsilon}}_{\theta}=(\tilde{\mathbf{X}}^{\intercal}\tilde{\mathbf{X}})^{-1}\tilde{\mathbf{X}}^{\intercal}\hat{\boldsymbol{\epsilon}}_{\theta}+o_{\mathbb{P}}(1/\sqrt{n}),$
where the last equality follows from $(\hat{\mathbf{X}}^{\intercal}\hat{\mathbf{X}})^{-1}-(\tilde{\mathbf{X}}^{\intercal}\tilde{\mathbf{X}})^{-1}=O_{\mathbb{P}}(1)$, $\hat{\mathbf{X}}-\tilde{\mathbf{X}}=o_{\mathbb{P}}(1)$, and $\hat{\boldsymbol{\epsilon}}_{\theta}=O_{\mathbb{P}}(n^{-1/2})$.
The OLS residual vector $\hat{\boldsymbol{\epsilon}}$ is
$\hat{\boldsymbol{\epsilon}}=\hat{\mathbf{Y}}-\hat{\mathbf{X}}\hat{\theta}=\hat{\boldsymbol{\epsilon}}_{\theta}-\hat{\mathbf{X}}(\hat{\theta}-\theta).$
Substituting the expansion for $\hat{\theta}-\theta$ gives
$\hat{\boldsymbol{\epsilon}}=\hat{\boldsymbol{\epsilon}}_{\theta}-\tilde{\mathbf{X}}(\tilde{\mathbf{X}}^{\intercal}\tilde{\mathbf{X}})^{-1}\tilde{\mathbf{X}}^{\intercal}\hat{\boldsymbol{\epsilon}}_{\theta}+o_{\mathbb{P}}(1/\sqrt{n})=(\mathbf{I}_{m}-\tilde{\mathbf{P}})\hat{\boldsymbol{\epsilon}}_{\theta}+o_{\mathbb{P}}(1/\sqrt{n}).$
Next, multiplying both sides by $\sqrt{n}$, we obtain
$\sqrt{n}\,\hat{\boldsymbol{\epsilon}}=(\mathbf{I}_{m}-\tilde{\mathbf{P}})\sqrt{n}\,\hat{\boldsymbol{\epsilon}}_{\theta}+o_{\mathbb{P}}(1)\xrightarrow{d}N(\mathbf{0},\boldsymbol{\Sigma}_{*}).$
The matrix $\boldsymbol{\Sigma}_{*}$ is singular of rank m2m-2.
Asymptotic distribution of $J$-Statistic. Let $\mathbf{B}=\mathbf{I}_{m}-\tilde{\mathbf{P}}$, then $\hat{\boldsymbol{\epsilon}}=\mathbf{B}\hat{\boldsymbol{\epsilon}}_{\theta}+o_{\mathbb{P}}(1/\sqrt{n})$. Since $\mathbf{B}$ is symmetric idempotent with $\operatorname{rank}(\mathbf{B})=m-2$, there exists an $m\times(m-2)$ matrix $\mathbf{V}$ whose columns form an orthonormal basis for the column space of $\mathbf{B}$, i.e.,
$\mathbf{V}^{\intercal}\mathbf{V}=\mathbf{I}_{m-2},\quad\mathbf{V}\mathbf{V}^{\intercal}=\mathbf{B}.$
Project the scaled residuals onto this subspace:
$\mathbf{z}=\mathbf{V}^{\intercal}(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})=\mathbf{V}^{\intercal}\mathbf{B}(\sqrt{n}\,\hat{\boldsymbol{\epsilon}}_{\theta})+o_{\mathbb{P}}(1)=\mathbf{V}^{\intercal}(\mathbf{V}\mathbf{V}^{\intercal})(\sqrt{n}\,\hat{\boldsymbol{\epsilon}}_{\theta})+o_{\mathbb{P}}(1)=\mathbf{V}^{\intercal}(\sqrt{n}\,\hat{\boldsymbol{\epsilon}}_{\theta})+o_{\mathbb{P}}(1).$
Hence, we have
$\mathbf{z}\xrightarrow{d}N(\mathbf{0},\,\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}),$
and $\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}$ is a positive definite $(m-2)\times(m-2)$ matrix.
In addition, observe that
$(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})^{\intercal}\mathbf{V}(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\mathbf{V}^{\intercal}(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})=\mathbf{z}^{\intercal}(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\mathbf{z}+o_{\mathbb{P}}(1),$
and since $\mathbf{z}\xrightarrow{d}N(\mathbf{0},\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})$, the standard quadratic form theory yields
$(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})^{\intercal}\mathbf{V}(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\mathbf{V}^{\intercal}(\sqrt{n}\,\hat{\boldsymbol{\epsilon}})\xrightarrow{d}\chi^{2}_{m-2}.$
Therefore, to establish that $J\xrightarrow{d}\chi^{2}_{m-2}$, it suffices to verify that
$\boldsymbol{\Sigma}_{*}^{+}=\mathbf{V}\,(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}.$
This can be shown by the following verification:
  • Recall that $\boldsymbol{\Sigma}_{*}=\mathbf{B}\boldsymbol{\Omega}\mathbf{B}=\mathbf{V}\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}\mathbf{V}^{\intercal}$, $\mathbf{V}^{\intercal}\mathbf{V}=\mathbf{I}_{m-2}$. By calculation,
    $\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}=\mathbf{V}\,(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}\mathbf{V}\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}\mathbf{V}^{\intercal}=\mathbf{V}\,(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}\mathbf{V}^{\intercal}=\mathbf{V}\mathbf{V}^{\intercal}=\mathbf{B}.$
    $\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}=\mathbf{V}\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}\mathbf{V}^{\intercal}\mathbf{V}\,(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}=\mathbf{V}\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V}(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}=\mathbf{V}\mathbf{V}^{\intercal}=\mathbf{B}.$
  • Verify $\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}=\boldsymbol{\Sigma}_{*}^{+}$. It holds by noting that $\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}=\mathbf{B}\boldsymbol{\Sigma}_{*}^{+}=\mathbf{B}\mathbf{V}\,(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}=\mathbf{V}\mathbf{V}^{\intercal}\mathbf{V}\,(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}=\mathbf{V}(\mathbf{V}^{\intercal}\boldsymbol{\Omega}\mathbf{V})^{-1}\,\mathbf{V}^{\intercal}=\boldsymbol{\Sigma}_{*}^{+}$.
  • Verify $\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}=\boldsymbol{\Sigma}_{*}$. It holds by noting that $\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}=\mathbf{B}\boldsymbol{\Sigma}_{*}=\mathbf{B}\mathbf{B}\boldsymbol{\Omega}\mathbf{B}=\mathbf{B}\boldsymbol{\Omega}\mathbf{B}=\boldsymbol{\Sigma}_{*}$. ($\mathbf{B}$ is idempotent matrix)
  • Verify $(\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+})^{\intercal}=\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}$ and $(\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*})^{\intercal}=\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}$. They hold by noting that $\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{*}^{+}=\mathbf{B}=\boldsymbol{\Sigma}_{*}^{+}\boldsymbol{\Sigma}_{*}$ and $\mathbf{B}$ is symmetrical.
Discussion: Incorporation of Covariates
In the main text, we focus on the non-covariate, nonparametric multi-trial identification framework. This is consistent with the ACCT application in Section 7 where no individual-level baseline covariates are available. At the same time, incorporating baseline covariates is important. Covariates may improve the plausibility of the transportability assumption by absorbing observed case-mix differences, and in many settings they are also needed either for confounding adjustment or for capturing effect heterogeneity that is not adequately represented in the marginal untreated potential outcomes. This importance is also reflected in subsequent methodological development: Shahn and Madigan [2025] extend our framework to settings with covariates. In this subsection, we provide a brief discussion on the implications of incorporating covariates, and refer to Shahn and Madigan [2025] for more details.
To clarify the technical implications of incorporating covariates, it is helpful to distinguish two roles of covariates.
Role 1: Confounding adjustment.
If treatment assignment is randomized only conditional on baseline covariates, or if one considers an observational analogue, then the arm-specific counterfactual risks within each trial (e.g., $\mathbb{P}(Y^{0}\mid G=g)$ and $\mathbb{P}(Y^{1}\mid G=g)$) must first be estimated after adjusting for covariates, for example by standardization, weighting, or doubly robust procedures. Our identifying equations then apply after replacing simple sample proportions by these covariate-adjusted risk estimates.
Role 2: Effect-heterogeneity adjustment.
A second and more direct role is that the key transportability assumption itself may only hold after conditioning on additional baseline covariates. In this case, Assumption 2 becomes a conditional invariance assumption.
We formalize this setting below for the binary-outcome case.
Assumption S1 (Conditional unconfoundedness of trials).
For all $g\in{\mathcal{G}}=\{1,\ldots,m\}$, (i) $A\perp\!\!\!\perp(Y^{0},Y^{1})\mid G=g,X$, and (ii) $0<\mathbb{P}(A=1\mid G=g,X)<1$ almost surely.
Assumption S2 (Conditional transportability of state transition probability).
$Y^{1}\perp\!\!\!\perp G\mid Y^{0},X$.
For each covariate value $x$ in the common support of $X$, define
$p_{1g}(x):=\mathbb{P}(Y=1\mid A=1,G=g,X=x),\qquad p_{0g}(x):=\mathbb{P}(Y=1\mid A=0,G=g,X=x).$
Under Assumption S1,
$p_{1g}(x)=\mathbb{P}(Y^{1}=1\mid G=g,X=x),\qquad p_{0g}(x)=\mathbb{P}(Y^{0}=1\mid G=g,X=x).$
Moreover, Assumption S2 implies that the following transitional probabilities do not depend on $G$:
$\pi_{1\mid 0}(x):=\mathbb{P}(Y^{1}=1\mid Y^{0}=0,X=x),\qquad\pi_{1\mid 1}(x):=\mathbb{P}(Y^{1}=1\mid Y^{0}=1,X=x),$
Hence, for every $g\in{\mathcal{G}}$ and every $x$,
$p_{1g}(x)=\pi_{1\mid 0}(x)\{1-p_{0g}(x)\}+\pi_{1\mid 1}(x)p_{0g}(x).$ (S1)
In addition to Assumptions S1S2, one also needs common support of $X$ across the relevant trials. Otherwise, the conditional transition law at a given covariate value cannot be learned from all trials. Moreover, we also need to extend the rank condition to incorporate the covariates.
Condition S1 (Overlap Condition).
For every $g=1,\dots,m$, $\mathbb{P}(G=g\mid X)>0$ almost surely.
Condition S2 (Conditional full-column rank).
For every $x$ in the common support of $X$, the matrix $\bigl(\mathbb{P}(Y^{0}=0\mid G=g,X=x),\,\mathbb{P}(Y^{0}=1\mid G=g,X=x)\bigr)_{m\times 2}$ has full column rank.
Theorem S1 (Covariate-adjusted binary outcome).
Under Assumptions S1S2 and Conditions S1S2, the conditional distributions $\mathbb{P}(Y^{1}\mid Y^{0},G=g,X=x)$ and $\mathbb{P}(Y^{1},Y^{0}\mid G=g,X=x)$ are identifiable for all $g\in{\mathcal{G}}$. Consequently, the marginal joint distribution for trial $g$ is identifiable through
$\mathbb{P}(Y^{1},Y^{0}\mid G=g)=\int\mathbb{P}(Y^{1},Y^{0}\mid G=g,X=x)\,dF_{X\mid G=g}(x).$
Theorem S1 shows that the law-of-total-probability argument in the main text extends pointwise in $X$. At the same time, it also makes clear how the required conditions change once covariates are introduced. In particular, the positivity and rank conditions now become conditional on $X$, and the identifying variation must remain after conditioning on covariates.
Low-dimensional discrete covariates.
If $X$ takes only a small number of discrete values, then estimation and inference are conceptually straightforward. One may proceed stratum by stratum. For each covariate stratum $X=x$, estimate $p_{0g}(x)$ and $p_{1g}(x)$, solve the same least-squares problem as in Section 4 of the main text to recover $\pi_{1\mid 0}(x)$ and $\pi_{1\mid 1}(x)$, reconstruct $\mathbb{P}(Y^{1},Y^{0}\mid G=g,X=x)$, and then standardize over $X\mid G=g$ to obtain $\mathbb{P}(Y^{1},Y^{0}\mid G=g)$.
Continuous or high-dimensional covariates.
When the relevant covariates are continuous or high-dimensional, the stratum-by-stratum strategy is no longer viable. Exact stratification leads to sparsity, the functions $\pi_{1\mid 0}(x)$ and $\pi_{1\mid 1}(x)$ become unknown infinite-dimensional objects, and the conditional arm-specific risks $p_{0g}(x)$ and $p_{1g}(x)$ must also be estimated as functions of $x$. As a result, one typically needs additional structure.
A natural approach is to impose finite-dimensional parametric models such as
$\pi_{1\mid 0}(x)=f(x;\beta^{\star}),\qquad\pi_{1\mid 1}(x)=h(x;\lambda^{\star}),$
where $\beta^{\star}$ and $\lambda^{\star}$ are finite-dimensional parameters and $f(\cdot;\beta)$ and $h(\cdot;\lambda)$ are known functions, for example generalized linear models. Under these models, equation (S1) becomes
$p_{1g}(x)=f(x;\beta^{\star})\{1-p_{0g}(x)\}+h(x;\lambda^{\star})p_{0g}(x).$
In this parametric regime, one no longer needs to impose the previous pointwise rank condition for every $x$; rather, it suffices that the observed data provide enough variation across trials to identify the finite-dimensional parameters $(\beta^{\star},\lambda^{\star})$. In other words, parametric identification requires weaker identifying variations across trials.
One possible estimation strategy is to first construct nuisance estimators $\hat{p}_{1g}(x)$ and $\hat{p}_{0g}(x)$, and then solve
$(\hat{\beta},\hat{\lambda})\in\arg\min_{\beta,\lambda}\frac{1}{n}\sum_{g=1}^{m}\sum_{i=1}^{n}I(G_{i}=g)\Bigl[\hat{p}_{1g}(X_{i})-f(X_{i};\beta)\{1-\hat{p}_{0g}(X_{i})\}-h(X_{i};\lambda)\hat{p}_{0g}(X_{i})\Bigr]^{2}.$
In this regime, however, the problem is no longer a direct extension of the nonparametric least-squares framework in the main text. The main challenge shifts from solving a finite-dimensional linear system to learning covariate-dependent nuisance objects and covariate-dependent transition laws. In particular, high-dimensional or flexible nuisance estimation typically calls for more structured semiparametric or debiased approaches. See Shahn and Madigan [2025] for one such development based on Neyman-orthogonal moments and sample splitting.
Overall, conditioning on measured baseline covariates may improve the plausibility of the transportability assumption, because some between-trial heterogeneity that would otherwise appear as violations of $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$ may instead be explained by observed covariates. On the other hand, this comes at a cost: the relevant overlap and rank conditions become conditional on $X$, and continuous or high-dimensional covariates require additional modeling and inference machinery. Analogous covariate-adjusted extensions in the principal-stratification setting are possible but would be substantially more involved, and we do not pursue them here.
Implementation of the Test for Assumption 2
In this section, we describe a feasible procedure for conducting the test of Assumption 2 using the $J$-statistic. The main practical challenge is estimating the asymptotic covariance matrix $\boldsymbol{\Sigma}_{*}$ of the infeasible sample moment vector $\sqrt{n}\hat{\boldsymbol{\epsilon}}$; see Section S1.8 for details. We use the bootstrap to obtain a consistent estimator $\hat{\boldsymbol{\Sigma}}_{*}^{\text{boot}}$, from which we compute the required pseudo-inverse. The steps below outline the procedure.
  1. 1.
    Obtain the OLS estimator $\hat{\theta}$, and compute the OLS residuals $\hat{\boldsymbol{\epsilon}}=\hat{\mathbf{Y}}-\hat{\mathbf{X}}\hat{\theta}$.
  2. 2.
    Use the bootstrap to estimate $\boldsymbol{\Sigma}_{*}$. This gives $\hat{\boldsymbol{\Sigma}}_{*}^{\text{boot}}$. The detailed procedure is as follows:
    1. (a)
      For $b=1,\dots,B$ (e.g., $B=500$):
      1. i.
        Generate a bootstrap sample by resampling the original observations with replacement. To avoid empty cells in the trials defined by $G$, we use a stratified bootstrap: resampling is performed separately within each trial $g$.
      2. ii.
        From the bootstrap sample, compute $\hat{Y}_{g}^{(b)}$, $\hat{X}_{g}^{(b)}$, and $\hat{\theta}^{(b)}$ in exactly the same way as from the original data.
      3. iii.
        Form the bootstrap residual vector $\hat{\boldsymbol{\epsilon}}^{(b)}$ with components
        $\hat{\epsilon}_{g}^{(b)}=\hat{Y}_{g}^{(b)}-\hat{X}_{g}^{\intercal}\hat{\theta}^{(b)}.$
    2. (b)
      Compute the sample covariance matrix of the $\sqrt{n}\,\hat{\boldsymbol{\epsilon}}^{(b)}$ vectors across the $B$ bootstrap replicates:
[…]
      where $\bar{\hat{\boldsymbol{\epsilon}}}=\frac{1}{B}\sum_{b=1}^{B}\hat{\boldsymbol{\epsilon}}^{(b)}$.
  3. 3.
    Compute the pseudo-inverse $\hat{\boldsymbol{\Sigma}}_{*}^{+}$ and then compute $J=n\cdot\hat{\boldsymbol{\epsilon}}^{\intercal}\hat{\boldsymbol{\Sigma}}_{*}^{+}\hat{\boldsymbol{\epsilon}}$. Finally, compare the $J$-statistic to the $\chi^{2}_{m-2}$ critical value.
Estimation Method Based on Corollary 1 and Theorem 4
In this section, we present the estimation method for the joint distribution $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ and the principal stratification average causal effect $\textup{PSACE}_{ab|g}$ based on Corollary 1 and Theorem 4, respectively.
S4.1 Estimation Method Based on Corollary 1
We first give the estimation method for $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ and $\textup{PSACE}_{ab|g}$ for $g\in\mathcal{G}$ based on Corollary 1. Assumption 4 implies that $\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b)$ for $a,b,c,d=0,1$ are invariant parameters across trials. The proposed estimation method consists of the following steps.
  • Step 1. estimate $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ for $a,b,c,d=0,1$. We denote the estimators as $\hat{\mathbb{P}}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b\mid G=g)$.
  • Step 2. estimate $\mathbb{P}(S^{1}=0,Y^{1}=0\mid S^{0}=a,Y^{0}=b)$ for $a,b=0,1$ by conducting a linear regression of $\hat{\mathbb{P}}(S^{1}=0,Y^{1}=0\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=0,Y^{0}=1|G=g),\hat{\mathbb{P}}(S^{0}=1,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=1,Y^{0}=1|G=g))$. Similarly, we can estimate $\mathbb{P}(S^{1}=0,Y^{1}=1\mid S^{0}=a,Y^{0}=b)$, $\mathbb{P}(S^{1}=1,Y^{1}=0\mid S^{0}=a,Y^{0}=b)$, and $\mathbb{P}(S^{1}=1,Y^{1}=1\mid S^{0}=a,Y^{0}=b)$ for $a,b=0,1$. We denote the estimator as $\hat{\mathbb{P}}(S^{1}=d,Y^{1}=d\mid S^{0}=a,Y^{0}=b)$.
  • Step 3. estimate the joint distribution $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$. Based on $\hat{\mathbb{P}}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b)$, the estimator of $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$ is given as
    $\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$
    $={}$ $\hat{\mathbb{P}}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b)\cdot\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b\mid G=g)$
  • Step 4. estimate $\text{PSACE}_{ab|g}$, which is defiend by
    $\text{PSACE}_{ab|g}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b,G=g)-\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b,G=g),~a=0,1;b=0,1.$
    The estimator for $\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b,G=g)$ is given as
    $\hat{\mathbb{P}}(Y^{1}=1\mid$ $S^{0}=a,S^{1}=b,G=g)={}\frac{\hat{\mathbb{P}}(Y^{1}=1,S^{0}=a,S^{1}=b\mid G=g)}{\hat{\mathbb{P}}(S^{0}=a,S^{1}=b\mid G=g)}$
    $={}$ $\frac{\sum_{d=0,1}\hat{\mathbb{P}}(Y^{1}=1,Y^{0}=d,S^{0}=a,S^{1}=b\mid G=g)}{\sum_{c,d=0,1}\hat{\mathbb{P}}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g)}$
    Likewise, we can estimate $\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b,G=g)$ with
    $\hat{\mathbb{P}}(Y^{0}=1\mid$ $S^{0}=a,S^{1}=b,G=g)={}\frac{\hat{\mathbb{P}}(Y^{0}=1,S^{0}=a,S^{1}=b\mid G=g)}{\hat{\mathbb{P}}(S^{0}=a,S^{1}=b\mid G=g)}$
    $={}$ $\frac{\sum_{d=0,1}\hat{\mathbb{P}}(Y^{0}=1,Y^{1}=d,S^{0}=a,S^{1}=b\mid G=g)}{\sum_{c,d=0,1}\hat{\mathbb{P}}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g)}$
Similar to the estimation method in Section S5.3, another robust approach to estimate the invariance parameters is to solve a least squares optimization problem under several equality and inequality constraints.
S4.2 Estimation Method Based on Theorem 4
We then present the estimation method for $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ and $\textup{PSACE}_{ab|g}$ for $g\in\mathcal{G}$ based on Theorem 4. Under Assumption 4,
$\pi_{cd|ab}=\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b),\quad a,b,c,d=0,1,$
are invariant across trials. When the monotonicity condition $S^{1}\geq S^{0}$ and $Y^{1}\geq Y^{0}$ holds, $\pi_{00|01}=\pi_{10|01}=\pi_{00|10}=\pi_{01|10}=\pi_{00|11}=\pi_{01|11}=\pi_{10|11}\equiv 0$, and $\pi_{11|11}=1$, as shown in Table S1. We can estimate the other invariant parameters and $\textup{PSACE}_{ab|g}$ with the following steps.
  • Step 1. estimate $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ for $a,b,c,d=0,1$. We denote the estimators as $\hat{\mathbb{P}}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b\mid G=g)$.
  • Step 2. estimate the other invariant parameters and the joint distribution $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$.
    • estimate $\pi_{00|00}$ by linear regression of $\hat{\mathbb{P}}(S^{1}=0,Y^{1}=0\mid G=g)$ on $\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g)$;
    • estimate $(\pi_{01|00},\pi_{01|01})$ by linear regression of $\hat{\mathbb{P}}(S^{1}=0,Y^{1}=1\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=0,Y^{0}=1|G=g))$;
    • estimate $(\pi_{10|00},\pi_{10|10})$ by linear regression of $\hat{\mathbb{P}}(S^{1}=1,Y^{1}=0\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=1,Y^{0}=0|G=g))$;
    • estimate $\pi_{11|00}$ with $\pi_{11|00}=1-\pi_{00|00}-\pi_{01|00}-\pi_{10|00}$, $\pi_{11|01}$ with $\pi_{11|01}=1-\pi_{01|01}$, and $\pi_{11|01}$ with $\pi_{11|10}=1-\pi_{10|10}$.
    Let $\hat{\pi}_{cd|ab}$ be the estimator of $\pi_{cd|ab}$, then the estimator of the joint distribution $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$ is given as
    $\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)=\hat{\pi}_{cd|ab}\cdot\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b\mid G=g)$
  • Step 4. estimate $\text{PSACE}_{ab|g}$. The estimator for $\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b,G=g)$ is
    $\hat{\mathbb{P}}(Y^{1}=1\mid$ $S^{0}=a,S^{1}=b,G=g)={}\frac{\hat{\mathbb{P}}(Y^{1}=1,S^{0}=a,S^{1}=b\mid G=g)}{\hat{\mathbb{P}}(S^{0}=a,S^{1}=b\mid G=g)}$
    $={}$ $\frac{\sum_{d=0,1}\hat{\mathbb{P}}(Y^{1}=1,Y^{0}=d,S^{0}=a,S^{1}=b\mid G=g)}{\sum_{c,d=0,1}\hat{\mathbb{P}}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g)}$
    Likewise, we can estimate $\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b,G=g)$.
Similar to the estimation method in Section S5.3, another robust approach to estimate the invariance parameters is to solve a least squares optimization problem under several equality and inequality constraints.
S4.3 Variance Estimation via Bootstrap
In both the simulation and the application, we use a nonparametric stratified bootstrap for variance estimation. The procedure is as follows:
  • Step 1: Resample within each stratum. For each bootstrap replication $b=1,\ldots,B$:
    • For each trial $g$, draw a bootstrap sample by sampling $n_{g}$ observations with replacement from the original $n_{g}$ observations in that trial.
    • Combine the resampled observations to form the complete bootstrap dataset.
  • Step 2: Recompute the estimator. Using each bootstrap dataset $b$ (b=1,,Bb=1,...,B), re-estimate the parameter $\theta$, denoted as $\hat{\theta}^{(b)}$.
  • Step 3: Variance estimation. After obtaining $\{\hat{\theta}^{(b)}\}_{b=1}^{B}$, compute the estimated variance:
    $\widehat{\mathrm{Var}}(\hat{\theta})=\frac{1}{B-1}\sum_{b=1}^{B}\left(\hat{\theta}^{(b)}-\bar{\theta}\right)^{2},\qquad\bar{\theta}=\frac{1}{B}\sum_{b=1}^{B}\hat{\theta}^{(b)}.$
Principal Causal Effects under Partial Monotonicity
In this section, we extend the results of Corollary 1 under partial monotonicity condition ($S^{1}\geq S^{0}$ or $Y^{1}\geq Y^{0}$, but not both).
S5.1 Under the Monotonicity Condition $S^{1}\geq S^{0}$
We first extend the results of Corollary 1 under $S^{1}\geq S^{0}$. Let
$\pi_{cd|ab}=\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b),\quad a,b,c,d=0,1,$
which are invariant parameters across trials by Assumption 4. When the monotonicity condition $S^{1}\geq S^{0}$ holds, $\pi_{00|10}=\pi_{01|10}=\pi_{00|11}=\pi_{01|11}=0$, and it has 12 possible values of $(a,b,c,d)$, as shown in Table S2.
Table S2: Parameters under Assumption 4 and the monotonicity condition $S^{1}\geq S^{0}$
$S^{0}=a$ $Y^{0}=b$ $S^{1}=c$ $Y^{1}=d$ Restriction Number of Free Parameters
0 0 0 0 $\sum_{c=0}^{1}\sum_{d=0}^{1}\pi_{cd|00}=1$ 3
0 0 0 1
0 0 1 0
0 0 1 1
0 1 0 0 $\sum_{c=0}^{1}\sum_{d=0}^{1}\pi_{cd|01}=1$ 3
0 1 0 1
0 1 1 0
0 1 1 1
1 0 1 0 $\sum_{d=0}^{1}\pi_{1d|10}=1$ 1
1 0 1 1
1 1 1 0 $\sum_{d=0}^{1}\pi_{1d|11}=1$ 1
1 1 1 1
On the other hand, under Assumption 4, we have the following system of equations
[…]
where $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ are identifiable quantities under Assumption 3.
The above system of equations include a total of $m\times 3$ linearly independent equations due to $\sum_{c,d}\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)=1$. However, it contains $8$ unknown free parameters as shown in the last column of Table S2. It seems that these invariant parameters $\pi_{cd|ab}$ for $a,b,c,d=0,1$ are identifiable when $m\geq 3$ under regular conditions. However, it is not this case due to the fact that the first two rows of the system of equations including only two unknown parameters. Formally, we have the following conclusion, as shown in Corollary S1.
Corollary S1.
If $S^{1}\geq S^{0}$, then under Assumptions 3, 4, and Condition 4(i), $\pi_{00|10}=\pi_{01|10}=\pi_{00|11}=\pi_{01|11}=0$, and $\{\pi_{00|00},\pi_{00|01},\pi_{01|00},\pi_{01|01}\}$ are identifiable.
In addition, observing that the third row of the system of equations has four free parameters. To identify them, it still requires Condition 3, Thus, we cannot relax Condition 3 for identifying the additional parameters, including $\{\pi_{10|00},\pi_{10|01},\pi_{10|10},\pi_{10|11},\pi_{11|00},\pi_{11|01},\pi_{10110},\pi_{11|11}\}$. As a result, we cannot relax the Condition 3 for identifying the joint distribution $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$ under the additional monotonicity condition $S^{1}\geq S^{0}$. Nevertheless, the additional monotonicity condition $S^{1}\geq S^{0}$ allows us to slightly simplify the estimation method, see Section S5.3 for details.
S5.2 Under the Monotonicity Condition $Y^{1}\geq Y^{0}$
We then extend the results of Corollary 1 under the monotonicity condition $Y^{1}\geq Y^{0}$. Under Assumption 4,
$\pi_{cd|ab}=\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b),\quad a,b,c,d=0,1,$
are invariant across trials. If $Y^{1}\geq Y^{0}$ holds, then $\pi_{00|01}=\pi_{10|01}\equiv 0$ and $\pi_{00|11}=\pi_{10|11}\equiv 0$. It has 12 possible values of $(a,b,c,d)$, as shown in Table S3.
Table S3: Parameters under Assumption 4 and the monotonicity condition $Y^{1}\geq Y^{0}$
$S^{0}=a$ $Y^{0}=b$ $S^{1}=c$ $Y^{1}=d$ Restriction Number of Free Parameters
0 0 0 0 $\sum_{c=0}^{1}\sum_{d=0}^{1}\pi_{cd|00}=1$ 3
0 0 0 1
0 0 1 0
0 0 1 1
0 1 0 1 $\sum_{c=0}^{1}\pi_{c1|01}=1$ 1
0 1 1 1
1 0 0 0 $\sum_{d=0}^{1}\sum_{d=0}^{1}\pi_{cd|10}=1$ 3
1 0 0 1
1 0 1 0
1 0 1 1
1 1 0 1 $\sum_{d=0}^{1}\pi_{c1|11}=1$ 1
1 1 1 1
On the other hand, we have the following system of equations
$\begin{cases}\mathbb{P}(S^{1}=0,Y^{1}=0|G=g)={}&\pi_{00|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{00|10}\cdot\mathbb{P}(S^{0}=1,Y^{0}=0|G=g),\\ \mathbb{P}(S^{1}=0,Y^{1}=1|G=g)={}&\pi_{01|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{01|01}\cdot\mathbb{P}(S^{0}=0,Y^{0}=1|G=g)\\ \hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad+{}&\pi_{01|10}\cdot\mathbb{P}(S^{0}=1,Y^{0}=0|G=g)+\pi_{01|11}\cdot\mathbb{P}(S^{0}=1,Y^{0}=1|G=g)\\ \mathbb{P}(S^{1}=1,Y^{1}=0|G=g)={}&\pi_{10|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{10|10}\cdot\mathbb{P}(S^{0}=1,Y^{0}=0|G=g)\\ \mathbb{P}(S^{1}=1,Y^{1}=1|G=g)={}&\pi_{11|00}\cdot\mathbb{P}(S^{0}=0,Y^{0}=0|G=g)+\pi_{11|01}\cdot\mathbb{P}(S^{0}=0,Y^{0}=1|G=g)\\ \hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad+{}&\pi_{11|10}\cdot\mathbb{P}(S^{0}=1,Y^{0}=0|G=g)+\pi_{11|11}\cdot\mathbb{P}(S^{0}=1,Y^{0}=1|G=g),\\ \qquad\text{for }g=1,...,m.\end{cases}$ (S3)
where $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ are identifiable quantities under Assumption 3.
Similar to the conclusion in Section S5.1, we have the following conclusion, as shown in Corollary S2.
Corollary S2.
If $Y^{1}\geq Y^{0}$, then under Assumptions 3, 4, and Condition 4(ii), $\pi_{00|01}=\pi_{10|01}=\pi_{00|11}=\pi_{10|11}\equiv 0$, and $\{\pi_{00|00},\pi_{00|10},\pi_{10|00},\pi_{10|10}\}$ are identifiable.
Under the additional monotonicity condition $Y^{1}\geq Y^{0}$, Corollary 1 also requires Condition 3 for identifying the full joint distribution $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$. Nevertheless, it can be used to simplify the estimation method for $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$ and $\text{PSACE}_{ab|g}$. See Section S5.3 for details.
S5.3 Estimation Method under Partial Monotonicity
We give a detailed description on the corresponding estimation method for $\mathbb{P}(S^{0},S^{1},Y^{0},Y^{1}\mid G=g)$ and $\textup{PSACE}_{ab|g}$ for $g\in\mathcal{G}$ based on the identification in Corollary 1, together with additional partial monotonicity $Y^{1}\geq Y^{0}$ or $S^{1}\geq S^{0}$. These additional monotonicity conditions simplify the identification following the analyses in Sections S5.1 and S5.2, so the estimation method can be also simplified accordingly. We mainly present the estimation method under $Y^{1}\geq Y^{0}$, as the method under $S^{1}\geq S^{0}$ is similar.
1. Under the monotonicity condition $Y^{1}\geq Y^{0}$.
Denote $\pi_{cd|ab}=\mathbb{P}(S^{1}=c,Y^{1}=d\mid S^{0}=a,Y^{0}=b),\quad a,b,c,d=0,1,$ as invariant parameters across trials. We set $\pi_{00|01}=\pi_{10|01}\equiv 0$ and $\pi_{00|11}=\pi_{10|11}\equiv 0$ due to $Y^{1}\geq Y^{0}$. Then, the estimation procedures are given as follows:
  • Step 1. estimate $\mathbb{P}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g)$ for $a,b,c,d=0,1$. We denote the estimators as $\hat{\mathbb{P}}(S^{1}=c,Y^{1}=d\mid G=g)$ and $\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b\mid G=g)$.
  • Step 2. estimate the other invariant parameters and the joint distribution $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$.
    • from equation (S3), we estimate $(\pi_{00|00},\pi_{00|10})$ by linear regression of $\hat{\mathbb{P}}(S^{1}=0,Y^{1}=0\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=1,Y^{0}=0|G=g))$;
    • from equation (S3), we estimate $(\pi_{01|00},\pi_{01|01},\pi_{01|10},\pi_{01|11})$ by linear regression of $\hat{\mathbb{P}}(S^{1}=0,Y^{1}=1\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=0,Y^{0}=1|G=g),\hat{\mathbb{P}}(S^{1}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=1,Y^{0}=1|G=g))$;
    • from equation (S3), we estimate $(\pi_{10|00},\pi_{10|10})$ by linear regression of $\hat{\mathbb{P}}(S^{1}=1,Y^{1}=0\mid G=g)$ on $(\hat{\mathbb{P}}(S^{0}=0,Y^{0}=0|G=g),\hat{\mathbb{P}}(S^{0}=1,Y^{0}=0|G=g))$;
    • estimate $\pi_{11|ab}$ with $\pi_{11|ab}=1-\pi_{00|ab}-\pi_{01|ab}-\pi_{10|ab}$ for $a,b=0,1$.
    Based on $\hat{\pi}_{cd|ab}$, the estimator of $\mathbb{P}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)$ is given as
    $\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b,S^{1}=c,Y^{1}=d\mid G=g)=\pi_{cd|ab}\cdot\hat{\mathbb{P}}(S^{0}=a,Y^{0}=b\mid G=g)$
  • Step 4. estimate $\text{PSACE}_{ab|g}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b,G=g)-\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b,G=g)$. The estimator for $\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b,G=g)$ is given as
    $\hat{\mathbb{P}}(Y^{1}=1\mid$ $S^{0}=a,S^{1}=b,G=g)={}\frac{\hat{\mathbb{P}}(Y^{1}=1,S^{0}=a,S^{1}=b\mid G=g)}{\hat{\mathbb{P}}(S^{0}=a,S^{1}=b\mid G=g)}$
    $={}$ $\frac{\sum_{d=0,1}\hat{\mathbb{P}}(Y^{1}=1,Y^{0}=d,S^{0}=a,S^{1}=b\mid G=g)}{\sum_{c,d=0,1}\hat{\mathbb{P}}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g)}$
    Likewise, we can estimate $\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b,G=g)$ with
    $\hat{\mathbb{P}}(Y^{0}=1\mid$ $S^{0}=a,S^{1}=b,G=g)={}\frac{\hat{\mathbb{P}}(Y^{0}=1,S^{0}=a,S^{1}=b\mid G=g)}{\hat{\mathbb{P}}(S^{0}=a,S^{1}=b\mid G=g)}$
    $={}$ $\frac{\sum_{d=0,1}\hat{\mathbb{P}}(Y^{0}=1,Y^{1}=d,S^{0}=a,S^{1}=b\mid G=g)}{\sum_{c,d=0,1}\hat{\mathbb{P}}(S^{0}=a,S^{1}=b,Y^{0}=c,Y^{1}=d\mid G=g)}$
Further Consideration. In the estimation procedures above, we do not impose the restriction that $\pi_{cd|ab}$ must belong to the interval [0, 1]. To address this issue, a straightforward approach is to use the restricted least-squares method in Step 2 instead of directly applying the standard least-squares method. However, this approach may not guarantee that $\sum_{c,d=0,1}\hat{\pi}_{cd|ab}=1$.
To avoid this problem, we can estimate the invariant parameters $\pi_{cd|ab}$ by solving a least squares programming under equality and inequality constraints. Specifically, we denote $\theta=(\pi_{00|00},\pi_{00|01},\pi_{00|10},\pi_{00|11},\pi_{01|00},\pi_{01|01},\pi_{01|10},\pi_{01|11},\pi_{10|00},\pi_{10|01},\pi_{10|10},\pi_{10|11},\pi_{11|00},\pi_{11|01},\pi_{11|10},\pi_{11|11})^{\intercal}$ as the vector of all invariant parameters, let
$\begin{cases}\tilde{Y}_{00}={}&(\mathbb{P}(S^{1}=0,Y^{1}=0\mid G=1),\mathbb{P}(S^{1}=0,Y^{1}=0\mid G=2),\cdots,\mathbb{P}(S^{1}=0,Y^{1}=0\mid G=m))^{\intercal}\\ \tilde{Y}_{01}={}&(\mathbb{P}(S^{1}=0,Y^{1}=1\mid G=1),\mathbb{P}(S^{1}=0,Y^{1}=1\mid G=2),\cdots,\mathbb{P}(S^{1}=0,Y^{1}=1\mid G=m))^{\intercal}\\ \tilde{Y}_{10}={}&(\mathbb{P}(S^{1}=1,Y^{1}=0\mid G=1),\mathbb{P}(S^{1}=1,Y^{1}=0\mid G=2),\cdots,\mathbb{P}(S^{1}=1,Y^{1}=0\mid G=m))^{\intercal}\\ \tilde{Y}_{11}={}&(\mathbb{P}(S^{1}=1,Y^{1}=1\mid G=1),\mathbb{P}(S^{1}=1,Y^{1}=0\mid G=2),\cdots,\mathbb{P}(S^{1}=1,Y^{1}=1\mid G=m))^{\intercal}\end{cases}$
$\tilde{X}$=
$\begin{pmatrix}\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=1)&\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=1)&\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=1)&\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=1)\\ \mathbb{P}(S^{0}=0,Y^{0}=0\mid G=2)&\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=2)&\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=2)&\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=2)\\ \vdots\\ \mathbb{P}(S^{0}=0,Y^{0}=0\mid G=m)&\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=m)&\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=m)&\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=m)\end{pmatrix}$,
${\bf\tilde{X}}=\begin{pmatrix}\tilde{X}&\bm{0}&\bm{0}&\bm{0}\\ \bm{0}&\tilde{X}&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&\tilde{X}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}&\tilde{X}\end{pmatrix}_{4m\times 16},\quad{\bf\tilde{Y}}=\begin{pmatrix}\tilde{Y}_{00}\\ \tilde{Y}_{01}\\ \tilde{Y}_{10}\\ \tilde{Y}_{11}\end{pmatrix}$.
Under the condition $Y^{1}\geq Y^{0}$, solving $\theta$ can be formulated as follows:
[…]
where
[…]
${\bf d}=\begin{pmatrix}0\\ 0\\ 0\\ 0\\ 1\\ 1\\ 1\\ 1\end{pmatrix}$,
${\bf E}=\begin{pmatrix}{\bf I}_{16\times 16}\\ -{\bf I}_{16\times 16}\end{pmatrix},\quad{\bf f}=\begin{pmatrix}{\bf 0}_{16\times 1}\\ -{\bf 1}_{16\times 1}\end{pmatrix}$,
where ${\bf I}_{16\times 16}$ is is the identity matrix of dimension $16\times 16$, ${\bf 0}_{16\times 1}$ is a $16$-dimensional vector with each element being 0, and ${\bf I}_{16\times 1}$ is a $16$-dimensional vector with each element being -1.
In equation (S4), ${\bf C}\theta={\bf d}$ corresponds to 8 equality constraints: $\pi_{00|01}=0,\pi_{00|11}=0,\pi_{10|01}=0,\pi_{10|11}=0$, $\sum_{c,d=0,1}\hat{\pi}_{cd|00}=1$, $\sum_{c,d=0,1}\hat{\pi}_{cd|01}=1$, $\sum_{c,d=0,1}\hat{\pi}_{cd|10}=1$, $\sum_{c,d=0,1}\hat{\pi}_{cd|11}=1$. In addition, ${\bf E}\theta\geq{\bf f}$ corresponds to 32 inequality constraints: $\pi_{cd|ab}\geq 0$, $\pi_{cd|ab}\leq 1$ for $a,b,c,d=0,1$.
We can solve the optimization problem (S4) using the lsei function in R package lsei [Wang et al., 2020].
2. Under the monotonicity condition $S^{1}\geq S^{0}$. The estimation method under $S^{1}\geq S^{0}$ is similar to that under $Y^{1}\geq Y^{0}$, and is therefore omitted to avoid redundancy.
Further Discussion under a Weaker Version of Assumption 4
If we relax Assumption 4 to Assumption S3 below, we can obtain the identifiability for $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$, as shown in Corollary S3.
Assumption S3.
(i) $G\perp\!\!\!\perp S^{1}\mid S^{0},Y^{0}$; (ii) $G\perp\!\!\!\perp Y^{1}\mid S^{0},Y^{0}$.
Corollary S3.
Under Assumption 3, Assumption S3, and Condition 3, the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$ for $g\in\mathcal{G}$ are identifiable.
It is noteworthy that $\mathbb{P}(S^{0},S^{1},Y^{1}\mid G=g)$ is not identifiable under the conditions outlined in Corollary S3, and therefore, we cannot identify $\text{PSACE}_{ab|g}$. Also, similar to Theorem 4, if we impose the monotonicity assumption ($Y^{1}\geq Y^{0}$ and $S^{1}\geq S^{0}$), we can relax Condition 3.
Corollary S4.
If $Y^{1}\geq Y^{0}$ and $S^{1}\geq S^{0}$, then under Assumptions 3, S3, and Condition 4, the joint distributions $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ and $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$ for $g\in\mathcal{G}$ are identifiable.
Proof of Corollary S3.
Let
$\pi_{c|ab}:=\mathbb{P}(S^{1}=c\mid S^{0}=a,Y^{0}=b)=\mathbb{P}(S^{1}=c\mid S^{0}=a,Y^{0}=b,G=g),\quad a,b,c=0,1,$
which are invariant parameters across trials.Under Assumption S3(i), we have that
$\mathbb{P}(S^{1}=1|G=g)=\sum_{a,b}\pi_{1|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b|G=g),\quad\text{for }g=1,...,m,$
where $\mathbb{P}(S^{1}=1|G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b|G=g)$ are identifiable quantities under Assumption 3. The above system of equations contain a total of $m$ linearly independent equations, and $4$ unknown free parameters (i.e., $\{(\pi_{1|ab},a,b=0,1\}$). Then under Assumption 3, S3(i), and Condition 3, $\pi_{1|ab}$ for $a=0,1;b=0,1$ are identifiable by solving the system of equations. This also implies the identifiability of $\mathbb{P}(S^{0},S^{1},Y^{0}|G=g)$ by noting that
$\mathbb{P}(S^{0}=a,S^{1}=1,Y^{0}=b|G=g)=\pi_{1|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b|G=g).$
similarly, we can show the identifiability of $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$.
Proof of Corollary S4.
Under Assumption S3(i), $\pi_{c|ab}:=\mathbb{P}(S^{1}=c\mid S^{0}=a,Y^{0}=b),\quad a,b,c=0,1$ are invariant parameters across trials. Due to $S^{1}\geq S^{0}$, we have $\pi_{0|10}=\pi_{0|11}=0$, which implies $\pi_{1|10}=\pi_{1|11}=1$. By Assumption S3(i),
$\mathbb{P}(S^{1}=1\mid G=g)=\sum_{a,b}\pi_{1|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g),$
$={}$ $\pi_{1|00}\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g)+\pi_{1|01}\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=g)$
$+\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=g)+\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=g)\qquad\text{for }g=1,...,m,$
where $\mathbb{P}(S^{1}=1|G=g)$ and $\mathbb{P}(S^{0}=a,Y^{0}=b|G=g)$ are identifiable under Assumption 3. The above system of equations contains a total of $m$ linearly independent equations, and $2$ unknown free parameters (i.e., $\{\pi_{1|00},\pi_{1|01}\}$). Thus, when Condition 4(i) holds, $\pi_{1|ab}$ for $a=0,1;b=0,1$ are identifiable by solving the system of equations. This also implies the identifiability of $\mathbb{P}(S^{0},S^{1},Y^{0}\mid G=g)$ by noting that $\mathbb{P}(S^{0}=a,S^{1}=1,Y^{0}=b\mid G=g)=\pi_{1|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g).$
Likewise, under Assumption S3(ii),
$\bar{\pi}_{d|ab}:=\mathbb{P}(Y^{1}=d\mid S^{0}=a,Y^{0}=b),\quad a,b,d=0,1,$
are invariant parameters across trials. Due to $Y^{1}\geq Y^{0}$, we have $\bar{\pi}_{0|01}=\bar{\pi}_{0|11}=0$, which implies $\bar{\pi}_{1|01}=\bar{\pi}_{1|11}=1$. By Assumption S3(ii),
$\mathbb{P}(Y^{1}=1\mid G=g)=\sum_{a,b}\bar{\pi}_{1|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g),$
$={}$ $\bar{\pi}_{1|00}\mathbb{P}(S^{0}=0,Y^{0}=0\mid G=g)+\mathbb{P}(S^{0}=0,Y^{0}=1\mid G=g)$
$+\bar{\pi}_{1|10}\mathbb{P}(S^{0}=1,Y^{0}=0\mid G=g)+\mathbb{P}(S^{0}=1,Y^{0}=1\mid G=g)\qquad\text{for }g=1,...,m.$
The above system of equations contains a total of $m$ linearly independent equations, and $2$ unknown free parameters (i.e., $\{\bar{\pi}_{1|00},\bar{\pi}_{1|10}\}$). Thus, when Condition 4(ii) holds, $\bar{\pi}_{1|ab}$ for $a=0,1;b=0,1$ are identifiable by solving the system of equations. This also implies the identifiability of $\mathbb{P}(S^{0},Y^{0},Y^{1}\mid G=g)$ by noting that $\mathbb{P}(S^{0}=a,Y^{0}=b,Y^{1}=1\mid G=g)=\bar{\pi}_{1|ab}\cdot\mathbb{P}(S^{0}=a,Y^{0}=b\mid G=g).$
Monotonicity-Based Route of Jiang et al. [2016]
In this section, we formally restate the monotonicity-based approach of Jiang et al. [2016] and provide an alternative proof of their identification theorem, which in turn motivates a least-squares estimator.
S7.1 Restatement of the Route of Jiang et al. [2016]
Jiang et al. [2016] adopt Assumption 3, together with the key Assumptions S4S5 and Condition S3, presented below.
Assumption S4.
$G\perp\!\!\!\perp Y^{a}\mid S^{0},S^{1}$ for $a=0,1$.
Assumption S5.
$S^{1}\geq S^{0}$.
Condition S3.
(i) The matrix $(\mathbb{P}(S^{0}=0,S^{1}=1\mid G=g),\mathbb{P}(S^{0}=1,S^{1}=1\mid G=g))_{m\times 2}$ is full-column rank; (ii) The matrix $(\mathbb{P}(S^{0}=0,S^{1}=1\mid G=g),\mathbb{P}(S^{0}=0,S^{1}=0\mid G=g))_{m\times 2}$ is full-column rank.
Assumption S4 means that there is no dependence between $G$ and $Y^{0}$ or $Y^{1}$ within the principal stratum defined by $(S^{0},S^{1})$. This implies that $\text{PSACE}_{ab|g}$ for $g\in{\mathcal{G}}$ are invariant across trials. In contrast, the assumptions in Section 5.1 allow different principal effects across trials.
The monotonicity condition in Assumption S5, together with the unconfoundedness Assumption 3, imply the identifiability of principal scores $\delta_{ab|g}=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$ for $a,b\in\{0,1\}$. Specifically, Assumption S5 implies $\delta_{10\mid g}=0$, thus the joint distribution $\mathbb{P}(S^{0},S^{1}\mid G=g)$ involves only three free parameters ($\delta_{00\mid g},\delta_{01\mid g},\delta_{11\mid g}$). These parameters satisfy three equations $\delta_{10\mid g}+\delta_{11\mid g}=\mathbb{P}(S^{0}=1\mid G=g),~\delta_{01\mid g}+\delta_{11\mid g}=\mathbb{P}(S^{1}=1\mid G=g),~\sum_{a=0}^{1}\sum_{b=0}^{1}\delta_{ab\mid g}=1.$ Solving these equations easily gives the identification of the principal scores.
Condition S3 is also a full rank condition, which, like Condition 4, also requires $m\geq 2$. Under Assumption 3 and the three above assumptions, Jiang et al. [2016] show the identifiability of $\text{PSACE}_{ab|g}$ in their Theorem 1. We reproduce their results in Theorem S2 below and provide an alternative proof (see Supplementary Material S1.6). This alternative proof will motivate our new estimator in Section S7.2.
Theorem S2.
Under Assumptions 3, S4, and S5, for g=1,,mg=1,...,m, we have that
(a) $\mathbb{P}(Y^{1}\mid S^{0},S^{1},G=g)$ is identifiable if Condition S3(i) holds.
(b) $\mathbb{P}(Y^{0}\mid S^{0},S^{1},G=g)$ is identifiable if Condition S3(ii) holds.
(c) $\text{PSACE}_{ab|g}$ for $a,b\in\{0,1\}$ are identifiable if Condition S3 holds.
In Theorem S2, the monotonicity condition $S^{1}\geq S^{0}$ in Assumption S5 is used to identify the principal scores as discussed above. One may wonder whether this condition can be replaced by an alternative condition, $G\perp\!\!\!\perp S^{1}\mid S^{0}$, that is also relevant for identifying the principal scores [^1] According to our theory in Section 3, the condition $G\perp\!\!\!\perp S^{1}\mid S_{0}$ is relevant for the identification of the joint distribution of $(S^{1},S^{0})$ and thus the principal scores.. Section S7.3 gives a negative answer to this question, showing the importance of the monotonicity condition. Moreover, Theorem 4 also considers another monotonicity condition $Y^{1}\geq Y^{0}$. This alternative monotonicity condition is not relevant here because the distributions $\mathbb{P}(Y^{1}\mid S^{0},S^{1},G=g)$ and $\mathbb{P}(Y^{0}\mid S^{0},S^{1},G=g)$ and the principal effects $\text{PSACE}_{ab|g}$ do not involve the joint distribution of $Y^{1}$ and $Y^{0}$.
Without the monotonicity condition $S^{1}\geq S^{0}$, Proposition 2 of Jiang et al. [2016] further shows that a necessary condition for local identifiability of the principal effects is $m\geq 3$. In contrast, our previous Corollary 1, which also considers identification without any monotonicity condition, establishes sufficient conditions for global identifiability when $m\geq 4$. In this sense, our new identification results provide some alternative conditions that lead to a stronger identification, which complements the findings of Jiang et al. [2016].
S7.2 Least-squares Estimation Based on Theorem S2
Under the identification in Theorem S2, Jiang et al. [2016] propose a Bayesian estimation method. In this part, we follow our new proof for Theorem S2 and extend the least-square estimation method in Section 4 to the setting of Theorem S2.
We first define the following parameters for $a,b\in\{0,1\}$:
$\pi_{1|ab}=\mathbb{P}(Y^{1}=1\mid S^{0}=a,S^{1}=b),~\tilde{\pi}_{1|ab}=\mathbb{P}(Y^{0}=1\mid S^{0}=a,S^{1}=b).$
These are key invariant parameters across trials. Clearly, under the monotonicity condition $S^{1}\geq S^{0}$, $\pi_{1|10}$ and $\tilde{\pi}_{1|10}$ are undefined and do not need estimation. Now we describe how to estimate $\beta:=(\pi_{1|00},\pi_{1|01},\pi_{1|11})$ and $\gamma:=(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01},\tilde{\pi}_{1|11})$.
Step 1: estimate the principal scores $\delta_{ab|g}:=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)$, and the probabilities $\mathbb{P}(Y^{1}=1\mid G=g)$, and $\mathbb{P}(Y^{0}=1\mid G=g)$. Specifically, for any given $g$, $\delta_{10|g}=0$ by Assumption S5, and
$\begin{cases}\delta_{11|g}={}&\mathbb{P}(S^{0}=1,S^{1}=1|G=g)=\mathbb{P}(S^{0}=1|G=g)=\mathbb{P}(S=1|A=0,G=g),\\ \delta_{01|g}={}&\mathbb{P}(S^{0}=0,S^{1}=1|G=g)=\mathbb{P}(S^{1}=1|G=g)-\delta_{11|g}\\ \qquad\qquad={}&\mathbb{P}(S=1|A=1,G=g)-\delta_{11|g},\\ \delta_{00|g}={}&1-\delta_{11|g}-\delta_{01|g}.\end{cases}$
In addition, $\mathbb{P}(Y^{1}=1\mid G=g)=\mathbb{P}(Y=1\mid G=g,A=1)$ and $\mathbb{P}(Y^{0}=1\mid G=g)=\mathbb{P}(Y=1\mid G=g,A=0)$. We can construct their estimators $\hat{\delta}_{11|g},\hat{\delta}_{01|g},\hat{\delta}_{00|g}$, $\hat{\mathbb{P}}(Y^{1}=1\mid G=g)$, and $\hat{\mathbb{P}}(Y^{0}=1\mid G=g)$ via replacing the population probabilities by the empirical frequencies and solving the resulting linear equations.
Step 2: estimate $\pi_{1|00}$ and $\tilde{\pi}_{1|11}$. Clearly, due to $S^{1}\geq S^{0}$, we have $\pi_{1|00}=\mathbb{P}(Y^{1}=1\mid S^{0}=0,S^{1}=0)=\mathbb{P}(Y^{1}=1\mid S^{1}=0)=\mathbb{P}(Y=1\mid S=0,A=1),$ and $\tilde{\pi}_{1|11}=\mathbb{P}(Y^{0}=1\mid S^{0}=1,S^{1}=1)=\mathbb{P}(Y^{0}=1\mid S^{0}=1)=\mathbb{P}(Y=1\mid S=1,A=0).$ These two can again be estimated by the empirical frequencies. We denote the estimators by $\hat{\pi}_{1|00}$ and $\hat{\tilde{\pi}}_{1|11}$ respectively.
Step 3: estimate $(\pi_{1|01},\pi_{1|11})$ and $(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01})$. Note that
$\mathbb{P}(Y^{1}=1|G=g)-\pi_{1|00}\delta_{00|g}={}$ $\pi_{1|01}\delta_{01|g}+\pi_{1|11}\delta_{11|g},\qquad g=1,...,m,$
$\mathbb{P}(Y^{0}=1|G=g)-\tilde{\pi}_{1|11}\delta_{11|g}={}$ $\tilde{\pi}_{1|00}\delta_{00|g}+\tilde{\pi}_{1|01}\delta_{01|g},\qquad g=1,...,m.$
Then we can estimate $(\pi_{1|01},\pi_{1|11})$ by running linear regression of $\hat{\mathbb{P}}(Y^{1}=1\mid G=g)-\hat{\pi}_{1|00}\hat{\delta}_{00|g}$ on $(\hat{\delta}_{01|g},\hat{\delta}_{11|g})$, and estimate $(\tilde{\pi}_{1|00},\tilde{\pi}_{1|01})$ by running linear regression of $\hat{\mathbb{P}}(Y^{0}=1\mid G=g)-\hat{\tilde{\pi}}_{1|11}\hat{\delta}_{11|g}$ on $(\hat{\delta}_{00|g},\hat{\delta}_{01|g})$. Let $(\hat{\pi}_{1|01},\hat{\pi}_{1|11})$ and $(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01})$ be the corresponding estimators, and denote $\hat{\beta}=(\hat{\pi}_{1|00},\hat{\pi}_{1|01},\hat{\pi}_{1|11})$ and $\hat{\gamma}=(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01},\hat{\tilde{\pi}}_{1|11})$.
Step 4: estimate principal effects $\text{PSACE}_{ab|g}$. The estimators are given by $\hat{\pi}_{1|ab}-\hat{\tilde{\pi}}_{1|ab}$ for $ab=00,01$, and 11. Due to the monotonicity condition $S^{1}\geq S^{0}$, $\text{PSACE}_{ab|g}$ for $ab=10$ is undefined and needs no estimation.
Similar to Theorem 3, we can show the large sample properties of the proposed estimators $\hat{\beta}:=(\hat{\pi}_{1|00},\hat{\pi}_{1|01},\hat{\pi}_{1|11})$ and $\hat{\gamma}:=(\hat{\tilde{\pi}}_{1|00},\hat{\tilde{\pi}}_{1|01},\hat{\tilde{\pi}}_{1|11})$, respectively. The corresponding results are presented in Theorem S3. For notational convenience, we slightly abuse notation for $X_{g}$, $Y_{g}$, and $C$ below.
Theorem S3.
It follows that
(a) $\sqrt{n}(\hat{\beta}-\beta)\xrightarrow{d}N(0,\sigma_{\beta}^{2}),$ where $\sigma_{\beta}^{2}$ is the covariance matrix of
$\begin{pmatrix}\dfrac{\mathbb{I}(Y_{i}=1,S_{i}=0,A_{i}=1)-\mathbb{P}(Y=1,S=0,A=1)}{\mathbb{P}(S=0,A=1)}\\ C^{-1}\left\{\dfrac{1}{m}\sum_{g=1}^{m}\Big(\phi_{g}(X_{i},A_{i},S_{i},Y_{i})Y_{g}+\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})X_{g}\Big)\right\}\end{pmatrix}$,
with $Y_{g}=\mathbb{P}(Y=1\mid G=g,A=1)-\pi_{1|00}\delta_{00|g}$, $X_{g}=(\delta_{01|g},\delta_{11|g})^{\intercal}$, $C=m^{-1}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}$,
$\varphi_{g}(X_{i},A_{i},S_{i},Y_{i})={}$ $\frac{\mathbb{I}(Y_{i}=1,A_{i}=1,G_{i}=g)-\mathbb{P}(Y=1,A=1,G=g)}{\mathbb{P}(A=1,G=g)}$
-{} $\pi_{1|00}\frac{\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)-\mathbb{P}(S=0,A=1,G=g)}{\mathbb{P}(A=1,G=g)}$
-{} $\delta_{00|g}\frac{\mathbb{I}(Y_{i}=1,S_{i}=0,A_{i}=1)-\mathbb{P}(Y=1,S=0,A=1)}{\mathbb{P}(S=0,A=1)},$
and
$\phi_{g}(X_{i},A_{i},S_{i},Y_{i})=$
$\begin{pmatrix}\dfrac{\mathbb{I}(S_{i}=1,A_{i}=1,G_{i}=g)-\mathbb{P}(S=1,A=1,G=g)}{\mathbb{P}(A=1,G=g)}-\dfrac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}\\ \dfrac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}\end{pmatrix}$.
(b) $\sqrt{n}(\hat{\gamma}-\gamma)\xrightarrow{d}N(0,\sigma_{\gamma}^{2}),$ where $\sigma_{\gamma}^{2}$ is the covariance matrix of
$\begin{pmatrix}C^{-1}\left\{\dfrac{1}{m}\sum_{g=1}^{m}\Big(\tilde{\phi}_{g}(X_{i},A_{i},S_{i},Y_{i})Y_{g}+\tilde{\varphi}_{g}(X_{i},A_{i},S_{i},Y_{i})X_{g}\Big)\right\}\\ \dfrac{\mathbb{I}(Y_{i}=1,S_{i}=1,A_{i}=0)-\mathbb{P}(Y=1,S=1,A=0)}{\mathbb{P}(S=1,A=0)}\end{pmatrix}$,
with $Y_{g}=\mathbb{P}(Y=1\mid A=0,G=g)-\tilde{\pi}_{1|11}\delta_{11|g}$, $X_{g}=(\delta_{00|g},\delta_{01|g})^{\intercal}$, $C=m^{-1}\sum_{g=1}^{m}X_{g}X_{g}^{\intercal}$,
$\tilde{\varphi}_{g}(X_{i},A_{i},S_{i},Y_{i})={}$ $\frac{\mathbb{I}(Y_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(Y=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}$
-{} $\tilde{\pi}_{1|11}\frac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}$
-{} $\delta_{11|g}\frac{\mathbb{I}(Y_{i}=1,S_{i}=1,A_{i}=0)-\mathbb{P}(Y=1,S=1,A=0)}{\mathbb{P}(S=1,A=0)}$
and
$\tilde{\phi}_{g}(X_{i},A_{i},S_{i},Y_{i})=$
$\begin{pmatrix}\dfrac{\mathbb{I}(S_{i}=0,A_{i}=1,G_{i}=g)-\mathbb{P}(S=0,A=1,G=g)}{\mathbb{P}(A=1,G=g)}\\ \dfrac{\mathbb{I}(S_{i}=1,A_{i}=1,G_{i}=g)-\mathbb{P}(S=1,A=1,G=g)}{\mathbb{P}(A=1,G=g)}-\dfrac{\mathbb{I}(S_{i}=1,A_{i}=0,G_{i}=g)-\mathbb{P}(S=1,A=0,G=g)}{\mathbb{P}(A=0,G=g)}\end{pmatrix}$.
The proof of Theorem S3 is provided in Section S1.7
S7.3 Further Discussion on Theorem S2 without Monotonicity
In this section, we examine the identifiability of $\mathbb{P}(Y^{1},S^{0},S^{1}|G=g)$ and $\mathbb{P}(Y^{0},S^{0},S^{1}|G=g)$, replacing the monotonicity assumption $S^{1}\geq S^{0}$ with the following Assumption S6. We primarily show that $\mathbb{P}(Y^{1},S^{0},S^{1}|G=g)$ and $\mathbb{P}(Y^{0},S^{0},S^{1}|G=g)$ are not identifiable in this case.
Assumption S6.
$S^{1}\perp\!\!\!\perp G\mid S^{0}$.
According the proposed method in Section 3.2 of the manuscript, under Assumptions 3 and S6,
$\mathbb{P}(S^{1}=b\mid S^{0}=a)$
are invariant parameters across trials, and they are identifiable if $(\mathbb{P}(S^{0}=0\mid G=g),\mathbb{P}(S^{0}=1\mid G=g))_{m\times 2}$ is full column rank. In addition, $\mathbb{P}(S^{0}=a\mid G=g)$ is identifiable under Assumption 3. Thus, the principal score $\mathbb{P}(S^{0},S^{1}\mid G=g)$ is identifiable with
$\delta_{ab|g}:=\mathbb{P}(S^{0}=a,S^{1}=b\mid G=g)=\mathbb{P}(S^{1}=b\mid S^{0}=a)\cdot\mathbb{P}(S^{0}=a\mid G=g),$
On the other hand, under Assumption S4,
$\pi_{1|ab}:=\mathbb{P}(Y^{1}=1|S^{0}=a,S^{1}=b,G=g)$
are also invariant parameters across trials. In addition, we have the following decomposion,
$\mathbb{P}(Y^{a}=1\mid G=g)={}$ $\pi_{1|00}\delta_{00|g}+\pi_{1|01}\delta_{01|g}+\pi_{1|10}\delta_{10|g}+\pi_{1|11}\delta_{11|g},\qquad g=1,...,m,$
It seems that the parameters $\{\pi_{1|ab}:a,b=0,1\}$ are identifiable by solving the above system of equations, under Condition S4 below.
Condition S4.
$m\geq 4$, the matrix $(\mathbb{P}(S^{0}=0,S^{1}=0|G=g),\mathbb{P}(S^{0}=0,S^{1}=1|G=g),\mathbb{P}(S^{0}=1,S^{1}=0|G=g),\mathbb{P}(S^{0}=1,S^{1}=1|G=g))_{m\times 4}$ is full column rank.
However, Condition S4 can never hold under Assumption S6 due to the collinearity among $\delta_{ab|g}$. Specifically,
$\frac{\delta_{01|g}}{\delta_{00|g}}=\frac{\mathbb{P}(S^{1}=1\mid S^{0}=0)\cdot\mathbb{P}(S^{0}=0\mid G=g)}{\mathbb{P}(S^{1}=0\mid S^{0}=0)\cdot\mathbb{P}(S^{0}=0\mid G=g)}=\frac{\mathbb{P}(S^{1}=1\mid S^{0}=0)}{\mathbb{P}(S^{1}=0\mid S^{0}=0)},$
which is a constant independent of $g$. Similarly, $\delta_{11|g}/\delta_{10|g}$ is a constant.
Discussion: Extension to General Treatments
In this section, we discuss extensions to more general treatment settings. For a multivalued treatment, one would need a compatible collection of transportability assumptions linking several potential outcomes, together with corresponding rank conditions, and the compatibility of these assumptions is nontrivial. As an example, suppose the treatment takes values in $\{0,1,2\}$, binary outcome $Y\in\{0,1\}$, and let $\{Y^{0},Y^{1},Y^{2}\}$ denote the corresponding potential outcomes. Following the proposed method in the manuscript, a plausible set of identifiability assumptions can be formulated as follows.
Assumption S7.
(i) $(Y^{0},Y^{1},Y^{2})\perp\!\!\!\perp A\mid G=g$, $0<\mathbb{P}(A=1\mid G=g)<1$;
(ii) $Y^{1}\perp\!\!\!\perp G\mid Y^{0}$;
(iii) The matrix $(\mathbb{P}(Y^{0}=0\mid G=g),\mathbb{P}(Y^{0}=1\mid G=g))_{m\times 2}$ has a full-column rank;
(iv) $Y^{2}\perp\!\!\!\perp G\mid(Y^{0},Y^{1})$;
(v) The matrix $(\mathbb{P}(Y^{0}=0,Y^{1}=0\mid G=g),\mathbb{P}(Y^{0}=0,Y^{1}=1\mid G=g),\mathbb{P}(Y^{0}=1,Y^{1}=0\mid G=g),\mathbb{P}(Y^{0}=1,Y^{1}=1\mid G=g))_{m\times 4}$ has a full-column rank.
Intuitively, Assumption S7(i) identifies the marginal distributions of the potential outcomes. Assumptions S7(i)–(iii), which are exactly the same as the conditions in Theorem 1 of the manuscript, identify the joint distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$. Following the same rationale, one might consider further using Assumptions S7(iv)–(v) to identify the conditional distribution $\mathbb{P}(Y^{2}\mid Y^{0},Y^{1},G=g)$ and hence the joint distribution $\mathbb{P}(Y^{0},Y^{1},Y^{2}\mid G=g)$. However, these assumptions are not compatible: Assumption S7(ii) in fact contradicts Assumption S7(v). Specifically, if Assumption S7(ii) holds, then
$\frac{\mathbb{P}(Y^{0}=0,Y^{1}=1\mid G=g)}{\mathbb{P}(Y^{0}=0,Y^{1}=0\mid G=g)}=\frac{\mathbb{P}(Y^{1}=1\mid Y^{0}=0)\cdot\mathbb{P}(Y^{0}=0\mid G=g)}{\mathbb{P}(Y^{1}=0\mid Y^{0}=0)\cdot\mathbb{P}(Y^{0}=0\mid G=g)}=\frac{\mathbb{P}(Y^{1}=1\mid Y^{0}=0)}{\mathbb{P}(Y^{1}=0\mid Y^{0}=0)},$
which is a constant independent of $g$. This violates the rank condition in Assumption S7(v).
One possible way to identify $\mathbb{P}(Y^{0},Y^{1},Y^{2}\mid G=g)$ is to replace Assumptions S7(ii)–(iii) with a monotonicity condition $Y^{1}\geq Y^{0}$. The associated results are shown in Lemma S2.
Lemma S2.
Under Assumptions S7(i), (iv), and (v), together with the condition $Y^{1}\geq Y^{0}$, $\mathbb{P}(Y^{0},Y^{1},Y^{2}\mid G=g)$ is identifiable.
Considering the role of the monotonicity condition, an alternative set of identifiability assumptions can be formulated as follows.
Assumption S8 (Second Set of Possible Assumptions).
(i) $(Y^{0},Y^{1},Y^{2})\perp\!\!\!\perp A\mid G=g$, $0<\mathbb{P}(A=1\mid G=g)<1$;
(ii) $Y^{2}\geq Y^{1}\geq Y^{0}$.
Proposition S1.
Under Assumption S8, $\mathbb{P}(Y^{0},Y^{1},Y^{2}\mid G=g)$ is identifiable.
Although sequential monotonicity ($Y^{2}\geq Y^{1}\geq Y^{0}$) can be used to identify the joint distribution of the potential outcomes, it does not follow the same rationale as the proposed method in the manuscript.
For continuous treatment, the potential outcome process $\{Y^{a}:a\in\mathcal{A}\}$ is infinite-dimensional. Our finite linear-system identification argument does not extend directly to identify the joint distribution of the whole potential outcome process.
Additional Numerical Results
S9.1 Additional Results for Simulation
Additional simulations for the just-identified case. We consider the just-identified case helps better evaluate the finite-sample performance of the proposed estimators. We include two additional simulation cases (C5)–(C6). These two cases follow the same data-generating processes as cases (C1)–(C2) in the manuscript, except that the number of trials is set to 2 to ensure just-identified settings. The results are presented in Table S5. As shown, the Bias is small and the CP95 values are close to 0.95, indicating that the proposed method performs well overall.
However, by comparing Table S5 with Table 1 of the manuscript, we observe that: (1) the SD and ESE are substantially larger than those in Table 1 (cases (C1)–(C2)) of the manuscript; and (2) The SE and ESE are very close in Table 1 even when $n_{g}=100$, whereas the SE and ESE in Table S5 are unstable when $n_{g}=100$, indicating that a larger sample size is required to obtain stable results in just-identified cases.
Table S4: Simulation results for cases (C5)–(C6).
$n_{g}=100$ $n_{g}=200$ $n_{g}=500$
Case $\theta$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
(C5) $\pi_{1|0}$ -0.014 0.289 0.312 0.964 -0.006 0.181 0.184 0.952 -0.001 0.104 0.107 0.945
$\pi_{1|1}$ 0.009 0.154 0.164 0.964 0.003 0.098 0.102 0.951 -0.001 0.059 0.060 0.948
(C6) $\pi_{1|0}$ -0.016 0.345 0.296 0.960 -0.008 0.166 0.166 0.958 -0.004 0.090 0.093 0.945
$\pi_{1|1}$ 0.006 0.161 0.146 0.965 0.004 0.089 0.088 0.950 0.001 0.049 0.050 0.952
Note: Bias and SD are the Monte Carlo bias and standard deviation over the 1000 simulations of the point estimates, ESE and CP95 are the estimated asymptotic variances and coverage proportions of the 95% confidence intervals based on 100 bootstraps, respectively.
Additional simulations on the violation of Assumption 2. We also perform an additional simulation design in which Assumption 2 is violated by allowing the transition probabilities $\mathbb{P}(Y^{1}=1\mid Y^{0},G=g)$ to vary between trials to different degrees. We consider the following data-generating process:
  • (C7): $\mathbb{P}(Y^{1}=1\mid Y^{0},G=g)=\text{expit}(Y^{0}-0.5+\gamma_{g})$ for g=1,,10g=1,...,10, $\Gamma\leq\gamma_{g}\leq\Gamma$, where $\text{expit}(x)=\exp(x)/\{1+\exp(x)\}$ is the standard logistic function, $\Gamma\geq 0$ is a constant, set to 0.1, 0.2, and 0.3 in our analysis. For each simulation, $\gamma_{g}$ is drawn from the uniform distribution $\text{Uniform}(-\Gamma,\Gamma)$. For each trial g=1,2,,10g=1,2,...,10, the potential outcome $Y^{0}$ follows from a Bernoulli distribution with $\mathbb{P}(Y^{0}=1\mid G=g)=0.5+(g-1)/30$, i.e., taking evenly spaced values at equal intervals from 0.5 to 0.8, the binary treatment $A$ is randomly assigned with probability $\mathbb{P}(A=1\mid G=g)=0.5$, the sample size is set to 200.
Table S5: Simulation results for case (C7).
$\Gamma=0.1$ $\Gamma=0.2$ $\Gamma=0.3$
Trial $\theta_{g}$ Bias SD ESE CP95 Bias SD ESE CP95 Bias SD ESE CP95
1 $\pi_{1|0,1}$ 0.038 0.106 0.108 0.937 0.043 0.124 0.109 0.924 -0.050 0.129 0.110 0.896
$\pi_{1|1,1}$ 0.006 0.058 0.059 0.929 0.015 0.068 0.059 0.856 -0.076 0.072 0.060 0.810
2 $\pi_{1|0,2}$ 0.020 0.106 0.108 0.942 0.055 0.124 0.109 0.914 0.077 0.129 0.110 0.906
$\pi_{1|1,2}$ -0.013 0.058 0.059 0.933 0.028 0.068 0.059 0.856 0.048 0.072 0.060 0.802
3 $\pi_{1|0,3}$ 0.039 0.106 0.108 0.936 0.018 0.124 0.109 0.905 -0.009 0.129 0.110 0.894
$\pi_{1|1,3}$ 0.007 0.058 0.059 0.930 -0.010 0.068 0.059 0.879 -0.040 0.072 0.060 0.818
4 $\pi_{1|0,4}$ -0.003 0.106 0.108 0.936 0.049 0.124 0.109 0.900 -0.018 0.129 0.110 0.885
$\pi_{1|1,4}$ -0.035 0.058 0.059 0.926 0.021 0.068 0.059 0.873 -0.048 0.072 0.060 0.807
5 $\pi_{1|0,5}$ 0.010 0.106 0.108 0.938 0.003 0.124 0.109 0.905 -0.040 0.129 0.110 0.878
$\pi_{1|1,5}$ -0.023 0.058 0.059 0.942 -0.026 0.068 0.059 0.876 -0.068 0.072 0.060 0.851
6 $\pi_{1|0,6}$ 0.044 0.106 0.108 0.939 0.008 0.124 0.109 0.899 0.074 0.129 0.110 0.868
$\pi_{1|1,6}$ 0.012 0.058 0.059 0.932 -0.021 0.068 0.059 0.881 0.045 0.072 0.060 0.855
7 $\pi_{1|0,7}$ 0.016 0.106 0.108 0.939 0.040 0.124 0.109 0.896 0.062 0.129 0.110 0.863
$\pi_{1|1,7}$ -0.017 0.058 0.059 0.936 0.012 0.068 0.059 0.893 0.033 0.072 0.060 0.877
8 $\pi_{1|0,8}$ 0.025 0.106 0.108 0.938 -0.025 0.124 0.109 0.895 0.049 0.129 0.110 0.863
$\pi_{1|1,8}$ -0.008 0.058 0.059 0.933 -0.052 0.068 0.059 0.902 0.018 0.072 0.060 0.886
9 $\pi_{1|0,9}$ 0.001 0.106 0.108 0.936 -0.017 0.124 0.109 0.889 0.011 0.129 0.110 0.858
$\pi_{1|1,9}$ -0.031 0.058 0.059 0.939 -0.044 0.068 0.059 0.912 -0.020 0.072 0.060 0.883
10 $\pi_{1|0,10}$ 0.021 0.106 0.108 0.931 -0.010 0.124 0.109 0.890 0.036 0.129 0.110 0.846
$\pi_{1|1,10}$ -0.012 0.058 0.059 0.939 -0.038 0.068 0.059 0.919 0.004 0.072 0.060 0.892
Average 0.005 0.082 0.083 0.936 0.002 0.096 0.084 0.893 0.004 0.101 0.085 0.862
Note: Bias and SD are the Monte Carlo bias and standard deviation over the 1000 simulations of the point estimates, ESE and CP95 are the estimated asymptotic variances and coverage proportions of the 95% confidence intervals based on 100 bootstraps, respectively.
For case (C7), $\Gamma\geq 0$ is the sensitivity parameter, with larger values indicating greater violations of Assumption 2. When $\Gamma>0$, there are no invariant parameters across trials. We let $\theta_{g}=(\pi_{1\mid 0,g},\pi_{1\mid 1,g})$ denote the state transition probabilities for each trial $g$, where $\pi_{1\mid 0,g}=\mathbb{P}(Y^{1}=1\mid Y^{0}=0,G=g)$ and $\pi_{1\mid 1,g}=\mathbb{P}(Y^{1}=1\mid Y^{0}=1,G=g)$. If we nonetheless treat them as invariant parameters and apply the proposed estimation method, the corresponding numerical results are shown in Table S5. From Table S5, we observe that as the degree of violation ($\Gamma$) increases, the point estimates of $\pi_{1\mid 0,g}$ and $\pi_{1\mid 1,g}$ become increasingly distorted, and the coverage of the nominal $95\%$ confidence intervals deteriorates. On average, however, the bias remains small, while CP95 declines from 0.925 to 0.849.
S9.2 Additional Results for Application
We estimate the joint distributions $\mathbb{P}(S^{0},S^{1}\mid G=g)$ and $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for $g\in{\mathcal{G}}$. Figure S1 displays the point estimates along with the corresponding pointwise 95% confidence intervals. From Figure S1, we observe that the point estimates of the joint distributions show only slight volatility across all trials, indicating low heterogeneity of these distributions among trials. Additionally, the estimated standard error of the point estimates is relatively small, ranging from 0.024 to 0.058. Moreover, it is noteworthy that the lower bounds of the 95% confidence interval for $\hat{\mathbb{P}}(S^{0}=1,S^{1}=0\mid G=g)$ for all trials are slightly greater than 0, suggesting that the monotonicity assumption (Assumption S5, $S^{1}\geq S^{0}$) may not hold. In contrast, the 95% confidence intervals for $\hat{\mathbb{P}}(Y^{0}=1,Y^{1}=0\mid G=g)$ cover 0 in all trials.
Figure S1: (a) Estimated joint distribution $\mathbb{P}(S^{0},S^{1}\mid G=g)$ for all $g\in{\mathcal{G}}$; (b) Estimated joint distribution $\mathbb{P}(Y^{0},Y^{1}\mid G=g)$ for all $g\in{\mathcal{G}}$;